A Complete Bibliography of The American Statistician: 2010–2019

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

07 October 2024 Version 1.29

Title word cross-reference

1 [Por19]. $29.99 [Sta16]. 2 × 2 [Che11, SVR11]. $79.95 [Sab19]. −X ± t_{α}(n − 1)(s/\sqrt{\{n\}}) [Zuo10]. F [ZPM^{+}11]. G [Zil19]. k [CCB16]. P [BS11, Goo19, Gre19, HO16, Ioa19, KH19, RM19a, BB19, Bet19, BGW^{+}19, Bro10, Col19, Dem16a, Fou20, Fra19, FBHW19, GSK19, IGRP17, KS19, Kme19, KW19, Mar17, MHWB19, Rou19, WL16b, Zil19]. P (X > Y) [Bro10]. p < 0.05 [KS19, Mat19, Tra19, WSL19]. Q [LFH16]. R [BG17, RM19a]. S [Gre19]. t [Bro10, DAEP18, Joh19, WL16a, Zuo10].

[Sab19].

-Means [CCB16]. -Observation [Por19]. -POD [CCB16]. -Squared [BG17, RM19a]. -Statistics [DAEP18]. -Test [Bro10, WL16a]. -Value [Bet19, Dem16a, FBHW19, GSK19, Fra19, MHWB19, BS11, KH19]. -Values [BB19, BGW^{+}19, Col19, Fou20, IGRP17, KS19, Kme19, KW19, Mar17, Rou19, WL16b, Goo19, Gre19, HO16, RM19a, Zil19].

157-169 [Muk14b]. 175th [CSBB^{+}15]. 182 [IIDdL14]. 1950s [Fie13a].

2 [Sab19]. 200th [Rod15]. 21st [RHGS^{+}19]. 21st-Century [RHGS^{+}19]. 24 [Sch10b]. 242 [IDE15]. 244 [Muk14a]. 260 [ANBB^{+}14]. 276 [Din14a]. 2nd [Ber19].

325 [Fri15].

55 [Fri15].

66 [MS14]. 67 [ANBB^{+}14, BXHH14, Hol14, IIDdL14, Muk14a, Muk14b]. 68 [Din14a, IDE15].

73 [Fou20].

9 [Sta16]. 978 [Sab19, Sta16]. 978-1-107-63768-9 [Sta16]. 978-1-49-877013-2 [Sab19].

Abandon [MGG^{+}19]. Absolute [Hon15a, RLS^{+}16]. Absurdity [GR13a]. Academic [PBD^{+}16, SFSM17, Van15, Wal18]. Account [Bro11]. Accuracy [KP11, Lom13, RM19b, WSF10]. Accurate [LCD10]. Achieve [HO17]. Acquisition [ZHM^{+}13]. ACT [All19]. ACT-Tested [All19]. Action [GJLW19]. Activity [WRH19]. Adding [Gre10, HR10]. Addressing [HC19, LWM15]. Adjusted [BG17, Chr17]. Adjustment [RM19a, VZH16]. Advanced [RST15]. Advantages [LCG16]. Advocating [TdCP17]. After [FBHW19, ZR10]. Against [FH19, HO16]. Aggressive [MT19]. agis [Fox10]. Agree [Che11]. Aid [WBNTF19]. AIDS [Car19]. Airbnb [BCG18]. Aldor [ANBB^{+}14]. Aldor-Noiman [ANBB^{+}14]. Algorithm [LP14]. Algorithms [LBR17, LDM10]. Aliases [HV18]. Allocation [WBNTF19, Wri12]. Also [Coh17, GR13a]. Alternating [Pat18]. Alternative [CJE19, Hay11, Hon12, Hon15b, KO12]. Alternatives [MLL10, RB18]. Am [KV18]. American [ANBB^{+}14, BXHH14, Din14a, Fou20, Fri15, Hol14, IDE15, IIDdL14, MS14, Muk14a, Muk14b, MM15]. Analogies [BGMA13, LAK16]. Analogy [Sti17]. Analyser [BJ16]. Analyses [FBHW19, MLL10, MPP19, TG18, WW11, WASB10]. Analysis [All19, AD18, Arm14b, BSC15, DKY17, GR19, KO12, KS15, Le 19, Lu19, Mat19, NL15, SF19, SVCB18, VR15, Wag16, Wan19, WYG17, YGP14, ZPL^{+}18]. Analytic [SVR11]. Analytical [MBM18]. Analytics [LHH15, PS17]. Analyzing [FGTM14]. André [Fd12]. André-Michel [Fd12]. Animation [Sti17]. Anniversary [CSBB^{+}15, Rod15]. ANOVA [Har19, LLS16, WGW12]. Anti [GR13b, TCC^{+}15]. Anti-Bayesian [GR13b]. Anti-Statistical [TCC^{+}15]. Antipoverty [YGP14]. Any [Dem19]. AP [FHP^{+}11]. Application [GE11, KKAB16, NR19]. Applications [HHC17, Lit19, Lom13, LGXB11, LPB13a, ZNY11]. Applied [FBHW19, GR13a, MGCR10, May13, Pog19, Sch10a, Wan10, YJCR15]. Applying [MGRL14]. Apportioning [Wri12]. Appraise [HRS10]. Approach [And15, BDP18, BL13, CFP18, Har14a, LHH15, Lu19, LLS16, NL15, dSPCdS17, Por16, PFG16, Ric10, Sam13, SD14]. Approval [RHGS^{+}19]. Approximate [And15, DH11, JTL17]. Approximately [Coh17]. Approximating [HS12]. Approximation [Sti13b]. Approximations [OW10]. Arab [Bro11]. Arbitrary [Bac19]. Arising [MG14]. Arm [IDE15, KF14]. Articulating [Gri13, Muk14b]. ASA [CSBB^{+}15, Coc15, Hub19, IGRP17, Kwa17, Mor17, Rod15, Was15, WL16b]. Ask [CJC19]. Assess [GP13, Muk14a, Pog19, WE13]. Assessing [And19, FBHW19, LLB16, NMR16]. Assessment [CP15, EIRS17, GZK^{+}11, MK15]. Assignments [NP16]. Associated [Fie13b]. Association [Arm16, LLB16, Smi11, MM15]. Association-Not-Causation [Smi11]. Asymptotic [Coh17, GD19, KOT12]. Attendance [Wie16]. Attributes [McN19]. Attribution [PFG16]. Audit [MHWB19]. Auditing [TdCP17]. Authentic [Gri15, NL15]. Authorship [PFG16]. Autocorrelated [LLS16]. Average [Hod17, KKAB16, Kit17]. Awards [ZC15].

B [BXHH14, WS15]. B-Spline [WS15]. Back [Coc15, Gri13, Hor15, Muk14b, BH11]. Back-Transformations [Gri13]. Bad [Ton19, VR15]. Balance [GE11, HRS10, MPP19]. Balanced [NR19]. Ban [FBHW19]. Bar [Chr17, PB16]. Bar-Gera [Chr17]. Barley [Wri13]. Baseball [FF18, SGBS17]. Based [BLOP19, CW11, DDF19, Har14a, KOT12, Kou12, KS15, Man13, Mar17, MLL10, ML14, Sam13, Sch12, TL12, TCC^{+}15, VR15, VTH14, WRH19, XQL16, YH15, YB10]. Basic [FBHW19]. Basis [LaM14]. Basketball [ACF16, SEC^{+}19, WB18]. Baye [Arm16, RM19b]. Bayes [GM16, GJLW19, Har19, Lav19, LLB16, Rou19, Wit17, Zig16]. Bayesian [BL13, FF18, Fie13a, GdBPP19, GR13b, GR13a, GGGV19, GG15, Gre10, Joh13, Le 19, LNM16, Lon10, May13, Sam13, SVCB18, Sti13a, WL16a, WIE13, WGW12, WYG17, ZR10, vDLMW18]. Bayesianity [JTL17]. Be [JJ14, SJD10, All19, Hub19, Muk14a, Rob19, WE13]. Been [Sam15]. Before [KS19]. Behave [Gre19]. Being [Fri15, LSO01]. Bender [Sta16]. Benefit [Pog19]. Benford [AP10, BH11, BS10b, Hil11]. Benjamin [Fou20]. Berger [Fou20]. Berserk [CG19]. Best [PBD^{+}16]. Beta [HV18]. Better [CJC19, Gri13, LAK16, Men10, Muk14b, SJD10]. Between [Bau18, BM10, Dem16b, DHW11, FHP^{+}11, LL13, LDM10, MT19, Sen12, WJ19]. Between-Group [LDM10]. Beyond [BDP18, HG12, KS19, PR10, Smi11, SB15, SLG19, WSL19]. Bias [LRR19, MPD19, PB16, XY16, YH15]. Big [LBR17, Sta16, VLZ17, ACF16]. Binary [IDE12, IDE15, KF14, MS14]. Binomial [Dem18, Eas15, FP12, JTL17, KMS10, LTN10, Lon10, Pat18, SD14, Sti13b, Vel15]. Binseqtest [IDE15, KF14]. Biostatistical [PBD^{+}16]. Biostatisticians [Sch10a]. Biostatistics [BU15, BV11, DWG15, LAK16, SFSM17, Wal18, WYG17, ZC15]. Birthday [HHC17, Sch10b]. Bivariate [AM14, Dem19, FR12, LC15, ZBGZ14]. Blackwell [GM16]. Blending [GdBPP19]. Blind [Loc19]. Block [NR19]. Blood [DWG15]. Blunders [Hil11]. Board [Ano12a, Ano14b, Ano15b, Ano16b, Ano17a]. Boca [Sab19]. Book [Ano15a, Ber19, Car19, Dre19, Hig19, Lit19, Lu19, Sab19, SF19, Sta16, Wan19]. Books [Ano10f, Ano10g, Ano10h, Ano11h, Ano11e, Ano11f, Ano11g, Ano12f, Ano12e, Ano12g, Ano12h, Ano13e, Ano13f, Ano13g, Ano13h, Ano14f, Ano14d, Ano14e, Ano15a, Ano15f, Ano15d, Ano15e, Ano16d, Ano16e, Ano16f, Ano16g, Ano17e, Ano17c, Ano17d, Ano18d, Ano18b, Ano18c, Ram17, SB14]. Bootstrap [Hes15, MGRL14]. Bootstrapping [Ang16]. Both [GA13]. Boundedness [FPW12]. Bounds [DH11, Fro12, LNM16, Por15]. Bowling [KO12]. Boxplot [WDCH18]. Bridges [FHP^{+}11, War17]. Bridging [ZSMS17]. Brief [EB18, MM15]. BUGS [Lu19]. Building [FHP^{+}11, RM19a, War17]. Bull [HLU19]. Bulut [Sab19]. Business [CW10, PS17, Wan19]. Buy [Dem16a].

C [MS14, EB18]. Calculating [FT10, LDM10]. Calculation [LP14, ML16, SLJ^{+}18]. Calculations [SLG19]. Calculator [Fd12]. Calibrated [LLB16]. Calibrating [Old16]. Cambridge [Sta16]. Can [AP10, HR10, JJ14, VM13, BXHH14]. Can’t [Dem16a]. Capstone [LRF11, SB15]. Care [JS11]. Career [ACHMW17, LVH^{+}17]. Case [HEJ11, JM18, Kha15, LR16, Pog19, RLS^{+}16, Wag16, WASB10, YH15, ZRK^{+}11]. Case-Control [ZRK^{+}11]. Cases [KOT12]. Categorical [ALR17, BR14, MH18, XY16]. Causal [AD18, Arm14b, Zig16]. Causation [Fri15, Smi11]. Causes [YH15]. Cautionary [HV18]. Celebrate [Rod15]. Celebrating [CSBB^{+}15]. Censored [Lu19]. Center [SFSM17]. Centering [Chr18a, Vel18a]. Centers [PBD^{+}16, Van15]. Central [Inl10, Kon11, NL18, Pat18, Zig16, BBM13]. Century [Joh13, RHGS^{+}19]. Certain [Vel15]. Certainty [LG19]. Chains [AKLS10]. Challenge [Bor14]. Challenges [Gol17, Hor15, KKM15, YJCR15]. Change [Gor16]. Changes [QL16, Tra19]. Changing [Hig19, Sti10]. Chapman [Sab19]. Characteristics [MT11]. Characterizing [DDV16]. Cheap [FR19]. Chebyshev [Coh15, SVG17]. Chi [Che11]. Chi-Square [Che11]. Choi [IIDdL14]. Choice [Man19, MT19]. Choosing [FT10]. Christensen [Din14a]. Christopher [Sab19]. Citation [AD18, Sch10a]. Class [Bor14, JM18, Wie16]. Classes [LC15, ZHM^{+}13]. Classical [GdBPP19]. Classification [CW11, GJLW19, ZLW13]. Classification-Based [CW11]. Classifier [BC12]. Classifiers [BXHH14, VM13]. Classroom [And15, CFP18, GZ17, WRH19]. Clean [MPP19]. Clinical [Pog19, Dre19]. Clinicians [LAK16]. Clopper [Eas15]. Closed [GP13, IDE12, LRR19, MS14, Wan10, YC17]. Closed-Form [LRR19, YC17]. CLT [AP10]. Clumping [KM19]. Cluster [LG19]. Cluster-Membership [LG19]. Clustering [CCB16, LG19, ZZT^{+}19]. Cochran [KV18]. Code [VR15]. Coefficient [Dem16b, OR14, Zha17, vDLMW18]. Coefficients [Dem18, Din14a, Pat18, SDE19, TOPC14]. Coherent [FIME17]. Cohort [WASB10]. Coincidences [Sch10b]. Collaborating [SFSM17]. Collaboration [DDF19, EL18, Loc19, Mic14, PBD^{+}16]. Collaborators [Ano10a, Ano11a, Ano12b, Ano13a, Ano14c, Ano15c, Ano16c, Ano17b, Ano19b, Ano19c]. College [FHP^{+}11]. Collide [Hal10]. Collinearity [Chr18a, GG15, Vel18a]. Combating [TCC^{+}15]. Come [HS17]. Comment [ANBB^{+}14, BXHH14, BB14, Chr14, Chr17, Chr18b, Chr18a, Chr20, Din14a, Fie13a, Fie13b, Fou20, Fre19, Fri15, Hol14, Hut18, Joh13, Kad13, LM14, McC14, MS14, Muk14a, Muk14b, Pea14, Sch13, Ste14, Sti13a, Zim14, VR15]. Commentary [Mor17]. Common [AM14, DHW11, FHMV19, Utt15]. Communicating [Was15]. Communication [DDF19]. Community [Sch12]. Community-Based [Sch12]. Comparative [Ric19]. Compare [RS13, AP10]. Comparing [FWW17, GJLW19, ME11, Sch10b, WKW19]. Comparison [ALR17, BHLE15, Fro12, GJLW19, HL10, LDM10, WB17]. Comparisons [MA19, YH15]. Competitions [NR19]. Complementary [Gui18]. Complicated [Man13]. Component [LDM10]. Computation [Jae16, LFS^{+}17, Nad15, NL18]. Computational [LHH15]. Compute [LNM16]. Computer [HKPT18, PW13]. Computing [ÇRR18, DH11, McN19, NL10, SLJ^{+}18]. Concentration [EIRS17]. Concepts [BGMA13]. Conceptual [GB15, WRH19]. Concerning [Col19]. Concerns [HEJ11]. Conclusions [YZS^{+}14]. Conditional [Din16]. Conditionality [Pri11]. Conference [ACF16]. Confidence [And15, Eas15, Fre10, FP12, FZ17, GE11, GA13, Hay14, HL10, HKM16, Hut18, Jae16, JTL17, Lan17, LTN10, LDM10, Man13, NMR16, PB16, Por19, SD14, VZH16, Wan18, WRH19, WKW19, Zha18, Zuo10]. Confounding [TK18]. Consequences [CG19]. Conservative [FR19]. Consider [Tra19]. Consistency [YB10]. Constants [Gre10]. Constrained [TG18]. Construction [And15, AM14, LaM14, LTN10, Wan18]. Consultation [Loc19, PBD^{+}16]. Consulting [Kha15, SB15, Van15, Zab13]. Contemporary [MTBG19]. Content [MHWB19]. Contest [Wie16]. Contests [PGCL14]. Context [Bet19, Gor11, IGRP17, MCM^{+}11, WL16b]. Contextualize [KS19]. Contingency [Gre10, KV18, SF19]. Continuous [Sch10b]. Contoured [ZBGZ14]. Contract [FKN^{+}17]. Contrary [Hub19]. Contrast [AP10]. Contributions [Dav11, Wal18]. Control [Bry18, CLR13, HC19, LC09, RS13, WYG17, ZRK^{+}11]. Controversial [KW19]. Controversies [FGTM14]. Convergence [BM17, DHW11]. Convolution [Pat18]. Cooperate [Hal10]. Coordinate [BR14]. Correcting [Kme19]. Correction [Ano10b, Ano14a, Ano16a, Ano19a, CLR13]. Correlated [Coh19, Fri15, GD19, HR10, LSO01, ZZT^{+}19]. Correlation [DH11, Dem16b, Dem19, FR12, Fri15, Muk10, WB17, ZBGZ14, vDLMW18]. Correlations [GS14]. Correspondence [ZPL^{+}18]. Corrigenda [Ano18a]. Corrigendum [O’N16]. Corrupt [Kme19]. Could [SJD10]. Count [KZ16, ZSMS17, Mac17, SMSZ17]. Counter [BBM13]. Counterexamples [Muk10]. Counts [Tuy19, Wan10]. Coup [HLU19]. Course [Bau15, BSC15, CFP18, DDF19, Hor13, LRF11, MGCR10, Mar17, SLG19, Wag16]. Courses [GZK^{+}11, Gri15, Kha15, Sch12, WYG17]. Covariance [GKP12, GP13, Hay11, Muk10]. Covariate [HRS10]. Coverage [SD14]. Cox [Hal10, KOT12]. CRC [Sab19]. Creating [VTK^{+}17]. Creativity [Wie16]. Credibility [Mat19]. Credible [LNM16]. Crime [Por16]. Crisis [ATG19, Rob19]. Criterion [GSK19]. Critical [BH12, CFP18]. Criticisms [Gre19]. Critique [Arm14b]. Cross [ESC18, HRS10, MCM^{+}11]. Cross-Cultural [ESC18]. Cross-Match [HRS10]. Cross-Validation [MCM^{+}11]. Crowdsourcing [EIS17]. Cryptology [Bro11]. Cultural [ESC18]. Current [PS17, WG11]. Curricula [HHH^{+}15, NL10]. Curriculum [ÇRR18, CP15, Cob15, DDF19, ESC18, Hes15, HH15, TCC^{+}15]. Curve [Gri13, HS12, LCG16, Muk14b, TL12]. Cutoff [Bac19].

D [ANBB^{+}14, BXHH14, Fou20, MS14, Muk14b, Sab19]. Daily [SEC^{+}19]. Dangerfield [Sti13a]. Dangers [SSS12]. Data [ALR17, Bau15, Bau18, Ber19, BCG18, BS10b, BHLE15, BW18, BR14, CCB16, Dem19, EL18, FHMV19, FR12, GKP12, GR19, Gön13, Gri15, HHH^{+}15, HI18, HH15, Kap18, KZ16, KS15, LBR17, Lu19, LLS16, LCG16, MPD19, Man19, MBM18, MH18, MPP19, Nie11, NL15, NP16, PP10, Sch10a, SJD10, Sta16, TL12, TdCP17, VR15, VZH16, Wag16, WDCH18, Wal18, WBL18, Wri13, XQL16, YB10, ZHM^{+}13, ZSMS17, ZPL^{+}18, Mac17, SMSZ17]. Data-Driven [VZH16]. Dataset [SVCB18]. Day [Sch10b]. Dealing [TK18]. Decision [BLOP19, EIS17, Lon10, Man19, MKR19, Ric10, RHGS^{+}19]. Decision-Making [BLOP19]. Decision-Theoretic [Ric10]. Decorated [BR14]. Decorrelation [KLS18]. Deeper [Men10]. Default [WGW12]. Defended [GR13a]. Defense [May13]. Define [GdBPP19]. Defining [VKP^{+}17]. Delta [HL10, Hoe12]. Demonstrated [Sam15]. Demonstration [RM19a]. Demystifying [Lo19]. Densities [ZBGZ14]. Density [HS12, VTH14]. Density-Based [VTH14]. Departments [Wal18]. Depend [Coh19]. Dependence [DHW11]. Dependent [GdBPP19, WASB10]. Depends [HO16]. Derivations [Hon14b]. Derivatives [Man13]. Derived [LRR19, YC17]. Descriptive [ATG19]. Design [BG17, BSC15, Chr17, KRSR12, MPP19]. Designs [Grö14, NR19, WGW12, XQL16]. Desire [CW10]. Desired [Fox10, GA18, vC10]. Desjardins [Sab19]. Detecting [MG14]. Determination [Zha17]. Determining [Che11]. Deutsch [MS14]. Developing [CFP18, LHH15, LVH^{+}17, YJCR15]. Development [Gol17, RHGS^{+}19]. Developments [Van15]. Deviation [Muk14a, SR16, WE13]. Diagnostics [Chr18a, Vel18a]. Dialogue [BM10]. Dice [PP10, Sun18]. Did [ER15]. Difference [Hon15a, KMS10]. Differences [JM18]. Digital [Sti13b]. Dimensional [GD19, HZ17, Jae16]. Direct [LP14, ML16]. Directions [KD17]. Directors [Coc15, Was15]. Dirichlet [DDV16]. Discard [SJD10]. Discipline [Kot10].

Discontinuation [KRSR12]. Discrepancies [WJ19]. Discrete [Fro12, LC15, Sch10b]. Discreteness [YH15]. Discussion [May13]. Disease [Wan10, WWGL19]. Disparate [MG14]. Display [Mis19]. Displaying [Ber19]. Dispute [FKN^{+}17]. Dissonant [YZS^{+}14]. Distance [WYG17]. Distribution [ALL10, Din14b, Din16, DB18, HEJ11, Hay14, LRR19, QL16, YC17, YB10, ZPM^{+}11]. Distributional [JMS19, MT11]. Distributions [BM17, Dem16b, HKM16, HV18, LC15, ZBGZ14, Zha18]. Diversity [HO17]. Divided [Muk14a, WE13]. Do [Bry18, Che11, Col19, FZ17, Hor13, Hut18, KV18, SVCB18, vC10, dC16]. Doctor [BU15]. Documenting [Wal18]. DOE [FGTM14]. Does [Hod17, Hol14, Kit17, ST13]. Dollar [FKN^{+}17]. Don’t [ATG19]. Doolittle [Arm16]. Draws [HMR13, LNM16]. Driven [Eas10, VZH16]. Dropout [ZPL^{+}18]. Dropouts [Gön13]. Drug [ER15, RHGS^{+}19]. Duration [FHMV19]. Dutter [Chr18b].

Early [ACHMW17, Bro11, SFSM17]. Early-Stage [SFSM17]. East [ACF16]. Easy [CLR13, LC09]. Ecology [Coh16]. ed. [Ber19]. Editor [Hil11, Ost11d, Ost11a, Ost11b, Ost11c, UL15, Ano10d, Ano10e, Ano12c, Ano12d, Ano13d, Ano13b, Ano13c, Coh10, CFR15, Hod17, HN10, Hon14a, Jon19, Les16, Mis15, Nad16, Ost12a, Ost12b, Ost13a, Ost13b, Ost14, Smi12, Sti16, TP19]. Editorial [Ano10a, Ano11a, Ano12a, Ano12b, Ano13a, Ano14b, Ano14c, Ano15b, Ano15c, Ano16c, Ano17a, Ano17b, Ano19b, Ano19c, Tra19, Ano16b]. Editors [Ano11b, Ano11c, Ano11d, Tra19]. eds [Sta16]. Education [FHP^{+}11, Hor15, PS17, RST15, Utt15]. Educational [Sab19, YH15]. Edwards [MS14]. Effect [Bro10, GSK19, HR10, Pog19, ZLW13]. Effective [PS18]. Effects [ALL10, CGM19, GR19, IIDdL14, KCE13, KD17, LL13, LDM10, SVR11, Zig16]. Efficiency [WBNTF19]. Efficient [CW11, LFS^{+}17, SLJ^{+}18]. Efforts [Hub19]. Eight [Nar12, VTK^{+}17]. Eight-Step [VTK^{+}17]. Elementary [BM17, Hon14b, VWB^{+}10]. Elicitation [JJ14, O’H19]. Else [CJC19]. Elteren [MLL10]. Emery [IIDdL14]. Emphasis [ESC18, Har14a]. Empirical [ALR17, GSK19, MCM^{+}11, VTH14]. Enables [Ton19]. Encouraging [Wie16]. Engagement [Sta16]. Engineering [GRTM16, HS17, TG18]. Enough [Zil19]. Enriching [WRH19]. Ensemble [KSW16]. Entropy [Hod17, KKAB16, Kit17]. Entwined [RS13]. EOV [Ano12a, Ano14b, Ano15b, Ano16b, Ano17a]. Equal [Wri12]. Equality [NMR16, ZXC12]. Equation [Nar12, TK18]. Equations [LRR19, WB17, YC17]. Equivalence [CLR13, FHMV19, LC09, Wri12]. Era [BV11, LBR17, Mat19]. Erich [Muk14a]. Error [BXHH14, Hol14, Kou12, RLS^{+}16, SEGMA18, ST13, Sto12, VWB^{+}10, VM13]. Errors [GD19, Gor11, HR10]. Estimated [Ang16, SVG17]. Estimates [KD17, LP14, SCLP17, YB10]. Estimating [Din14a, FR12, GA18, LCG16, Pat19, PP10, TOPC14, WK13, WB17]. Estimation [AKLS10, DF16, Fie13b, HEJ11, IIDdL14, KRSR12, KCE13, KD17, LPB13a, PW13, Roc16, VZH16, ZLW13, Zig16]. Estimator [GM16, GD19, Kou12, ML16]. Estimators [HS12, LL13, LRR19, YC17]. Ethical [HG12, LMZ15]. Ethics [ESC18, LMZ15]. Euler [Dav11]. Evaluating [DMM19, LR16, RLS^{+}16, Tra19, Wal18]. Evaluation [KOT12]. Even [ML16]. Event [PW13]. Evidence [BJ16, BLOP19, GSK19, HO16, Hub19, Joh19, RM19b, VR15]. Evidence-Based [BLOP19, VR15]. Exact [Che11, DKY17, FP12, Lan17, ZXC12]. Exactly [Gre19]. Exam [BM10]. Examination [Lom13, WJ19]. Example [BS10b, GM16, Gre10, LDM10, WIE13]. Examples [BBM13, LCD10, Lu19]. Excel [dRFF17]. Excuse [Bry18]. Executive [Coc15, Was15]. Expanding [RST15]. Expansion [FWH^{+}13]. Expect [ATG19]. Expectation [CJE19, Hon12, Hon14b, Hon15b, Kon11, Lo19]. Expectations [Hon15a]. Expected [KV18, Rob19]. Experience [GZ17, Kot10]. Experiences [Gri15]. Experimental [MPP19]. Experiments [BG17, BSC15, Chr17, MLL10]. Expert [BLOP19, GS18, O’H19]. Expires [HLU19]. Explaining [BGMA13, GRTM16, OW10, Sti17]. Explanation [AP10, BH11]. Explorations [NL15]. Exploring [FHMV19, Gol17]. Exponential [DF16, HKM16, Zha18]. Expose [NL15]. Exposure [WASB10]. Expression [NL18]. Extended [PL17]. Extending [EB18]. Exterior [RS13]. Extreme [Che11]. Eyes [LFH16].

Face [PP10]. Facets [Utt15]. Facilitating [SLJ^{+}18]. Factor [Har19]. Factorial [Fro12, Grö14, SVR11]. Factorizations [FHB^{+}13]. Factors [GJLW19, LLB16, Rou19]. Faculty [BM10, Gol17, SFSM17, ZC15]. Fails [DNBJC18]. Fairness [KO12, RM19a]. False [Col19, Sam15]. Families [HV18]. Family [HKM16]. Fantasy [SEC^{+}19]. Fast [NL18]. Fay [IDE15]. Fear [CW10]. Feared [Fox10, vC10]. Feasible [DDF19, Dem19]. Few [BBM13, MT11]. Fewer [SVR11]. Fieller [HL10]. Figure [EA11]. Filling [Sti17, XQL16]. Filter [KSW16]. Finding [SD14]. Findings [Pog19, Sam15]. Finite [Dem18]. Finiteness [Muk10]. First [FVMU10, Fd12, HG12, Joh13]. Fisher [Che11, Hal10, Le 19, Pri11, Ric10]. Fitting [WG11]. Five [BGMA13, MKR19, Tra19]. Five-Decision [MKR19]. Fixed [BG17, CGM19, Chr17, Fre10, HR10]. Fixed-Design [BG17, Chr17]. Fixed-Effects [CGM19]. Fixed-Width [Fre10]. FL [Sab19]. Florent [FVMU10]. Flux [May13]. Following [ZRK^{+}11]. Force [SB15]. Forecasting [Bor14, TL18]. Foreign [WBNTF19]. Forget [Hor13]. Form [LRR19, Wan10, YC17]. Formal [HHP19]. Formula [CJE19, Hay11, Hon12, Hon15b, Lo19]. Formulas [KOT12]. Formulation [Ric10]. Forward [HS10, Hor15]. Fostering [GB15]. Foulley [Fou20]. Framework [BHLE15, Gri15]. Frameworks [Sta16]. Framing [GR19]. Fraud [EIRS17]. Free [KRSR12]. Frequency [QL16]. Frequentist [JTL17, Lav19]. Frey [Hut18]. Friedman [Fri15]. Friends [MBM18]. Fruitful [LM14]. Fun [JJ14]. Function [AM14, CJE19, DF16, EIRS17, Fra19, Inl10]. Functions [Man13, Nad15, Pat19]. Future [BU15, Kot10, Van15]. Fuzzy [AP10].

GAISE [DCF^{+}17]. Galton [Gor16]. Game [KS15, PL17, TO16]. Game-Based [KS15]. Gamma [LRR19, YC17]. Gap [ZSMS17]. Gauss [Nad15, dSPCdS17]. Gaussian [DF16, DB18]. gem [BH11]. General [BS10b, LC15, ZPM^{+}11]. Generalized [FPW12, GM16, WS15, WB17, Zha17, ZSMS17]. Generating [Inl10]. Generation [BGW^{+}19, HH15, Lyn16, Rod15]. Generative [WWGL19]. Genetic [LLB16]. Genomic [Sam15]. Genomics [MCM^{+}11]. Geometric [HL10, Hon14b, dSPCdS17, Xu14]. Geometrical [Hay11, ML14]. Gera [Chr17]. Gertrude [Hal10]. Get [Cun12]. Getting [Goo19]. Giants [Sne19]. Gibbs [BL13]. Gini [LCG16, TL12]. Goalie [BS10a]. Gone [CG19]. Good [Sta16, Ton19]. Google [CMW15]. Google-Scale [CMW15]. Gosset [Zil19]. Grâce [HLU19]. Graduate [HO17, LMZ15, WYG17]. Gram [LaM14]. Grammatical [Hol14, ST13]. Grand [LL13]. Graph [FVMU10]. Graphical [ANBB^{+}13, ANBB^{+}14, ML14, SDE19]. Graphics [JJ14]. Greater [CG19, YGP14]. Greece [Mic14]. Griffith [Muk14b]. Ground [Cob15]. Group [LDM10, LCG16, MA15, MA19, ZNY11]. Grouped [LCG16, TL12]. Groups [RS13]. Growing [Rod15]. Guarantee [FZ17, Hut18]. Guerry [Fd12]. Guide [HI18, VTK^{+}17]. Guidelines [CP15, HG12]. Guinnessometrics [Zil19]. Guns [CCFJ18].

H [Arm16, Sab19, TO16]. H-O-R-S-E [TO16]. Hall [Sab19]. Hall/CRC [Sab19]. Hand [BD11, Sto12]. Handbook [Sab19]. Handicapping [KO12]. Handle [Tuy19]. Hard [Goo19, LG19]. Hartford [YGP14]. Hazard [Bed14]. Head [Lyn16]. Heads [Was15]. Health [JS11, PBD^{+}16, RM19a, SFSM17]. Health-Care [JS11]. Hear [Hor13]. Heart [DWG15, Lyn16]. Helen [Sta16]. Helpful [dRFF17]. Helping [LAK16]. Helps [BCG18]. Heterogeneity [PGA12]. Hidden [SSS12]. Hierarchical [ALL10, JS11]. High [FHP^{+}11, HZ17, ZPL^{+}18]. High-Dimensional [HZ17]. Higher [CJE19, PS17]. Higher-Order [CJE19]. Hilbert [NSSLC12]. Histogram [HS12, TL12]. Histogram-Based [TL12]. Historical [Gor11, SVCB18]. History [Gui18, KS19, MM15, Sti10]. HIV [Car19]. HIV/AIDS [Car19]. Hoc [LR16]. Hockey [BS10a]. Holds [Dem18]. Home [FF18]. Homogeneous [LP16]. Hot [Sto12]. Hours [Sch10b]. House [Wri12]. Human [EA11, Fie13b, LPB13a]. Humans [ZHM^{+}13]. Hybrid [GSK19]. Hyperbolic [Din14b]. Hyperbolic-Secant [Din14b]. Hypercubically [ZBGZ14]. Hypergeometric [Nad15]. Hypersphere [Le 19]. Hypotheses [HO16]. Hypothesis [BHLE15, FR19, FIME17, Man19, Par19, WL16a, WGW12].

ICAR [LH12]. Idea [HS17]. Ideas [Fie13a, NP16]. Identification [JS11]. Identifying [Sch10a]. Identities [Vel15]. Identity [JMS19, LL13]. if [ATG19]. Ignatova [MS14]. II [MT19, SEGMA18]. Image [Coh15, KKAB16]. Immer [Wri13]. Impact [Loc19, MG14, Was15]. Implications [Van15]. Importance [GRTM16, Kot10]. Improper [TL10]. Improvable [GM16]. Improve [FR19, Hub19]. Improved [WDCH18, WSF10]. Improvement [GM16]. Improving [ACF16, BB19, Fou20, Wie16]. Imputation [ALR17, HMR13, WJ19, XY16, ZR10]. Include [BU15]. Including [WKW19]. Incomplete [GKP12, NR19]. Incorporating [Kha15, VZH16]. Increase [BXHH14, ER15, VM13]. Increased [WSF10]. Increasing [Gib19, Nie11]. Independence [BDP18, Muk10, VTH14]. Independent [BXHH14, HMR13, VM13]. Indeterminacy [ALL10]. Index [Ano10c, Ano19d, LCG16, TL12]. Indices [QL16]. Individual [LG19, PGA12]. Inducing [Dem19]. Industry [Sne19]. Inefficient [GM16]. Inequality [Coh15, LB19, PS18, SVG17]. Infer [MKR19]. Infer.NET [WW11]. Inference [AD18, BL13, Bil19, Bro11, BLOP19, CGM19, Fra19, GR13a, HHP19, IDE12, KM19, LM14, LCD10, Mar17, MS14, Rob19, RHGS^{+}19, Sti13a, Sun18, Ton19, ZR10, vDvDG^{+}19, vDLMW18]. Inferences [IDE15, KF14, LR16, Rob19]. Inferential [ATG19]. Influence [DWG15]. Inform [TO16]. Information [VZH16]. Informative [JJ14]. Infrastructure [ÇRR18]. Infusing [Gri15]. Initial [Lom13]. Innovations [Fie13b]. Innovative [CFP18]. Institute [DWG15]. Institutions [PS17]. Instrumental [YZS^{+}14]. Integer [Dem18]. Integrated [Lo19]. Intensified [BM10]. Interaction [PGA12]. Interactions [HZ17]. Interactive [JJ14, WRH19]. Interdisciplinary [BV11, LRZG^{+}10, LHH15]. Interpolated [FZ17, Hut18]. Interpolation [TL12]. Interpretation [Hay11]. Interpretations [ML14]. Interpreting [SDE19]. Interquartile [Muk14a, WE13]. Interrater [QL16]. Interrelation [Sen12]. Interval [And15, Lu19, Por19, Roc16, VZH16, Wan18, Zuo10]. Interval-Censored [Lu19]. Intervals [DF16, Fre10, FP12, FZ17, GA13, Hay14, HL10, HKM16, Hut18, JTL17, Lan17, LTN10, Man13, NMR16, PB16, Wan10, WRH19, WKW19]. Interventions [ER15]. Intriguing [JMS19]. Intro [GA13]. Introduced [XY16]. Introduction [BGW^{+}19, EB18, IDE15, KF14, Wan19, WBL18]. Introductory [BU15, CFP18, DCF^{+}17, Gri13, Kha15, MGCR10, MHWB19, Muk14b, Sol10]. Intuitive [dSPCdS17]. Invariance [Por19]. Invariant [JRP11, KD17]. Invented [Hoe12]. Inverse [DF16, HHC17]. Investigations [GE11]. Investigators [LAK16]. iPad [Lom13]. ISBN [Sab19, Sta16]. Isolated [Bau18]. Issue [HH15, WBL18, WKW19]. Issues [KS15].

J [Fou20, Hol14, IDE15]. Jensen [LB19]. Joint [ZZT^{+}19, Zig16]. Joke [Wie16]. Judgment [BLOP19]. Julia [Sta16]. Just [MT11].

Kalman [KSW16]. Kendall [vDLMW18]. Kernel [NSSLC12]. Key [McN19, Sch10a]. Kim [IIDdL14]. Kirk [IDE15]. Knaeble [Chr18b]. Know [BU15, BV11, DMM19, Hes15]. Knowledge [O’H19]. Known [BM17, CJC19, FR12, FVMU10]. Kurtosis [JRP11, SB11, Wes14].

L [ANBB^{+}14, Hol14, IDE15, Muk14a]. Landscape [AD18, PS17]. Lane [Sta16]. Langford [Fri15]. Langren [FVMU10]. Laplace [DB18]. Large [MTBG19, Zil19]. Large-Scale [MTBG19]. Late [Cob15]. Latin [FKN^{+}17]. Law [Coh16, Coh17, Dem18, Hon15a, BH11, AP10, Hil11]. Leadership [Gib19, Gol17]. Leaf [Mis19]. League [FF18, SGBS17]. Learned [YJCR15]. Learning [CP15, DMM19, GZK^{+}11, NP16, Sch12, WYG17]. Learnt [Ioa19]. Least [GD19, KD17]. Least-Square [GD19, KD17]. Legal [Zab13]. Lens [ZPL^{+}18]. Lessons [Bau18, YJCR15]. Letter [Ano12c, Ano12d, Ano13b, Ano13c, Coh10, CFR15, Hil11, Hod17, HN10, Hon14a, Jon19, Les16, Mis15, Nad16, Smi12, Sti16, TP19]. Letters [Ano10d, Ano10e, Ano11b, Ano11c, Ano11d, Ano13d]. Level [CP15, Dem19, HS10, JM18]. Levels [GdBPP19, ME11]. Leveraging [EA11]. Lévy [Inl10]. Li [vC10]. life [JM18]. Lifetime [Wan19]. Likelihood [DF16, GM16, GP13, HEJ11, Jae16, LP14, LWM15, LRR19, LCD10, ML16, ML14, PL17, VTH14, YC17, YB10, ZXC12]. Likelihood-Based [ML14]. Limit [BBM13, Inl10]. Limitations [RM19a]. Limited [HHP19]. Limiting [BM17]. Limits [Eas15, GE11, LDM10]. Lindeberg [Inl10]. Linear [CGM19, GD19, HZ17, LaM14, WS15, WG11, ZPM^{+}11, Zha17]. Lines [Bau18]. Linkage [Por16]. Links [DHW11]. Little [Cob15, Zil19]. Live [SVCB18]. Located [YGP14]. Location [DKY17, HS12, SB11]. Location-Scale [DKY17]. Logistic [IIDdL14, KCE13, MPD19, OR14]. Logit [SDE19]. Long [HL15, SVCB18]. Longitude [FVMU10]. Longitudinal [BR14, GKP12, Gön13, ZPL^{+}18]. Look [Coc15, WIE13]. Looking [Hor15]. Lorenz [LCG16, TL12]. Love [HR10]. Low [Grö14, ME11]. Low-Order [Grö14]. Lower [DH11]. Lung [DWG15].

M [BXHH14, Fri15, IDE15]. M. [Arm16]. Madness [FT10]. Magnitude [And19]. Major [FF18, MK15, SGBS17]. Majority [BXHH14, VM13]. Majors [LRF11]. Make [BCG18, Hol14, ST13]. Making [BLOP19, Lon10, RHGS^{+}19]. Malawi [WBNTF19]. Mann [DNBJC18, FH19]. Many [KV18, Utt15]. Mapping [WWGL19]. March [FT10]. Marginal [AM14, GD19, GS14]. Marginally [Joh19, ZBGZ14]. Marginals [Dem19]. Markov [dSPCdS17, AKLS10, Coh15, TO16]. Maryland [BH12]. Master [JM18]. Match [HRS10, RS13]. Matched [HRS10, RS13, WYG17]. Matching [DKY17, LGXB11, Por15, ZRK^{+}11]. Materials [Ano10f, Ano10g, Ano10h, Ano11h, Ano11e, Ano11f, Ano11g, Ano12f, Ano12e, Ano12g, Ano12h, Ano13e, Ano13g, Ano13h, Ano14f, Ano14d, Ano14e, Ano15a, Ano15f, Ano15d, Ano15e, Ano16d, Ano16e, Ano16f, Ano16g, Ano17e, Ano17c, Ano17d, Ano18d, Ano18b, Ano18c, Ram17, SB14, Ano13f]. Mathematical [GB15, Gri13, Hor13, Muk14b, BH11]. Mathematics [Dav11]. Matrices [GP13]. Matrix [FHB^{+}13]. Matters [Muk10]. Maximal [HO16]. Maximizing [Por15]. Maximum [GM16, HEJ11, LP14, ML16, YB10]. MCMC [LNM16, Wit17]. Me [Bry18]. Mean [And15, Coh17, GKP12, Kou12, Lan17, LL13, Pat19, Por19, RLS^{+}16, SR16, Sen12, SVG17, Xu14, dC16, FWH^{+}13]. Mean-Covariance [GKP12]. Means [CCB16, DF16, GS14, LCG16, NMR16, SCLP17]. Measure [Hod17, KKAB16, Kit17, PS18]. Measurement [Sab19, Sto12, VWB^{+}10]. Measures [Arm16, JRP11, LG19]. Mechanism [MPD19]. Median [SR16]. Medians [DNBJC18]. Medical [EIRS17, ZC15]. Medicare [GE11]. Meeting [Wag16]. Meier [Zab13]. Members [Rod15]. Membership [LG19]. Memory [HL15]. Meng [vC10]. Mentor [LVH^{+}17, War17]. Mentoring [ACHMW17, Gol17, HO17, Kwa17, Mor17, Sha17, VLZ17, VTK^{+}17, VKP^{+}17]. Mere [Cob15]. Mess [HR10]. Messy [MPP19]. Meta [BJ16]. Meta-Analyser [BJ16]. Method [CCB16, Hor13, Tuy19, Hoe12]. Methods [ALR17, Car19, IIDdL14, JM18, KCE13, LLS16, MGRL14, MCM^{+}11, May13, TCC^{+}15, VWB^{+}10, ZNY11]. Mexican [ER15]. Michael [FVMU10]. Michel [Fd12]. Might [Rob19]. Military [ER15]. Million [FKN^{+}17]. Millions [Ioa19]. Minimax [MA15]. Minimizing [Kou12]. Minimum [Lan17, Zha18]. Misleading [Gre19]. Misleads [Pog19]. Missing [CCB16, MPD19, YB10]. Missteps [Smi11]. Misunderstood [Arm16]. Misuse [Kme19]. Mixed [CGM19, WG11]. Mixture [CW11, DB18, FHMV19, RB18, ZBGZ14]. Model [And19, DAEP18, DKY17, FF18, Har14a, JS11, KD17, LRZG^{+}10, LL13, LDM10, Tuy17, Tuy19, WS15, ZPM^{+}11]. Model-Based [Har14a]. Model-Invariant [KD17]. Modeling [Bor14, CG19, Nar12, TO16, WBNTF19, WB18]. Models [ALL10, CGM19, CW11, FHMV19, GKP12, GGGV19, GD19, KOT12, LaM14, LH12, LLB16, LFS^{+}17, Mac17, PW13, RLS^{+}16, SDE19, TK18, WWGL19, WG11, Zha17, SMSZ17]. Modern [McN19]. Modes [DHW11]. Modified [FH19, Hor13]. Moment [Bry18, Fro12, GR13b, Inl10, O’N14, Por15]. Moment-Matching [Por15]. Moments [CJE19, Kon11, MT11, NL18]. Monotone [WS15]. Moore [Hor13]. Moore-Method [Hor13]. Morris [BXHH14]. Mosaic [Grö14]. Most [BCG18, Sam15]. Movement [WB18]. Moving [HS10, Mat19, WSL19]. MSLT [MT19]. MSLT-II [MT19]. Much [CJC19]. Multi [FKN^{+}17]. Multi-Million [FKN^{+}17]. Multimodal [DDF19]. Multimodel [BL13]. Multinomial [Chr20, Sza19, Tuy17, Tuy19]. Multiple [ALR17, Fie13b, HMR13, IIDdL14, KCE13, LPB13a, Smi11, WJ19, XY16, ZR10, vDvDG^{+}19]. Multiplicity [CLR13, LC09]. Multirater [QL16]. Multiresolution [LM14]. Multistage [IDE12, MS14]. Multivariate [Din16, SVG17, Wag16, ZPM^{+}11]. Musings [Goo19]. Must [BU15]. Must-Include [BU15]. My [KV18].

N [Fri15, Hol14]. Nano [BM10]. Nano-Project [BM10]. National [DWG15]. Nature [EA11, Sam13]. Navigating [SEC^{+}19]. Near [MT19, NR19]. Near-Balanced [NR19]. Near-Optimal [MT19]. Necessary [HMR13]. Need [Cob15, GA13, Har14a, TdCP17]. Needed [LNM16, YGP14]. Needs [Wag16]. Negative [Nie11]. Nested [MA19]. Net [Roc16]. Network [AD18]. News [PR11]. Next [HS10, HH15, KKM15, Lyn16, Rod15, Was15, vC10]. Neyman [Wri12]. NIH [ZC15]. Nissenbaum [Sta16]. No [ATG19, BH11]. Noiman [ANBB^{+}14]. Nominal [ZRK^{+}11]. Nominees [WASB10]. Non [BM17, Coh17]. Non-Asymptotic [Coh17]. Non-normal [BM17]. Nonbipartite [LGXB11]. Noncentral [Le 19]. Noncollapsibility [Gre10]. Nondenominational [Har14a]. Nonidentifiability [WIE13]. Nonignorable [MPD19]. Noninformative [SSS12]. Nonmonotonicity [FPW12]. Nonnormal [Dem16b]. Nonobvious [Tra19]. Nonparametric [FZ17, Hut18, Ric19]. Nonprofit [YGP14]. Nontransitivity [Fri15]. Normal [BM17, FR12, Gri13, HS12, Muk14b, Por19, SB15, Sti17, YB10, ZXC12, ZBGZ14]. Normal-Distribution-Based [YB10]. Normality [ANBB^{+}13, ANBB^{+}14, Muk14a, SEGMA18, WE13]. Note [ALL10, Arm16, Chr18a, Dem16b, GS14, HZ17, HV18, HHC17, LL13, LRR19, MA15, ME11, Sun18, Tuy17, UL15, Vel18a, ZR10]. Notes [Ost11d, Ost11a, Ost11b, Ost11c, Ost12a, Ost12b, Ost13a, Ost13b, Ost14]. Nothing [Eas15]. Notions [Muk10]. Novice [GS18]. Nuisance [Ang16]. Null [FH19, HO16]. Number [Coh19, LNM16]. Numbers [LMZ15, SR16]. NY [Sta16].

O [Fou20, TO16]. Objective [GJLW19]. Observation [And15, Por19]. Observations [SVR11]. Obtained [DF16]. Occurrences [Din14b]. OEIS [Dem18]. Offensive [WB18]. Office [Coc15]. Ogden [Din14a]. Okan [Sab19]. Old [HLU19]. Olympic [EA11]. Omissions [WK13]. One [AP10, And15, CLR13, LC09, LDM10, ML14]. One-Sided [CLR13, LC09]. One-Way [LDM10]. Online [KS15, YJCR15]. Only [GR13a]. Opportunities [Gol17, Hor15, KKM15]. Optimal [GdBPP19, KLS18, LGXB11, MT19, SD14, TO16, Zuo10]. Optimality [Por19]. Optimization [Dre19]. Optimum [Wri12]. Oracle [MCM^{+}11]. Order [CJE19, Grö14]. Ordered [LTN10, XY16]. Ordonnateur [Fd12]. Organization [BW18]. Organizations [YGP14]. Oscar [WASB10]. Other [Bac19, Lav19]. Our [Gib19, Hig19, LFH16, Rod15, Was15, Cob15]. Outcome [Chr20, Sza19]. Outcomes [CP15, KM19]. Overdispersed [GS14]. Overlap [NMR16]. Overlapping [WKW19]. Overview [WG11]. Owens [Fri15].

P [IDE15, Sta16]. Package [IDE15, KF14]. Packages [FGTM14, IIDdL14, KCE13, Nar12]. Packaging [MBM18]. Pairs [HRS10, JM18]. Papers [Ioa19, Sch10a]. Paradox [GZ17, SJD10, Wan18, Arm14b, Gre10, LM14, Pea14]. Paradoxical [Din14a, TOPC14]. Parallel [BR14]. Parameter [BG17, Chr17, HEJ11, HKM16, HMR13, MKR19, Zha18]. Parameters [Ang16]. Partial [PP10]. Partitional [LG19]. Passing [GR13b]. Passion [Eas10]. Passion-Driven [Eas10]. Patients [KRSR12]. Pattern [Pri11]. Paul [Zab13]. Peakedness [Wes14]. Pearl [Arm14b]. Pearson [Eas15]. Penalized [LP14, ML16, NSSLC12]. Penalized-Likelihood [ML16]. Penalties [WB17]. Penn [YJCR15]. Perceived [GR13a]. Percentage [YH15]. Percentage-Based [YH15]. Performance [MCM^{+}11]. Perils [MPP19]. Permutation [ZPM^{+}11]. Perspective [AD18, LVH^{+}17]. Perspectives [vDvDG^{+}19]. Petkova [Din14a]. Phi [Dem16b]. Philosophical [May13]. Phlegon [Mis19]. Pin [KO12]. Pioneers [Hal10, Sne19]. Pitches [SGBS17]. Pitfalls [WASB10]. Pivot [Ang16]. Pivotal [Tou17]. Place [KH19]. Placement [RST15]. Plan [SVR11]. Player [WB18]. Plays [TO16]. Plot [BR14]. Plots [Grö14, LFH16, Old16, Sti17]. Plus [GSK19, ML14]. POD [CCB16]. Point [HO16, Sam13, SEC^{+}19, SCLP17]. Poisson [And15, DF16, LP16, Sti13b]. Poker [BD11]. Police [BH12]. Policing [Men10]. Policy [RM19a]. Polyplot [SB11]. Pooling [XQL16]. Popes [SVCB18]. Population [FZ17, GA18, Hut18]. Populations [WKW19, ZXC12]. Portmanteau [BDP18]. Positive [Col19]. Positively [Fri15, LSO01]. Post [LR16, Mat19, Tra19]. Post-Hoc [LR16]. Poster [NR19]. Power [ANBB^{+}13, ANBB^{+}14, CGM19, FR19, FH19, Gri13, KOT12, LFH16, ME11, Muk14b, Was15]. pp [Sab19, Sta16]. Practical [DH11, IDE15, KF14, KS15, Lu19, Par19, Pog19]. Practice [GS18, Hub19, HG12, LAK16, LVH^{+}17, Pog19, Tra19]. Practices [HKPT18, PBD^{+}16]. Precision [And19, BS11]. Predicting [FF18]. Prediction [Har14a, PGCL14, RLS^{+}16, Wan10]. Predictive [Bil19, RM19b]. Prepared [BV11]. Preparing [HHH^{+}15, SB15]. Preprocessing [ZHM^{+}13]. Preserving [NMR16]. Presidents [Coc15]. Press [Cun12, Sab19, Sta16]. Previously [DHW11]. Primer [WKW19]. Primes [Coh16]. Principle [Pri11]. Principles [LAK16, MHWB19]. Prior [DKY17, JJ14, Lon10, VZH16]. Priors [DDV16, SSS12, TL10, Tuy17]. Privacy [Sta16]. Probabilistic [Pat18, Vel15]. Probabilities [BD11, Coh15, LTN10, PP10, Wan19]. Probability [AM14, DKY17, Gor11, Lo19, PL17, Por15, SD14, WSF10, ZLW13]. Probably [Coh16]. Problem [EIS17, HHC17, LWM15, MA15]. Problems [O’N14]. Procedure [DNBJC18, MKR19, SD14]. Procedures [GP13, MA19, WJ19, WG11]. Process [BM10, IGRP17, LP16, WL16b, ZSMS17]. Processes [Lit19, MG14]. Product [GS14]. Production [FF18]. Profession [HS10, Kme19, Was15]. Professional [VKP^{+}17, WB18]. Program [CP15, LRZG^{+}10, LHH15, MK15, VTK^{+}17, YJCR15]. Programs [HO17, LMZ15]. Progression [KRSR12]. Progression-Free [KRSR12]. Project [BM10, DDF19]. Project-Based [DDF19]. Projections [Grö14].

Projects [MTBG19, Smi11]. Prominent [Mic14]. Promise [PGCL14]. Promoter [Roc16]. Promotion [Wal18]. Proof [EIS17, Inl10, Pat18]. Proofs [Vel15]. Propensity [Zig16]. Properties [GD19, Rob19]. Property [Fri15, LSO01, Xu14]. Proportion [Fre10, JTL17]. Proportions [KMS10, Wri12]. Proposal [Col19]. Proposals [FGTM14]. Proposed [GSK19]. Protecting [VR15]. Providers [JS11]. Pseudo [YB10]. Psychological [MTBG19]. Psychology} [FBHW19]. Psychometrics [Sab19]. Public [All19, Sta16]. Publication [CG19]. Publishing [Loc19]. Pulling [BS10a]. Purpose [IGRP17, WL16b]. Putting [KH19].

Quadratic [Sam13]. Qualifying [BM10]. Qualitative [LHH15, PGA12]. Qualitative-Quantitative-Qualitative [LHH15]. Quality [HC19, SGBS17]. Quantile [Old16, Sti17]. Quantile-Quantile [Old16, Sti17]. Quantiles [FZ17, Hay14, Hut18]. Quantitative [LHH15, Car19]. Quantity [Tou17]. Quid [Fox10]. Quincunx [PR10]. Quintessential [PR10]. Quo [Fox10].

R [ANBB^{+}14, Chr17, Din14a, IDE15, Lu19, MS14, Muk14a, TO16, BCG18, EB18, GGGV19, KF14, MBM18, MH18, Ber19, Sab19, Dre19]. R-Squared [Chr17, GGGV19]. R.I.P. [Wes14]. Radical [Hig19]. Random [ALL10, Coh19, Fro12, GA18, GS14, IIDdL14, KCE13, LL13, LDM10, YB10]. Random-Effects [ALL10]. Randomistas [Hig19]. Randomization [KM19]. Randomized [FP12, GE11, KRSR12]. Range [Muk14a, NMR16, WE13]. Range-Preserving [NMR16]. Rank [MLL10, RB18, vDLMW18]. Rank-Based [MLL10]. Rao [GM16]. Rare [PW13]. Rates [BXHH14, FWW17, Lon10, VM13]. Ratio [BM17, Bed14, GP13, Jae16, VTH14, ZXC12]. Raton [Sab19]. Rcpp [EB18]. Re [VKP^{+}17]. Re-Defining [VKP^{+}17]. Readmission [ZRK^{+}11]. Real [BHLE15, JM18, KS15]. Real-life [JM18]. Real-World [KS15]. Really [Sam15]. Reanalysis [BH12]. Reasonably [Rob19]. Recognition [Pri11]. Recommendations [BB19, DCF^{+}17, Fou20, Kme19]. Reconsidering [MT19]. Reduced [LFS^{+}17]. Reduction [MPD19]. Reemerged [Fie13a]. Reference [Bor14]. Reformulations [Bed14]. Region [YGP14]. Regions [Jae16, LTN10]. Regression [DAEP18, Din14a, GGGV19, GD19, Gor16, HZ17, IIDdL14, JM18, KCE13, LFS^{+}17, MPD19, NSSLC12, OR14, RM19a, SDE19, Smi11, TOPC14, Xu14]. Regressions [KZ16]. Regressograms [GKP12]. Rejoinder [Arm14a, GR13b, Har14b, Kwa17, LPB13b, Men10]. Relabeling [CW11]. Relationship [Bro10, Dem16b]. Relative [RLS^{+}16]. Relevance [HC19]. Reliability [MGRL14, QL16]. Remark [Hon12, Hon15b]. Remedies [ZLW13]. REML [LDM10]. Renovation [Cob15]. Repeated [Hol14, ST13]. Replacement [IDE12, MS14]. Replication [ATG19, MTBG19]. Replications [Har19]. Reply [Din15, FWH^{+}14, Gri14, IDE14, KE14, Muk14a, SD15, ST14, SMSZ17, TOPC15, VM14, Vel18b, WE14, BXHH14, Hol14, MS14, Muk14b].

Reporting [HKPT18]. Representation [DB18]. Representatives [Wri12]. Reproducibility [Bil19, BS11, HC19]. Reproducibly [MBM18]. Reproducing [NSSLC12]. Required [VLZ17]. Requires [Bet19]. Requiring [CG19, SVR11]. Resampling [Hes15]. Research [CG19, HKPT18, HC19, Kme19, LRZG^{+}10, LPB13a, MTBG19, NL15, Pog19, Sam15, Car19]. Researchers [Hig19]. Resequencing [MGCR10]. Resolution [Gre19, LM14, Rob19]. Resolving [Bor14, FKN^{+}17]. Resource [DMM19]. Response [FP12, IDE15, IGRP17, KF14, Kit17, Sam13, WS15, vC10]. Responses [IDE12, MPD19, MS14]. Responsiveness [HC19]. Restrictions [Por15]. Result [Din14a, TOPC14]. Results [And19, Gui18, Har19, Loc19, MA19, O’N14]. Resurrecting [Arm14b]. Rethink [Cob15]. Rethinking [GZK^{+}11, LBR17]. Reversals [KD17]. Reversible [AKLS10, CGM19]. Reversion [Gor16]. Review [Ano15a, Ber19, Car19, Dre19, Hig19, Lit19, Lu19, Nar12, Ric19, Sab19, SF19, Sta16, Wan19]. Reviewers [Tra19]. Reviews [Ano10f, Ano10g, Ano10h, Ano11h, Ano11e, Ano11f, Ano11g, Ano12f, Ano12e, Ano12g, Ano12h, Ano13e, Ano13f, Ano13g, Ano13h, Ano14f, Ano14d, Ano14e, Ano15a, Ano15f, Ano15d, Ano15e, Ano16d, Ano16e, Ano16f, Ano16g, Ano17e, Ano17c, Ano17d, Ano18d, Ano18b, Ano18c, Ram17, SB14]. Revisited [LBR17]. Revisiting [Arm16, MA19, Wri13]. Rid [Goo19]. Right [Hol14, ST13]. Rights [Fie13b, LPB13a]. Rigorous [LH12]. Rillig [VR15]. Risk [Col19, GE11, RM19a, TG18]. RJMCMC [BL13]. Robust [QL16]. Robustness [HS12, MA19, Sti10]. Rodney [Sti13a]. Role [BLOP19, HHP19, War17, Zig16]. Ronald [Hal10]. Root [Fos11]. Rootograms [KZ16]. Rounding [XY16]. Rule [Fos11, KV18]. Run [FF18].

S [ANBB^{+}14, BXHH14, IIDdL14, TO16, Cun12]. Same [Sch10b]. Same-Day [Sch10b]. Sample [CGM19, FR12, GdBPP19, GJLW19, GA18, Hay11, HO16, KOT12, Kon11, LDM10, NL18, Sen12, SCLP17, WL16a, WSF10]. Sample-Size-Dependent [GdBPP19]. Sampling [BL13, GA18, O’N14, Wri12]. Samurai [XQL16]. SAS [Lu19, Wan19]. SAS^{T M} [WJ19]. Satisfy [Coh16, Coh17]. Scale [CMW15, DKY17, HS12, HKM16, MTBG19, TL18]. Scenarios [vDvDG^{+}19]. Schedule [FT10]. Schmidt [LaM14]. School [All19, FHP^{+}11, ZPL^{+}18]. Schools [ZC15]. Schwertman [Fri15]. Science [Bau15, BV11, CJC19, Goo19, HHH^{+}15, HI18, Kap18, Loc19, Ton19, Wal18, WBL18]. Scientific [Bil19, CG19, HHP19, HC19, Ioa19, O’H19]. Scientists [Bau18]. Scope [RST15]. Score [FPW12, Roc16, SLJ^{+}18]. Scores [Zig16]. Searches [BH12, WK13]. Secant [Din14b]. Second [BSC15, BGW^{+}19]. Second-Generation [BGW^{+}19]. Secret [FVMU10]. Section [Ost11d, Ost11a, Ost11b, Ost11c, Ost14, Ost12a, Ost12b, Ost13a, Ost13b]. See [ANBB^{+}13, ANBB^{+}14]. Segal [Hol14]. Segmentation [KKAB16]. Selection [DAEP18, GG15, MG14, WB17]. Self [DMM19, Old16].

Self-Calibrating [Old16]. Self-Learning [DMM19]. Semiparametric [FF18]. Sensitivity [SCLP17]. Sequences [Dem18]. Sequential [Fre10, IDE12, IDE15, KF14, MS14, ZNY11]. Serial [BDP18]. Series [Ber19, HL15, Kha15, RLS^{+}16]. Set [SR16]. Sets [Zha18]. Settings [Fie13b]. Several [Hay14, ZRK^{+}11]. Share [EL18]. Sharpening [LB19]. Sharper [Coh15]. Shaved [PP10, Sun18]. Short [HL15]. Should [BU15, Cun12, Gre19, Hes15, Man19, Muk14a, WE13]. Shoulders [Sne19]. Shrinkage [ZLW13]. Sided [CLR13, LC09]. sight [BH11]. Sign [OR14]. Signed [RB18]. Signed-Rank [RB18]. Significance [Fos11, GdBPP19, KS19, MGG^{+}19, ME11, Pog19]. Significant [HLU19, Joh19]. Similarity [GP13]. Simple [AM14, DKY17, DF16, GA18, LL13, LG19, Pat18, PS18, VTH14, WL16a, WIE13, vDvDG^{+}19, BH11]. Simpson [Arm14b, GZ17, Gre10, LM14, Pea14]. Simulation [GS14, HKPT18, IIDdL14, KCE13, KOT12, Man13, Sam13, TG18, TCC^{+}15]. Simulation-Based [KOT12, Man13, Sam13, TCC^{+}15]. Simulations [KS15]. Simultaneous [Hay14, HS12]. Single [IDE15, KF14]. Single-Arm [IDE15, KF14]. Size [Bro10, CGM19, GdBPP19, GSK19, GA18, HO16, MT19, Pog19, WSF10]. Sizes [FR12]. Skating [EA11]. Skewed [GA18, VZH16]. Skewness [JRP11, SB11]. Skewness-Invariant [JRP11]. Sleuthing [EA11]. Sliders [JJ14]. Slope [Nie11]. Small [FR12, KOT12, Lon10, LDM10]. Smaller [KV18]. Smoothing [WS15]. Social [FBHW19]. Society [Kot10]. Software [FGTM14, HEJ11, KP11, LDM10, Nar12, WG11, ZNY11]. Sojourning [LP16]. Solution [MA15]. Some [DHW11, Gre19, Gui18, O’N14, TdCP17, Hub19]. Sometimes [Hal10]. Space [Ber19, XQL16]. Space-Filling [XQL16]. Space-Time [Ber19]. Spaces [NSSLC12]. Sparse [ZRK^{+}11]. Spatial [Ber19, TK18, YGP14]. Spatially [HR10]. Spatially-Correlated [HR10]. Special [HH15, WBL18]. Specification [LH12]. Specificity [SCLP17]. Specified [GS14]. Specifying [SSS12]. Spline [WS15]. Spread [SB11]. Spreadsheet [JJ14, KP11]. Spreadsheets [BW18]. SPSS^{T M} [WJ19]. Square [Che11, FKN^{+}17, Fos11, GD19, KD17]. Square-Root [Fos11]. Squared [BG17, Chr17, Kou12, GGGV19, RM19a]. Stabilizing [KMS10]. Stage [MA15, MG14, SFSM17]. Stages [ACHMW17]. Stand [Sne19]. Standard [Muk14a, SR16, WE13]. Starting [SEC^{+}19]. State [BH12, YJCR15]. Statement [IGRP17, WL16b]. Statewide [ZPL^{+}18]. Stationary [Sam13]. Statistical [ACHMW17, And19, BGMA13, Bor14, BHLE15, Bro11, BLOP19, CG19, ÇRR18, DDF19, DHW11, EA11, FGTM14, Fra19, FBHW19, FVMU10, Fd12, GS18, Hal10, HEJ11, HKPT18, HS17, Hor15, HHP19, Hub19, HG12, IIDdL14, KP11, KO12, Kha15, KCE13, KS15, LBR17, Loc19, Lom13, LGXB11, Man19, Mar17, MM15, McN19, MGG^{+}19, MG14, Por16, Pri11, Rob19, Sch10a, SB15, SJD10, SLG19, TO16, TCC^{+}15, Ton19, Van15, VWB^{+}10, WW11, vDvDG^{+}19, SF19]. Statistically [HLU19]. Statistician [All19, BXHH14, Fri15, Hol14, IIDdL14, Mic14, MS14, ANBB^{+}14, Din14a, Fou20, IDE15, Muk14a, Muk14b]. Statisticians [Lyn16, VKP^{+}17]. Statistics [ATG19, BU15, BSC15, CJC19, CMW15, CFP18, Coh16, Dav11, DAEP18, DCF^{+}17, DMM19, Eas10, ESC18, FWH^{+}13, FPW12, FHP^{+}11, GZK^{+}11, Gib19, Gol17, Goo19, GB15, Gri13, Gri15, HHH^{+}15, Hes15, HS10, Hor13, Hor15, HH15, Joh13, Joh19, KKM15, Kot10, LRF11, LMZ15, LVH^{+}17, MGCR10, MHWB19, MK15, ME11, ML14, Muk14b, NL10, NL15, PR11, dRFF17, Ric19, RST15, Sch12, Sne19, Sol10, Sti13a, Utt15, VLZ17, Wal18, Wan18, YJCR15, ZHM^{+}13]. Statistique [Fd12]. Stats [GA13, Kap18]. Steampunk [BJ16]. Stefan [Sta16]. Stem [Mis19]. Stem-and-Leaf [Mis19]. Step [VTK^{+}17]. Stepwise [RM19a]. Stochastic [ZSMS17, Lit19]. Stodden [Sta16]. Stop [Kme19]. Strategies [BS10a, EIS17, Par19, SFSM17]. Strategy [TO16]. Stratified [MLL10]. Strength [FT10, RM19b]. Strengthen [KS15]. strikes [BH11]. Stringency [CG19]. Strong [FH19]. Structural [Nar12, TK18]. Structure [WB17]. Structured [SVR11]. Student [GZK^{+}11, KS15, Smi11, Wag16, WRH19, Le 19]. Students [All19, BV11, BM10, GA13, Gri13, GRTM16, HHH^{+}15, HH15, Muk14b, PS17]. Studies [BHLE15, HKPT18, IDE15, Kha15, KF14, ZHM^{+}13]. Study [EA11, HEJ11, IIDdL14, JM18, KCE13, LR16, RLS^{+}16, Wag16, WASB10, YH15, ZRK^{+}11]. Style [BJ16]. Subgroup [LR16]. Subjective [GJLW19, O’H19]. Submissions [Tra19]. Succeed [BV11]. Success [SFSM17]. Successful [Hub19]. Sudoku [XQL16]. Sudoku-Based [XQL16]. Sufficiency [LBR17, Rou19]. Sufficient [Wan18]. Suicides [CCFJ18]. Sum [Coh19]. Summands [Coh19]. Supplant [Man19]. Supporting [GSK19]. Surface [Sam13]. Surgery [ZRK^{+}11]. Surveillance [MT19]. Survey [GR19, HKPT18, Mic14]. Survival [CJE19, KRSR12, SVCB18, Wan19, WASB10, Lu19]. Sustaining [VTK^{+}17]. Switching [Pog19]. Swivelling [BD11]. Symbolic [SLJ^{+}18]. Symmetrizing [DF16]. Synchronous [WYG17]. System [MGRL14, ZPL^{+}18]. Systems [Fie13b, KO12, LPB13a]. Szabo [Chr20].

T [Din14a, Din16]. Table [KV18]. Tables [Che11, Gre10, SF19]. Tail [Coh15, Lo19]. Takes [Sch12]. Tale [FHB^{+}13, KV18]. Talk [Bry18]. Tangent [BC12]. Target [BG17, Chr17]. Tarpey [Din14a]. Taught [ZHM^{+}13]. Taylor [Coh16, Coh17, Dem18, FWH^{+}13]. Teachers [Hes15]. Teaching [Ano10f, Ano10g, Ano10h, Ano11h, Ano11e, Ano11f, Ano11g, Ano12f, Ano12e, Ano12g, Ano12h, Ano13e, Ano13f, Ano13g, Ano13h, Ano14f, Ano14d, Ano14e, Ano15a, Ano15f, Ano15d, Ano15e, Ano16d, Ano16e, Ano16f, Ano16g, Ano17e, Ano17c, Ano17d, Ano18d, Ano18b, Ano18c, BBM13, ÇRR18, CMW15, DDF19, ESC18, EIS17, GS18, HI18, HH15, Kap18, NP16, Par19, PS17, dRFF17, Ram17, RM19b, SB14, Sol10, Wag16]. Teams [FT10]. Technique [BM17]. Templates [dRFF17]. Ten [KO12]. Ten-Pin [KO12]. Tent [VLZ17]. Terms [Coc15]. Test [ANBB^{+}13, ANBB^{+}14, AP10, Bro10, Che11, Chr20, DNBJC18, FR19, FH19, HRS10, HL15, MLL10, ME11, ML14, RB18, SCLP17, Sza19, VTH14, WL16a, WGW12, ZXC12]. Tested [All19].

Testing [BDP18, BHLE15, CLR13, FIME17, GP13, KS19, LC09, MA15, MA19, Man19, MKR19, MPP19, Par19, Ric10, SVR11, WL16a, YZS^{+}14, YH15]. Tests [CLR13, FWW17, LC09, MG14, SEGMA18, ZPM^{+}11]. Tetrachoric [Dem16b]. Textbooks [DCF^{+}17, Ric19]. Their [FBHW19, Gre19, Gri13, Muk14b, Van15, Coc15]. Them [Sch12]. Themes [Utt15]. Theorem [GJLW19, Pat19, Zig16, Arm16, FWH^{+}13, Inl10, dSPCdS17, RM19b]. Theorems [BBM13]. Theoretic [Ric10]. Theory [EIS17, Gri13, Man19, Muk14b, Lit19]. There [ATG19, Eas15]. They’ll [DMM19]. Think [HHH^{+}15, HH15]. Thinkers [CFP18]. Thinking [Bau15, Men10, SLG19, TCC^{+}15, Ton19]. Third [Arm14b]. Thoughts [Kot10]. Three [BB19, Din14b, Fou20, HEJ11, Jae16, LCD10, ML14]. Three-Dimensional [Jae16]. Three-Parameter [HEJ11]. Threshold [Bet19]. Throughout [ÇRR18, TCC^{+}15]. Time [Ber19, HS17, HL15, Kha15, RLS^{+}16, WASB10]. Time-Dependent [WASB10]. Tolerance [GA13]. Too [Cob15]. Tool [McN19, dRFF17, SDE19]. Tools [ÇRR18, GdBPP19]. Topics [BU15, MGCR10]. Torres [Hol14]. Tough [HLU19]. Tournament [ACF16]. Traditional [Sch12, WYG17]. Training [ACHMW17, LMZ15, Lyn16, Men10]. Transformations [Gri13, Muk14b]. Transitive [Fri15, LSO01]. Treated [KRSR12]. Treatment [Man19, MT19, PGA12]. Trend [Chr20, Sza19]. Trial [Dre19, KRSR12, Man19, MT19]. Trick [FR19]. Trinomial [SDE19]. Trivia [WRH19]. Trivia-Based [WRH19]. Try [Zil19]. Tutorial [NSSLC12]. Twenty [BGMA13, Joh13]. Twenty-Five [BGMA13]. Twin [Coh16]. Two [Bed14, CLR13, FHMV19, FWW17, FHB^{+}13, GJLW19, Hal10, Hon14b, Jae16, KMS10, LC09, LC15, LL13, MA15, MG14, RS13, WL16a, WASB10, ZXC12, Zha18, vDvDG^{+}19]. Two- [Jae16]. Two-Parameter [Zha18]. Two-Sample [GJLW19, WL16a]. Two-Stage [MA15, MG14]. Type [LDM10, SEGMA18].

U.S. [LMZ15, Wri12, ZC15]. Ultrahigh [GD19]. Ultrahigh-Dimensional [GD19]. Unattributed [Arm16]. Unbiased [GM16, LL13]. Unbounded [LWM15]. Uncertain [CJC19]. Uncertainty [And19, Hod17, KKAB16, Kit17, Lon10]. Undergraduate [BU15, Cob15, Hes15, HH15, Kot10, LRF11, LRZG^{+}10, LHH15, MK15, TCC^{+}15, Wag16, War17]. Undergraduates [Bau15, NL15, SB15, Wit17]. Underlying [Dem16b]. Understand [Hor13, LAK16]. Understanding [GB15, KSW16, KS15, LWM15, Pea14, WRH19]. Unidimensional [MKR19]. Unified [ML14, PFG16]. Uniform [ZBGZ14]. Unique [KD17, Mic14]. Univariate [WDCH18]. Universal [Ang16]. Universe [Tra19]. University [Sta16, YJCR15]. Unknown [Hay14]. Unrecognized [DHW11]. Unusual [JS11]. Upper [DH11]. Use [BB19, DMM19, EIRS17, Fou20, GSK19, ZPL^{+}18]. Used [FBHW19, Muk14a, WE13]. Useful [BBM13, Bed14, Dav11, Grö14, O’N14, Tou17, TdCP17]. Using [And15, CLR13, CJE19, Dre19, EIS17, FF18, GE11, HRS10, IDE15, JM18, KS19, KF14, KZ16, KS15, LC09, LCG16, MBM18, NMR16, RLS^{+}16, RS13, Sab19, Sch10a, SLJ^{+}18, TCC^{+}15, WW11, Wan18, Wan19].

vadis [Fox10]. Valid [Gre19]. Validation [MCM^{+}11]. Validity [LR16, YZS^{+}14]. Value [Bet19, BS11, Bro10, Dem16a, FWH^{+}13, FBHW19, Gib19, GSK19, KH19, MKR19, Pat19, Fra19, MHWB19]. Values [BB19, BGW^{+}19, Col19, Fou20, Goo19, Gre19, HO16, Ioa19, IGRP17, KS19, Kme19, KV18, KW19, Mar17, RM19a, Rou19, WL16b, Zil19]. VanDerwerken [Fre19]. Vardeman [BXHH14]. Variability [GRTM16]. Variable [Arm14b, GG15, YZS^{+}14]. Variables [Coh19, Fro12, GS14, TG18, XY16, ZZT^{+}19, ZRK^{+}11]. Variance [Coh17, Kou12, KMS10, LP14, LL13, LDM10, ML16, Sen12, SVG17]. Variances [FR12, SCLP17]. Variational [OW10]. Variations [LFH16]. Velilla [Chr18a]. Veresoglou [VR15]. Verification [PFG16]. Version [Bry18]. Versus [Sch12, Sch10b, WYG17]. Via [IDE15, BJ16, KF14, LP14, LM14, Mat19]. Victoria [Sta16]. Violence [ER15]. Visibility [Gib19]. Visualization [NP16]. Visualizations [WKW19]. Visualizing [Gön13, Grö14, KZ16, SEGMA18, SR16, SB11]. Volume [Ano10c, Zha18]. Voting [BXHH14, VM13]. vs [JM18].

W [ANBB^{+}14]. Wald [LDM10]. Wallet [PL17]. Wants [All19]. War [ER15]. Warr [Muk14a]. Water [Sti17]. Water-Filling [Sti17]. Way [DH11, LDM10]. Weibull [HEJ11]. Weighing [BJ16]. Where [YGP14, VKP^{+}17]. White [Bau18]. Whitening [KLS18]. Whitney [DNBJC18, FH19]. Who [All19, Hoe12, PS17, Rod15, Sch12, Sol10, VKP^{+}17]. Whose [HS17]. Width [Fre10]. Wikipedia [DMM19]. Wilcoxon [DNBJC18, FH19, RB18]. Will [BV11, Hub19, Rod15]. Winning [BD11]. Within [Gri15, HV18, PB16, Sch10b]. Within-the-Bar [PB16]. Without [IDE12, MS14]. Women [Gol17]. Work [MBM18, Mic14, SB15]. World [CG19, KS15, SEC^{+}19, VR15, WSL19, Hig19]. Wrangling [MH18]. Wrong [Eas15, TdCP17].

Xiao [vC10]. Xiao-Li [vC10]. xix [Sta16]. xxiii [Sab19].

Y.-K [IIDdL14]. Years [HG12, KKM15, Utt15, vC10]. Yes [May13]. York [Sta16].

Zero [KM19, Tuy19]. Zhang [Hut18].

References

Abdul-Chani:2016:IBE

[ACF16] Christopher Abdul-Chani and Jesse Frey. Improving the Big East Conference Basketball Tournament. The American Statistician, 70(4):342–349, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anderson-Cook:2017:SME

[ACHMW17] Christine M. Anderson-Cook, Michael S. Hamada, Leslie M. Moore, and Joanne R. Wendelberger. Statistical mentoring at early training and career stages. The American Statistician, 71 (1):6–14, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

An:2018:LCI

[AD18] Weihua An and Ying Ding. The landscape of causal inference: Perspective from citation network analysis. The American Statistician, 72(3):265–277, 2018. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://amstat.

tandfonline.com/doi/full/10.1080/00031305.2017.1360794.

Annis:2010:ERM

[AKLS10] David H. Annis, Peter C. Kiessler, Robert Lund, and Tara L. Steuber. Estimation in reversible Markov chains. The American Statistician, 64(2):116–120, May 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Alonso:2010:NIR

[ALL10] Ariel Alonso, Saskia Litière, and Annouschka Laenen. A note on the indeterminacy of the random-effects distribution in hierarchical models. The American Statistician, 64(4):318–324, November 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Allen:2019:WWS

[All19] Jeff Allen. Who wants to be a statistician? An analysis of ACTtested public school students. The American Statistician, 73(3): 253–263, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://amstat.tandfonline.com/doi/
full/10.1080/00031305.2017.1419143.

Akande:2017:ECM

[ALR17] Olanrewaju Akande, Fan Li, and Jerome Reiter. An empirical comparison of multiple imputation methods for categorical data. The American Statistician, 71(2):162–170, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2016.1277158.

Aoudia:2014:SCB

[AM14] Djilali Ait Aoudia and Éric Marchand. On a simple construction of a bivariate probability function with a common marginal. The American Statistician, 68(3):170–173, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Aldor-Noiman:2013:PSN

[ANBB^{+}13] Sivan Aldor-Noiman, Lawrence D. Brown, Andreas Buja, Wolfgang Rolke, and Robert A. Stine. The power to see: A new graphical test of normality. The American Statistician, 67(4): 249–260, 2013. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See comment [ANBB^{+}14].

Aldor-Noiman:2014:ANB

[ANBB^{+}14] Sivan Aldor-Noiman, Lawrence D. Brown, Andreas Buja, Wolfgang Rolke, and Robert A. Stine. Comment: Aldor-Noiman, S., Brown, L. D., Buja, A., Rolke, W., and Stine, R. A. (2013), “The Power to See: A New Graphical Test of Normality,” The American Statistician, 67, 249–260. The American Statistician, 68(4): 318, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See [ANBB^{+}13].

Andersson:2015:CAC

[And15] Per Gösta Andersson. A classroom approach to the construction of an approximate confidence interval of a Poisson mean using one observation. The American Statistician, 69(3):160–164, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00031305.2015.1056830.

Anderson:2019:ASR

[And19] Andrew A. Anderson. Assessing statistical results: Magnitude, precision, and model uncertainty. The American Statistician, 73 (S1):118–121, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/00031305.2018.1537889.

Angus:2016:BUP

[Ang16] John E. Angus. Bootstrapping a universal pivot when nuisance parameters are estimated. The American Statistician, 70(1):100– 107, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2010:EC

[Ano10a] Anonymous. 2010 editorial collaborators. The American Statistician, 64(4):364–365, November 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2010:C

[Ano10b] Anonymous. Correction. The American Statistician, 64(2):195, May 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2010:IV

[Ano10c] Anonymous. Index to volume 64. The American Statistician, 64 (4):366–368, November 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2010:LEa

[Ano10d] Anonymous. Letters to the Editor. The American Statistician, 64 (3):276, August 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2010:LEb

[Ano10e] Anonymous. Letters to the Editor. The American Statistician, 64 (4):362–363, November 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2010:RBTa

[Ano10f] Anonymous. Reviews of books and teaching materials. The American Statistician, 64(1):88–96, February 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2010:RBTb

[Ano10g] Anonymous. Reviews of books and teaching materials. The American Statistician, 64(3):268–275, August 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2010:RBTc

[Ano10h] Anonymous. Reviews of books and teaching materials. The American Statistician, 64(4):357–361, November 2010. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2011:EC

[Ano11a] Anonymous. 2011 editorial collaborators. The American Statistician, 65(4):298–299, November 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2011:LEa

[Ano11b] Anonymous. Letters to the Editors. The American Statistician, 65(1):69, February 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2011:LEb

[Ano11c] Anonymous. Letters to the Editors. The American Statistician, 65(2):140, May 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2011:LEd

[Ano11d] Anonymous. Letters to the Editors. The American Statistician, 65(3):209–211, August 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2011:RBTa

[Ano11e] Anonymous. Reviews of books and teaching materials. The American Statistician, 65(1):61–68, February 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2011:RBTb

[Ano11f] Anonymous. Reviews of books and teaching materials. The American Statistician, 65(2):136–139, May 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2011:RBTc

[Ano11g] Anonymous. Reviews of books and teaching materials. The American Statistician, 65(3):201–208, August 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2011:RBT

[Ano11h] Anonymous. Reviews of books and teaching materials. The American Statistician, 65(4):283–297, November 2011. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2012:EBE

[Ano12a] Anonymous. Editorial board EOV. The American Statistician, 66 (4):ebi, 2012. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2012:EC

[Ano12b] Anonymous. Editorial collaborators. The American Statistician, 66(4):246–247, 2012. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2012:LEa

[Ano12c] Anonymous. Letter to the Editor. The American Statistician, 66 (2):152–153, 2012. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/00031305.2012.697678.

Anonymous:2012:LEb

[Ano12d] Anonymous. Letter to the Editor. The American Statistician, 66 (4):243–245, 2012. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2012:RBTa

[Ano12e] Anonymous. Reviews of books and teaching materials. The American Statistician, 66(1):67–74, 2012. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2012:RBT

[Ano12f] Anonymous. Reviews of books and teaching materials. The American Statistician, 66(2):139–151, 2012. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://amstat.

tandfonline.com/doi/full/10.1080/00031305.2012.695972.

Anonymous:2012:RBTb

[Ano12g] Anonymous. Reviews of books and teaching materials. The American Statistician, 66(3):201–206, 2012. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2012:RBTc

[Ano12h] Anonymous. Reviews of books and teaching materials. The American Statistician, 66(4):239–242, 2012. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2013:EC

[Ano13a] Anonymous. Editorial collaborators. The American Statistician, 67(4):267–268, 2013. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2013:LEb

[Ano13b] Anonymous. Letter to the Editor. The American Statistician, 67 (2):114, 2013. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2013:LEc

[Ano13c] Anonymous. Letter to the Editor. The American Statistician, 67 (3):190, 2013. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2013:LEa

[Ano13d] Anonymous. Letters to the Editor. The American Statistician, 67(1):64–66, 2013. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2013:RBTa

[Ano13e] Anonymous. Reviews of books and teaching materials. The American Statistician, 67(1):60–63, 2013. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2013:RBTb

[Ano13f] Anonymous. Reviews of books and teaching Materials. The American Statistician, 67(2):109–113, 2013. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2013:RBTc

[Ano13g] Anonymous. Reviews of books and teaching materials. The American Statistician, 67(3):183–189, 2013. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2013:RBTd

[Ano13h] Anonymous. Reviews of books and teaching materials. The American Statistician, 67(4):261–265, 2013. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2014:C

[Ano14a] Anonymous. Correction. The American Statistician, 68(2):132, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See [Arm14b].

Anonymous:2014:EBE

[Ano14b] Anonymous. Editorial board EOV. The American Statistician, 68 (4):ebi, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2014:EC

[Ano14c] Anonymous. Editorial collaborators. The American Statistician, 68(4):319–320, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2014:RBTa

[Ano14d] Anonymous. Reviews of books and teaching materials. The American Statistician, 68(2):117–124, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2014:RBTb

[Ano14e] Anonymous. Reviews of books and teaching materials. The American Statistician, 68(3):212–219, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2014:RBT

[Ano14f] Anonymous. Reviews of books and teaching materials. The American Statistician, 68(4):307–315, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2015:BRR

[Ano15a] Anonymous. Book review: Reviews of books and teaching materials. The American Statistician, 69(1):53–61, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2015:EBE

[Ano15b] Anonymous. Editorial board EOV. The American Statistician, 69 (4):ebi, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2015:EC

[Ano15c] Anonymous. Editorial collaborators. The American Statistician, 69(4):435–438, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2015:RBTa

[Ano15d] Anonymous. Reviews of books and teaching materials. The American Statistician, 69(2):149–153, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2015:RBTb

[Ano15e] Anonymous. Reviews of books and teaching materials. The American Statistician, 69(3):244–252, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/00031305.2015.1068616.

Anonymous:2015:RBT

[Ano15f] Anonymous. Reviews of books and teaching materials. The American Statistician, 69(4):425–434, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2016:C

[Ano16a] Anonymous. Correction. The American Statistician, 70(2):225, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2016:EBE

[Ano16b] Anonymous. Editorial Board EOV. The American Statistician, 70(4):ebi, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2016:EC

[Ano16c] Anonymous. Editorial collaborators. The American Statistician, 70(4):435–437, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2016:RBTa

[Ano16d] Anonymous. Reviews of books and teaching materials. The American Statistician, 70(1):120–126, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2016:RBTb

[Ano16e] Anonymous. Reviews of books and teaching materials. The American Statistician, 70(2):221–223, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2016:RBTc

[Ano16f] Anonymous. Reviews of books and teaching materials. The American Statistician, 70(3):313–??, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2016:RBTd

[Ano16g] Anonymous. Reviews of books and teaching materials. The American Statistician, 70(4):424–433, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2017:EBE

[Ano17a] Anonymous. Editorial board EOV. The American Statistician, 71(4):??, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://amstat.tandfonline.com/doi/
full/10.1080/00031305.2017.1400355.

Anonymous:2017:EC

[Ano17b] Anonymous. Editorial collaborators. The American Statistician, 71(4):376–377, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/00031305.2017.1395629.

Anonymous:2017:RBTa

[Ano17c]  Anonymous.  Reviews of books and teaching materials.

The American Statistician, 71(2):182–186, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://amstat.tandfonline.com/doi/full/10.1080/
00031305.2017.1325631.

Anonymous:2017:RBTb

[Ano17d]  Anonymous.  Reviews of books and teaching materials.

The American Statistician, 71(3):282–289, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://amstat.tandfonline.com/doi/full/10.1080/
00031305.2017.1367180.

Anonymous:2017:RBT

[Ano17e]  Anonymous.  Reviews of books and teaching materials.

The American Statistician, 71(4):369–372, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://amstat.tandfonline.com/doi/full/10.1080/
00031305.2017.1395630.

Anonymous:2018:C

[Ano18a] Anonymous. Corrigenda. The American Statistician, 72(4):394, 2018. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2018.1523641.

Anonymous:2018:RBTa

[Ano18b]  Anonymous.  Reviews of books and teaching materials.

The American Statistician, 72(1):105–113, 2018. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://amstat.tandfonline.com/doi/full/10.1080/
00031305.2018.1444855.

Anonymous:2018:RBTb

[Ano18c]  Anonymous.  Reviews of books and teaching materials.

The American Statistician, 72(2):206–212, 2018. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://amstat.tandfonline.com/doi/full/10.1080/
00031305.2018.1469927.

Anonymous:2018:RBT

[Ano18d]  Anonymous.  Reviews of books and teaching materials.

The American Statistician, 72(3):295–299, 2018. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://amstat.tandfonline.com/doi/full/10.1080/
00031305.2018.1496649.

Anonymous:2019:C

[Ano19a] Anonymous. Correction. The American Statistician, 73(4):420, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2019.1660112.

Anonymous:2019:ECa

[Ano19b] Anonymous. Editorial collaborators. The American Statistician, 73(1):106–108, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/00031305.2018.1538832.

Anonymous:2019:ECb

[Ano19c] Anonymous. Editorial collaborators. The American Statistician, 73(4):420–421, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2019.1680048.

Anonymous:2019:I

[Ano19d] Anonymous. Index. The American Statistician, 73(1):94–104, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2018.1538846.

Aldous:2010:WCO

[AP10] David Aldous and Tung Phan. When can one test an explanation? Compare and contrast Benford’s Law and the fuzzy CLT. The American Statistician, 64(3):221–227, August 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Armistead:2014:R

[Arm14a] Timothy Armistead. Rejoinder. The American Statistician, 68(1): 30–31, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Armistead:2014:RTV

[Arm14b] Timothy W. Armistead. Resurrecting the third variable: A critique of Pearl’s causal analysis of Simpson’s Paradox. The American Statistician, 68(1):1–7, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See comment [Pea14, LM14], reply [Arm14a], and correction [Ano14a].

Armistead:2016:MUR

[Arm16] Timothy W. Armistead. Misunderstood and unattributed: Revisiting M. H. Doolittle’s measures of association, with a note on Bayes’ Theorem. The American Statistician, 70(1):63–73, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Amrhein:2019:ISD

[ATG19] Valentin Amrhein, David Trafimow, and Sander Greenland. Inferential statistics as descriptive statistics: There is no replication crisis if we don’t expect replication. The American Statistician, 73 (S1):262–270, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/00031305.2018.1543137.

Bacchetti:2019:OAC

[Bac19] Peter Bacchetti. The other arbitrary cutoff. The American Statistician, 73(4):413–414, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2019.1654920.

Baumer:2015:DSC

[Bau15] Ben Baumer. A data science course for undergraduates: Thinking with data. The American Statistician, 69(4):334–342, 2015. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Baumer:2018:LBW

[Bau18] Benjamin S. Baumer. Lessons from between the white lines for isolated data scientists. The American Statistician, 72(1):66–71, 2018. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2017.1375985.

Berry:2014:C

[BB14] Donald A. Berry and Scott M. Berry. Comment. The American Statistician, 68(2):88–89, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See [Har14a].

Benjamin:2019:TRI

[BB19] Daniel J. Benjamin and James O. Berger. Three recommendations for improving the use of p-values. The American Statistician, 73 (S1):186–191, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/00031305.2018.1543135. See [Fou20].

Bagui:2013:FCE

[BBM13] Subhash C. Bagui, Dulal K. Bhaumik, and K. L. Mehra. A few counter examples useful in teaching Central Limit Theorems. The American Statistician, 67(1):49–56, 2013. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Berrendero:2012:TC

[BC12] José R. Berrendero and Javier Cárcamo. The tangent classifier. The American Statistician, 66(3):185–194, 2012. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bion:2018:HRH

[BCG18] Ricardo Bion, Robert Chang, and Jason Goodman. How R helps Airbnb make the most of its data. The American Statistician, 72(1):46–52, 2018. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/00031305.2017.1392362.

Barr:2011:SPW

[BD11] G. D. I. Barr and K. Dowie. Swivelling probabilities — the winning hand at poker. The American Statistician, 65(3):170–176, August 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bagnato:2018:TSI

[BDP18] Luca Bagnato, Lucio De Capitani, and Antonio Punzo. Testing for serial independence: Beyond the portmanteau approach. The American Statistician, 72(3):219–238, 2018. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2016.1264314.

Bedrick:2014:TUR

[Bed14] Edward J. Bedrick. Two useful reformulations of the hazard ratio. The American Statistician, 68(1):36–41, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bergen:2019:BRD

[Ber19] Silas Bergen. Book review: Displaying Time Series, Spatial, and Space-Time Data with R, 2nd ed. The American Statistician, 73 (3):310–311, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/00031305.2019.1641357.

Betensky:2019:VRC

[Bet19] Rebecca A. Betensky. The p-value requires context, not a threshold. The American Statistician, 73(S1):115–117, 2019.

CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2018.1529624.

Bar-Gera:2017:TPA

[BG17] Hillel Bar-Gera. The target parameter of adjusted R-squared in fixed-design experiments. The American Statistician, 71(2):112– 119, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://amstat.tandfonline.com/doi/full/
10.1080/00031305.2016.1200489. See comment [Chr17].

Behar:2013:TFA

[BGMA13] Roberto Behar, Pere Grima, and Lluís Marco-Almagro. Twentyfive analogies for explaining statistical concepts. The American Statistician, 67(1):44–48, 2013. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Blume:2019:ISG

[BGW^{+}19] Jeffrey D. Blume, Robert A. Greevy, Valerie F. Welty, Jeffrey R. Smith, and William D. Dupont. An introduction to secondgeneration p-values. The American Statistician, 73(S1):157–167, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2018.1537893.

Berger:2011:BLS

[BH11] Arno Berger and Theodore P. Hill. Benford’s law strikes back: no simple explanation in sight for mathematical gem. The Mathematical Intelligencer, 33(1):85–91, March 2011. CODEN MAINDC. ISSN 0343-6993 (print), 1866-7414 (electronic). See letter [Hil11].

Barbe:2012:CRM

[BH12] Philippe Barbe and William C. Horrace. A critical reanalysis of Maryland State Police searches. The American Statistician, 66 (1):1–7, 2012. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Boulesteix:2015:SFH

[BHLE15] Anne-Laure Boulesteix, Robert Hable, Sabine Lauer, and Manuel J. A. Eugster. A statistical framework for hypothesis testing in real data comparison studies. The American Statistician, 69 (3):201–212, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00031305.2015.1005128.

Billheimer:2019:PIS

[Bil19] Dean Billheimer. Predictive inference and scientific reproducibility. The American Statistician, 73(S1):291–295, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2018.1518270.

Bowden:2016:WES

[BJ16] Jack Bowden and Chris Jackson. Weighing evidence “Steampunk” style via the meta-analyser. The American Statistician, 70(4):385–394, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Barker:2013:BMI

[BL13] Richard J. Barker and William A. Link. Bayesian multimodel inference by RJMCMC: A Gibbs sampling approach. The American Statistician, 67(3):150–156, 2013. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Brownstein:2019:REJ

[BLOP19] Naomi C. Brownstein, Thomas A. Louis, Anthony O’Hagan, and Jane Pendergast. The role of expert judgment in statistical inference and evidence-based decision-making. The American Statistician, 73(S1):56–68, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://amstat.

tandfonline.com/doi/full/10.1080/00031305.2018.1529623.

Blitzstein:2010:NPQ

[BM10] Joseph Blitzstein and Xiao-Li Meng. Nano-project qualifying exam process: An intensified dialogue between students and faculty. The American Statistician, 64(4):282–290, November 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bagui:2017:CKD

[BM17] Subhash Bagui and K. L. Mehra. Convergence of known distributions to limiting normal or non-normal distributions: An elementary ratio technique. The American Statistician, 71(3):265–271, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2017.1322001.

Bordley:2014:RCF

[Bor14] Robert F. Bordley. Reference class forecasting: Resolving its challenge to statistical modeling. The American Statistician, 68(4): 221–229, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Burgin:2014:DPC

[BR14] Reto Bürgin and Gilbert Ritschard. A decorated parallel coordinate plot for categorical longitudinal data. The American Statistician, 68(2):98–103, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Browne:2010:TVR

[Bro10] Richard H. Browne. The t-test p value and its relationship to the effect size and P (X > Y). The American Statistician, 64(1):30– 33, February 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Broemeling:2011:AES

[Bro11] Lyle D. Broemeling. An account of early statistical inference in Arab cryptology. The American Statistician, 65(4):255–257, November 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bryan:2018:EMD

[Bry18] Jennifer Bryan. Excuse me, do you have a moment to talk about version control? The American Statistician, 72(1):20–27, 2018. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2017.1399928.

Beaudoin:2010:SPG

[BS10a] David Beaudoin and Tim B. Swartz. Strategies for pulling the goalie in hockey. The American Statistician, 64(3):197–204, August 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Block:2010:GEB

[BS10b] Henry W. Block and Thomas H. Savits. A general example for Benford data. The American Statistician, 64(4):335–339, November 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Boos:2011:VPR

[BS11] Dennis D. Boos and Leonard A. Stefanski. P -value precision and reproducibility. The American Statistician, 65(4):213–221, November 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Blades:2015:SCS

[BSC15] Natalie J. Blades, G. Bruce Schaalje, and William F. Christensen. The second course in statistics: Design and analysis of experiments? The American Statistician, 69(4):326–333, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Baldi:2015:WYF

[BU15] Brigitte Baldi and Jessica Utts. What your future doctor should know about statistics: Must-include topics for introductory undergraduate biostatistics. The American Statistician, 69(3):231– 240, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00031305.2015.1048903.

Begg:2011:BSP

[BV11] Melissa D. Begg and Roger D. Vaughan. Are biostatistics students prepared to succeed in the era of interdisciplinary science? (and how will we know?). The American Statistician, 65(2):71–79, May 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Broman:2018:DOS

[BW18] Karl W. Broman and Kara H. Woo. Data organization in spreadsheets. The American Statistician, 72(1):2–10, 2018. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2017.1375989.

Baker:2014:VBM

[BXHH14] Stuart Baker, Jian-Lun Xu, Ping Hu, and Peng Huang. Vardeman, S. B. and Morris, M. D. (2013), “Majority Voting by Independent Classifiers can Increase Error Rates,” The American Statistician, 67, 94–96: Comment by Baker, Xu, Hu, and Huang and Reply. The American Statistician, 68(2):125–126, 2014. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See [VM13].

Carnegie:2019:BRQ

[Car19] Nicole Bohme Carnegie. Book review: Quantitative Methods for HIV/AIDS Research. The American Statistician, 73(2):209–210, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2019.1603473.

Chi:2016:PMM

[CCB16] Jocelyn T. Chi, Eric C. Chi, and Richard G. Baraniuk. k-POD: A method for k-means clustering of missing data. The American Statistician, 70(1):91–99, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cerqueira:2018:GS

[CCFJ18] Daniel Cerqueira, Danilo Coelho, Marcelo Fernandes, and Jony Pinto Junior. Guns and suicides. The American Statistician, 72(3):289–294, 2018. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/00031305.2017.1419144.

Cheng:2018:ICA

[CFP18] Sherri Cheng, Mark Ferris, and Jessica Perolio. An innovative classroom approach for developing critical thinkers in the introductory statistics course. The American Statistician, 72(4):354– 358, 2018. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://amstat.tandfonline.com/doi/full/
10.1080/00031305.2017.1305293.

Cook:2015:LE

[CFR15] R. Dennis Cook, Liliana Forzani, and Adam Rothman. Letter to the Editor. The American Statistician, 69(3):253–254, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00031305.2015.1053522.

Campbell:2019:WRG

[CG19] Harlan Campbell and Paul Gustafson. The world of research has gone berserk: Modeling the consequences of requiring “Greater statistical stringency” for scientific publication. The American Statistician, 73(S1):358–373, 2019. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2018.1555101.

Chi:2019:PSS

[CGM19] Yueh-Yun Chi, Deborah H. Glueck, and Keith E. Muller. Power and sample size for fixed-effects inference in reversible linear mixed models. The American Statistician, 73(4):350–359, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2017.1415972.

Chen:2011:DCS

[Che11] Yung-Pin Chen. Do the chi-square test and Fisher’s exact test agree in determining extreme for 2 × 2 tables? The American Statistician, 65(4):239–245, November 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Christensen:2014:C

[Chr14] Ronald Christensen. Comment. The American Statistician, 68(1): 13–17, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Christensen:2017:CTP

[Chr17] Ronald Christensen. Comment on “The Target Parameter of Adjusted R-Squared in Fixed-Design Experiments” by Bar-Gera (2017). The American Statistician, 71(4):373–375, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2017.1358215. See [BG17].

Christensen:2018:CNC

[Chr18a] Ronald Christensen. Comment on “A Note on Collinearity Diagnostics and Centering” by Velilla (2018). The American Statistician, 72(1):114–117, 2018. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://amstat.

tandfonline.com/doi/full/10.1080/00031305.2017.1392896.

See [Vel18a] and reply [Vel18b].

Christensen:2018:CKD

[Chr18b]  Ronald Christensen.  Comment on Knaeble and Dutter

(2017). The American Statistician, 72(4):392–393, 2018. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2016.1278036. See [KD17].

Christensen:2020:CTT

[Chr20] Ronald Christensen. Comment on “Test for Trend With a Multinomial Outcome” by Szabo (2019). The American Statistician, 74(3):313–314, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00031305.2020.1763835. See [Sza19].

Calin-Jageman:2019:NSB

[CJC19] Robert J. Calin-Jageman and Geoff Cumming. The new statistics for better science: Ask how much, how uncertain, and what else is known. The American Statistician, 73(S1):271–280, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2018.1518266.

Chakraborti:2019:HOM

[CJE19] Subhabrata Chakraborti, Felipe Jardim, and Eugenio Epprecht. Higher-order moments using the survival function: The alternative expectation formula. The American Statistician, 73(2):191– 194, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://amstat.tandfonline.com/doi/full/
10.1080/00031305.2017.1356374.

Caffo:2013:CSM

[CLR13] Brian Caffo, Carolyn Lauzon, and Joachim Röhmel. Correction to “Easy Multiplicity Control in Equivalence Testing Using Two One-Sided Tests”. The American Statistician, 67(2):115– 116, 2013. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See [LC09].

Chamandy:2015:TSG

[CMW15] Nicholas Chamandy, Omkar Muralidharan, and Stefan Wager. Teaching statistics at Google-scale. The American Statistician, 69(4):283–291, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cobb:2015:MRT

[Cob15] George Cobb. Mere renovation is too little too late: We need to rethink our undergraduate curriculum from the ground up. The American Statistician, 69(4):266–282, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cochran:2015:APE

[Coc15] James J. Cochran. ASA Presidents and Executive Directors look back on their terms in office. The American Statistician, 69(2): 79–85, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cohen:2010:LE

[Coh10] Michael P. Cohen. Letter to the Editor. The American Statistician, 64(2):192, May 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cohen:2015:MIC

[Coh15] Joel E. Cohen. Markov’s inequality and Chebyshev’s inequality for tail probabilities: A sharper image. The American Statistician, 69(1):5–7, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cohen:2016:SPP

[Coh16] Joel E. Cohen. Statistics of primes (and probably twin primes) satisfy Taylor’s law from ecology. The American Statistician, 70 (4):399–404, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cohen:2017:NAM

[Coh17] Michael P. Cohen. Non-asymptotic mean and variance also approximately satisfy Taylor’s law. The American Statistician, 71 (2):187, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://amstat.tandfonline.com/doi/
full/10.1080/00031305.2017.1286261.

Cohen:2019:SRN

[Coh19] Joel E. Cohen. Sum of a random number of correlated random variables that depend on the number of summands. The American Statistician, 73(1):56–60, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://amstat.tandfonline.com/doi/full/10.1080/
00031305.2017.1311283.

Colquhoun:2019:FPR

[Col19] David Colquhoun. The false positive risk: A proposal concerning what to do about p-values. The American Statistician, 73(S1):

192–201, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://amstat.tandfonline.com/doi/
full/10.1080/00031305.2018.1529622.

Chance:2015:CGL

[CP15] Beth Chance and Roxy Peck. From curriculum guidelines to learning outcomes: Assessment at the program level. The American Statistician, 69(4):409–416, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cetinkaya-Rundel:2018:ITT

[ÇRR18] Mine Çetinkaya-Rundel and Colin Rundel. Infrastructure and tools for teaching computing throughout the statistical curriculum. The American Statistician, 72(1):58–65, 2018. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2017.1397549.

Chuang-Stein:2015:CAA

[CSBB^{+}15] Christy Chuang-Stein, Narayanaswamy Balakrishnan, Marcus Berzofsky, Amy Herring, Fred Hulting, John McKenzie, Dionne Price, Stephen Stigler, George Williams, and Ronald Wasserstein. Celebrating the 175th anniversary of ASA. The American Statistician, 69(2):64–67, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cunningham:2012:SGM

[Cun12] J. Kelly Cunningham. Should S get more press? The American Statistician, 66(4):237, 2012. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cleary:2010:BDF

[CW10] Rick Cleary and Sam Woolford. The business of desire and fear. The American Statistician, 64(1):21–22, February 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cron:2011:ECB

[CW11] Andrew J. Cron and Mike West. Efficient classification-based relabeling in mixture models. The American Statistician, 65(1):16– 20, February 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Derryberry:2018:MSR

[DAEP18] DeWayne Derryberry, Ken Aho, John Edwards, and Teri Peterson. Model selection and regression t-statistics. The American Statistician, 72(4):379–381, 2018. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://amstat.

tandfonline.com/doi/full/10.1080/00031305.2018.1459316.

David:2011:ECM

[Dav11] Herbert A. David. Euler’s contributions to mathematics useful in statistics. The American Statistician, 65(1):37–42, February 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Ding:2018:GMR

[DB18] Peng Ding and Joseph K. Blitzstein. On the Gaussian mixture representation of the Laplace distribution. The American Statistician, 72(2):172–174, 2018. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://amstat.

tandfonline.com/doi/full/10.1080/00031305.2017.1291448.

deCarvalho:2016:MWD

[dC16] Miguel de Carvalho. Mean, what do you mean? The American Statistician, 70(3):270–??, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Dunn:2017:IST

[DCF^{+}17] Peter K. Dunn, Michael D. Carey, Michael B. Farrar, Alice M. Richardson, and Christine McDonald. Introductory statistics textbooks and the GAISE recommendations. The American Statistician, 71(4):326–335, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://amstat.tandfonline.com/doi/full/10.1080/
00031305.2016.1251972.

Davidson:2019:TCS

[DDF19] Mario A. Davidson, Charlene M. Dewey, and Amy E. Fleming. Teaching communication in a statistical collaboration course: A feasible, project-based, multimodal curriculum. The American Statistician, 73(1):61–69, 2019. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2018.1448890.

Diniz:2016:CDP

[DDV16] Marcio A. Diniz, Jasper De Bock, and Arthur Van Camp. Characterizing Dirichlet priors. The American Statistician, 70(1):9– 17, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Demidenko:2016:VYC

[Dem16a] Eugene Demidenko. The p-value you can’t buy. The American Statistician, 70(1):33–38, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Demirtas:2016:NRB

[Dem16b] Hakan Demirtas. A note on the relationship between the phi coefficient and the tetrachoric correlation under nonnormal underlying distributions. The American Statistician, 70(2):143–148, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Demers:2018:TLH

[Dem18] Simon Demers. Taylor’s law holds for finite OEIS integer sequences and binomial coefficients. The American Statistician, 72 (4):376–378, 2018. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/00031305.2017.1422439.

Demirtas:2019:IFL

[Dem19] Hakan Demirtas. Inducing any feasible level of correlation to bivariate data with any marginals. The American Statistician, 73 (3):273–277, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/00031305.2017.1379438.

Diaz-Frances:2016:SEI

[DF16] Eloísa Díaz-Francés. Simple estimation intervals for Poisson, exponential, and inverse Gaussian means obtained by symmetrizing the likelihood function. The American Statistician, 70(2):171– 180, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Demirtas:2011:PWC

[DH11] Hakan Demirtas and Donald Hedeker. A practical way for computing approximate lower and upper correlation bounds. The American Statistician, 65(2):104–109, May 2011. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Du:2011:PUL

[DHW11] Yeting Du, Christopher Hundt, and David B. Wolfson. Previously unrecognized links between statistical dependence and some common modes of convergence. The American Statistician, 65(1):55– 60, February 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Ding:2014:CTO

[Din14a] Peng Ding. Comment on Tarpey, T., Ogden, R. T., Petkova, E., and Christensen R. (2014), “A Paradoxical Result in Estimating Regression Coefficients,” The American Statistician, 68, 271–276. The American Statistician, 68(4):316, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See [TOPC14].

Ding:2014:TOH

[Din14b] Peng Ding. Three occurrences of the hyperbolic-secant distribution. The American Statistician, 68(1):32–35, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Ding:2015:R

[Din15] Peng Ding. Reply. The American Statistician, 69(3):255–256, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00031305.2015.1056615.

Ding:2016:CDM

[Din16] Peng Ding. On the conditional distribution of the multivariate t distribution. The American Statistician, 70(3):293–??, 2016. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

DiCiccio:2017:SAE

[DKY17] Thomas J. DiCiccio, Todd A. Kuffner, and G. Alastair Young. A simple analysis of the exact probability matching prior in the location-scale model. The American Statistician, 71(4):302–304, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2016.1255662.

Dunn:2019:EWS

[DMM19] Peter K. Dunn, Margaret Marshman, and Robert McDougall. Evaluating Wikipedia as a self-learning resource for statistics: You know they’ll use it. The American Statistician, 73(3):224– 231, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://amstat.tandfonline.com/doi/full/
10.1080/00031305.2017.1392360.

Divine:2018:WMW

[DNBJC18] George W. Divine, H. James Norton, Anna E. Barón, and Elizabeth Juarez-Colunga. The Wilcoxon–Mann–Whitney procedure fails as a test of medians. The American Statistician, 72(3):278– 286, 2018. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://amstat.tandfonline.com/doi/full/
10.1080/00031305.2017.1305291.

Dressler:2019:BRC

[Dre19] Emily Dressler. Book review: Clinical Trial Optimization Using R. The American Statistician, 73(2):210–211, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2019.1603479.

Quintela-del-Rio:2017:ETH

[dRFF17] Alejandro Quintela del Río and Mario Francisco-Fernández. Excel templates: A helpful tool for teaching statistics. The American Statistician, 71(4):317–325, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://amstat.

tandfonline.com/doi/full/10.1080/00031305.2016.1186115.

Pereira:2017:IGA

[dSPCdS17] Leandro da Silva Pereira, Lucas Monteiro Chaves, and Devanil Jaques de Souza. An intuitive geometric approach to the Gauss Markov Theorem. The American Statistician, 71(1):67– 70, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

DeMets:2015:IBN

[DWG15] David L. DeMets, Janet Turk Wittes, and Nancy L. Geller. The influence of biostatistics at the national heart, lung, and blood institute. The American Statistician, 69(2):108–120, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Emerson:2011:SSL

[EA11] John W. Emerson and Taylor B. Arnold. Statistical sleuthing by leveraging human nature: a study of Olympic figure skating. The American Statistician, 65(3):143–148, August 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Easterling:2010:PDS

[Eas10] Robert G. Easterling. Passion-driven statistics. The American Statistician, 64(1):1–5, February 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Easterling:2015:TNW

[Eas15] Robert Easterling. There’s nothing wrong with Clopper–Pearson binomial confidence limits. The American Statistician, 69(2):154– 155, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Eddelbuettel:2018:ERC

[EB18] Dirk Eddelbuettel and James Joseph Balamuta. Extending R with C++: A brief introduction to Rcpp. The American Statistician, 72(1):28–36, 2018. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/00031305.2017.1375990.

Ekin:2017:UCF

[EIRS17] Tahir Ekin, Francesca Ieva, Fabrizio Ruggeri, and Refik Soyer. On the use of the concentration function in medical fraud assessment. The American Statistician, 71(3):236–241, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2017.1292955.

Esteves:2017:TDT

[EIS17] Luís Gustavo Esteves, Rafael Izbicki, and Rafael Bassi Stern. Teaching decision theory proof strategies using a crowdsourcing problem. The American Statistician, 71(4):336–343, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2016.1264316.

Ellis:2018:HSD

[EL18] Shannon E. Ellis and Jeffrey T. Leek. How to share data for collaboration. The American Statistician, 72(1):53–57, 2018.

CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2017.1375987.

Espinosa:2015:DMI

[ER15] Valeria Espinosa and Donald B. Rubin. Did the military interventions in the Mexican Drug War increase violence? The American Statistician, 69(1):17–27, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Elliott:2018:TES

[ESC18] Alan C. Elliott, S. Lynne Stokes, and Jing Cao. Teaching ethics in a statistics curriculum with a cross-cultural emphasis. The American Statistician, 72(4):359–367, 2018. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2017.1307140.

Fricker:2019:ASA

[FBHW19] Ronald D. Fricker Jr., Katherine Burke, Xiaoyan Han, and William H. Woodall. Assessing the statistical analyses used in Basic and Applied Social Psychology after their p-value ban. The American Statistician, 73(S1):374–384, 2019. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2018.1537892.

Friendly:2012:AMG

[Fd12] Michael Friendly and Nicolas de Sainte Agathe. André-Michel Guerry’s Ordonnateur Statistique: The first statistical calculator? The American Statistician, 66(3):195–200, 2012. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Fellingham:2018:PHR

[FF18] Gilbert W. Fellingham and Jared D. Fisher. Predicting home run production in major league baseball using a Bayesian semiparametric model. The American Statistician, 72(3):253–264, 2018. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2017.1401959.

Fontdecaba:2014:ADS

[FGTM14] Sara Fontdecaba, Pere Grima, and Xavier Tort-Martorell. Analyzing DOE with statistical software packages: Controversies and proposals. The American Statistician, 68(3):205–211, 2014. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Fong:2019:MWM

[FH19] Youyi Fong and Ying Huang. Modified Wilcoxon–Mann–Whitney test and power against strong null. The American Statistician, 73(1):43–49, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/00031305.2017.1328375.

Fogel:2013:TTM

[FHB^{+}13] Paul Fogel, Douglas M. Hawkins, Chris Beecher, George Luta, and S. Stanley Young. A tale of two matrix factorizations. The American Statistician, 67(4):207–218, 2013. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Fader:2019:EET

[FHMV19] Peter S. Fader, Bruce G. S. Hardie, Daniel McCarthy, and Ramnath Vaidyanathan. Exploring the equivalence of two common mixture models for duration data. The American Statistician, 73 (3):288–295, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/00031305.2018.1543134.

Franklin:2011:ASB

[FHP^{+}11] Christine Franklin, Brad Hartlaub, Roxy Peck, Richard Scheaffer, David Thiel, and Katherine Tranbarger Freier. AP statistics: Building bridges between high school and college statistics education. The American Statistician, 65(3):177–182, August 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Fienberg:2013:CBI

[Fie13a] Stephen E. Fienberg. Comment: Bayesian ideas reemerged in the 1950s. The American Statistician, 67(1):7–8, 2013. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Fienberg:2013:CIA

[Fie13b] Stephen E. Fienberg. Comment: Innovations associated with multiple systems estimation in human rights settings. The American Statistician, 67(4):201–202, 2013. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Fossaluza:2017:CHT

[FIME17] Victor Fossaluza, Rafael Izbicki, Gustavo Miranda da Silva, and Luís Gustavo Esteves. Coherent hypothesis testing. The American Statistician, 71(3):242–248, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://amstat.tandfonline.com/doi/full/10.1080/
00031305.2016.1237893.

Fairley:2017:RMM

[FKN^{+}17] William B. Fairley, Peter J. Kempthorne, Julie Novak, Scott McGarvie, Steve Crunk, Bee Leng Lee, and Alan J. Salzberg. Resolving a multi-million dollar contract dispute with a latin square. The American Statistician, 71(3):249–258, 2017. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2016.1256231.

Foster:2011:SSR

[Fos11] Colin Foster. The significance of a square-root rule. The American Statistician, 65(4):222, November 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Foulley:2020:BDJ

[Fou20] Jean-Louis Foulley. Benjamin, D. J., and Berger, J. O. (2019), “Three Recommendations for Improving the Use of p-Values”, The American Statistician, 73, 186–191: Comment by Foulley. The American Statistician, 74(1):101–102, 2020. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2019.1668850. See [BB19].

Fox:2010:DFQ

[Fox10] David R. Fox. Desired and feared— quo vadis or quid agis? The American Statistician, 64(1):6–9, February 2010. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Frey:2012:EBC

[FP12] Jesse Frey and Andrés Pérez. Exact binomial confidence intervals for randomized response. The American Statistician, 66(1):8– 15, 2012. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Field:2012:BNG

[FPW12] C. A. Field, Zhen Pang, and A. H. Welsh. On the boundedness and nonmonotonicity of generalized score statistics. The American Statistician, 66(2):92–98, 2012. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://amstat.

tandfonline.com/doi/full/10.1080/00031305.2012.703888.

Fosdick:2012:ECB

[FR12] Bailey K. Fosdick and Adrian E. Raftery. Estimating the correlation in bivariate normal data with known variances and small sample sizes. The American Statistician, 66(1):34–41, 2012. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Fisher:2019:CTI

[FR19] Thomas J. Fisher and Michael W. Robbins. A cheap trick to improve the power of a conservative hypothesis test. The American Statistician, 73(3):232–242, 2019. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2017.1395364.

Fraser:2019:VFS

[Fra19] D. A. S. Fraser. The p-value function and statistical inference. The American Statistician, 73(S1):135–147, 2019. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2018.1556735.

Frey:2010:FWS

[Fre10] Jesse Frey. Fixed-width sequential confidence intervals for a proportion. The American Statistician, 64(3):242–249, August 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Frey:2019:CV

[Fre19] Jesse Frey. Comment on VanDerwerken (2019). The American Statistician, 73(4):411–412, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00031305.2019.
1604433.

Friedman:2015:NCC

[Fri15] Emil M. Friedman. Nontransitivity, correlation, and causation: Langford, E., Schwertman, N., and Owens, M. (2001), “Is the Property of Being Positively Correlated Transitive?” The American Statistician, 55, 322–325: Comment by Friedman. The American Statistician, 69(3):257, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/00031305.2015.1056382.

See [LSO01].

From:2012:CMF

[Fro12] Steven G. From. A comparison of the moment and factorial moment bounds for discrete random variables. The American Statistician, 66(4):214–216, 2012. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Fearnhead:2010:CSS

[FT10] Paul Fearnhead and Benjamin M. Taylor. Calculating strength of schedule, and choosing teams for march madness. The American Statistician, 64(2):108–115, May 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Friendly:2010:FKS

[FVMU10] Michael Friendly, Pedro Valero-Mora, and Joaquín Ibáñez Ulargui. The first (known) statistical graph: Michael Florent van Langren and the “Secret” of Longitude. The American Statistician, 64(2):185–191, May 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Feng:2013:MVT

[FWH^{+}13] Changyong Feng, Hongyue Wang, Yu Han, Yinglin Xia, and Xin M. Tu. The Mean Value Theorem and Taylor’s expansion in statistics. The American Statistician, 67(4):245–248, 2013. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Feng:2014:R

[FWH^{+}14] Changyong Feng, Hongyue Wang, Yu Han, Yinglin Xia, and Xin M. Tu. Reply. The American Statistician, 68(3):220, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Fan:2017:CTT

[FWW17] Chunpeng Fan, Lin Wang, and Lynn Wei. Comparing two tests for two rates. The American Statistician, 71(3):275–281, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2016.1246263.

Frey:2017:WDI

[FZ17] Jesse Frey and Yimin Zhang. What do interpolated nonparametric confidence intervals for population quantiles guarantee? The American Statistician, 71(4):305–309, 2017. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2016.1226952. See comment [Hut18].

Gitlow:2013:ISS

[GA13] Howard Gitlow and Hernan Awad. Intro stats students need both confidence and tolerance (intervals). The American Statistician, 67(4):229–234, 2013. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gregoire:2018:EDS

[GA18] Timothy G. Gregoire and David L. R. Affleck. Estimating desired sample size for simple random sampling of a skewed population. The American Statistician, 72(2):184–190, 2018. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2017.1290548.

Green:2015:FCU

[GB15] Jennifer L. Green and Erin E. Blankenship. Fostering conceptual understanding in mathematical statistics. The American Statistician, 69(4):315–325, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Goh:2019:APM

[GD19] Gyuhyeong Goh and Dipak K. Dey. Asymptotic properties of marginal least-square estimator for ultrahigh-dimensional linear regression models with correlated errors. The American Statistician, 73(1):4–9, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/00031305.2017.1302359.

Gannon:2019:BBC

[GdBPP19] Mark Andrew Gannon, Carlos Alberto de Bragança Pereira, and Adriano Polpo. Blending Bayesian and classical tools to define optimal sample-size-dependent significance levels. The American Statistician, 73(S1):213–222, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://amstat.tandfonline.com/doi/full/10.1080/
00031305.2018.1518268.

Gilliland:2011:URC

[GE11] Dennis Gilliland and Don Edwards. Using randomized confidence limits to balance risk: An application to Medicare investigations. The American Statistician, 65(3):149–153, August 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Ghosh:2015:BVS

[GG15] Joyee Ghosh and Andrew E. Ghattas. Bayesian variable selection under collinearity. The American Statistician, 69(3):165– 173, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00031305.2015.1031827.

Gelman:2019:RSB

[GGGV19] Andrew Gelman, Ben Goodrich, Jonah Gabry, and Aki Vehtari. R-squared for Bayesian regression models. The American Statistician, 73(3):307–309, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://amstat.

tandfonline.com/doi/full/10.1080/00031305.2018.1549100.

Gibson:2019:LSI

[Gib19] Eric W. Gibson. Leadership in statistics: Increasing our value and visibility. The American Statistician, 73(2):109–116, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2017.1336484.

Gonen:2019:COS

[GJLW19] Mithat Gönen, Wesley O. Johnson, Yonggang Lu, and Peter H. Westfall. Comparing objective and subjective Bayes factors for the two-sample comparison: The classification theorem in action. The American Statistician, 73(1):22–31, 2019.

CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2017.1322142.

Garcia:2012:RMC

[GKP12] Tanya P. Garcia, Priya Kohli, and Mohsen Pourahmadi. Regressograms and mean-covariance models for incomplete longitudinal data. The American Statistician, 66(2):85–91, 2012. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Galili:2016:EIR

[GM16] Tal Galili and Isaac Meilijson. An example of an improvable Rao–Blackwell improvement, inefficient maximum likelihood estimator, and unbiased generalized Bayes estimator. The American Statistician, 70(1):108–113, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Golbeck:2017:MFW

[Gol17] Amanda L. Golbeck. Mentoring faculty women in statistics: Exploring challenges and opportunities for leadership development. The American Statistician, 71(1):47–54, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gonen:2013:VLD

[Gön13] Mithat Gönen. Visualizing longitudinal data with dropouts. The American Statistician, 67(2):97–103, 2013. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Goodman:2019:WGR

[Goo19] Steven N. Goodman. Why is getting rid of P -values so hard? Musings on science and statistics. The American Statistician, 73(S1):26–30, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/00031305.2018.1558111.

Gorroochurn:2011:EPH

[Gor11] Prakash Gorroochurn. Errors of probability in historical context. The American Statistician, 65(4):246–254, November 2011. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gorroochurn:2016:GCF

[Gor16] Prakash Gorroochurn. On Galton’s change from “Reversion” to “Regression”. The American Statistician, 70(3):227–??, 2016.

CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Greselin:2013:CLR

[GP13] Francesca Greselin and Antonio Punzo. Closed likelihood ratio testing procedures to assess similarity of covariance matrices. The American Statistician, 67(3):117–128, 2013. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gelman:2013:SOD

[GR13a] Andrew Gelman and Christian P. Robert. “Not Only Defended But Also Applied”: The perceived absurdity of Bayesian inference. The American Statistician, 67(1):1–5, 2013. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gelman:2013:RAB

[GR13b] Andrew Gelman and Christian P. Robert. Rejoinder: The anti-Bayesian moment and its passing. The American Statistician, 67(1):16–17, 2013. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Goldin:2019:ASD

[GR19] Jacob Goldin and Daniel Reck. The analysis of survey data with framing effects. The American Statistician, 73(3):264–272, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2017.1407358.

Greenland:2010:SPA

[Gre10] Sander Greenland. Simpson’s Paradox from adding constants in contingency tables as an example of Bayesian noncollapsibility. The American Statistician, 64(4):340–344, November 2010. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Greenland:2019:VVB

[Gre19] Sander Greenland. Valid P -values behave exactly as they should: Some misleading criticisms of P -values and their resolution with S-values. The American Statistician, 73(S1):106–114, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2018.1529625.

Griffith:2013:BAN

[Gri13] Daniel A. Griffith. Better articulating normal curve theory for introductory mathematical statistics students: Power transformations and their back-transformations. The American Statistician, 67(3):157–169, 2013. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See comment [Muk14b] and reply [Gri14].

Griffith:2014:R

[Gri14] Daniel A. Griffith. Reply. The American Statistician, 68(1):67– 69, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See [Gri13, Muk14b].

Grimshaw:2015:FIA

[Gri15] Scott D. Grimshaw. A framework for infusing authentic data experiences within statistics courses. The American Statistician, 69(4):307–314, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gromping:2014:MPU

[Grö14] Ulrike Grömping. Mosaic plots are useful for visualizing low-order projections of factorial designs. The American Statistician, 68(2): 108–116, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Grima:2016:EIV

[GRTM16] Pere Grima, Lourdes Rodero, and Xavier Tort-Martorell. Explaining the importance of variability to engineering students. The American Statistician, 70(2):138–142, 2016. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Guerra:2014:NSO

[GS14] Matthew W. Guerra and Justine Shults. A note on the simulation of overdispersed random variables with specified marginal means and product correlations. The American Statistician, 68(2):104– 107, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Greenhouse:2018:TSP

[GS18] Joel B. Greenhouse and Howard J. Seltman. On teaching statistical practice: From novice to expert. The American Statistician, 72(2):147–154, 2018. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/00031305.2016.1270230.

Goodman:2019:PHE

[GSK19] William M. Goodman, Susan E. Spruill, and Eugene Komaroff. A proposed hybrid effect size plus p-value criterion: Empirical evidence supporting its use. The American Statistician, 73(S1): 168–185, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://amstat.tandfonline.com/doi/
full/10.1080/00031305.2018.1564697.

Guilbaud:2018:SCH

[Gui18] Olivier J. M. Guilbaud. Some complementary history and results. The American Statistician, 72(3):300–301, 2018. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2018.1448892.

Gou:2017:ESP

[GZ17] Jiangtao Gou and Fengqing (Zoe) Zhang. Experience Simpson’s paradox in the classroom. The American Statistician, 71(1):61– 66, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Garfield:2011:RAS

[GZK^{+}11] Joan Garfield, Andrew Zieffler, Daniel Kaplan, George W. Cobb, Beth L. Chance, and John P. Holcomb. Rethinking assessment of student learning in statistics courses. The American Statistician, 65(1):1–10, February 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hall:2010:RFG

[Hal10] Nancy S. Hall. Ronald Fisher and Gertrude Cox: Two statistical pioneers sometimes cooperate and sometimes collide. The American Statistician, 64(3):212–220, August 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Harville:2014:NME

[Har14a] David A. Harville. The need for more emphasis on prediction: A “Nondenominational” model-based approach. The American Statistician, 68(2):71–83, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See comments [Ste14, Zim14, McC14, BB14] and rejoinder [Har14b].

Harville:2014:R

[Har14b] David A. Harville. Rejoinder. The American Statistician, 68(2): 89–92, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See [Har14a].

Harms:2019:BFR

[Har19] Christopher Harms. A Bayes factor for replications of ANOVA results. The American Statistician, 73(4):327–339, 2019. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2018.1518787.

Hayes:2011:GIA

[Hay11] Kevin Hayes. A geometrical interpretation of an alternative formula for the sample covariance. The American Statistician, 65(2): 110–112, May 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hayter:2014:SCI

[Hay14] A. J. Hayter. Simultaneous confidence intervals for several quantiles of an unknown distribution. The American Statistician, 68 (1):56–62, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hubbard:2019:QCS

[HC19] Douglas W. Hubbard and Alicia L. Carriquiry. Quality control for scientific research: Addressing reproducibility, responsiveness, and relevance. The American Statistician, 73(S1):46– 55, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://amstat.tandfonline.com/doi/full/
10.1080/00031305.2018.1543138.

Harper:2011:CAM

[HEJ11] William V. Harper, Ted G. Eschenbach, and Thomas R. James. Concerns about maximum likelihood estimation for the threeparameter Weibull distribution: Case study of statistical software. The American Statistician, 65(1):44–54, February 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hesterberg:2015:WTS

[Hes15] Tim C. Hesterberg. What teachers should know about the bootstrap: Resampling in the undergraduate statistics curriculum.

The American Statistician, 69(4):371–386, 2015. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hurwitz:2012:EGS

[HG12] Shelley Hurwitz and John S. Gardenier. Ethical guidelines for statistical practice: The first 60 years and beyond. The American Statistician, 66(2):99–103, 2012. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Horton:2015:TNG

[HH15] Nicholas J. Horton and Johanna S. Hardin. Teaching the next generation of statistics students to “Think with data”: Special issue on statistics and the undergraduate curriculum. The American Statistician, 69(4):259–265, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hwang:2017:NIB

[HHC17] Wen-Han Hwang, Richard Huggins, and Lu-Fang Chen. A note on the inverse birthday problem with applications. The American Statistician, 71(3):191–201, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://amstat.tandfonline.com/doi/full/10.1080/
00031305.2016.1255657.

Hardin:2015:DSS

[HHH^{+}15] J. Hardin, R. Hoerl, Nicholas J. Horton, D. Nolan, B. Baumer, O. Hall-Holt, P. Murrell, R. Peng, P. Roback, D. Temple Lang, and M. D. Ward. Data science in statistics curricula: Preparing students to “Think with data”. The American Statistician, 69 (4):343–353, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hubbard:2019:LRF

[HHP19] Raymond Hubbard, Brian D. Haig, and Rahul A. Parsa. The limited role of formal statistical inference in scientific inference. The American Statistician, 73(S1):91–98, 2019. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2018.1464947.

Hicks:2018:GTD

[HI18] Stephanie C. Hicks and Rafael A. Irizarry. A guide to teaching data science. The American Statistician, 72(4):382–391, 2018.

CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2017.1356747.

Higgs:2019:BRR

[Hig19] Megan D. Higgs. Book review: Randomistas: How Radical Researchers Are Changing Our World. The American Statistician, 73(4):416–417, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2019.1676111.

Hill:2011:LEB

[Hil11] Theodore P. Hill. Letter to the editor: Benford’s Law blunders. The American Statistician, 65(2):141, May 2011. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.tandfonline.com/doi/pdf/10.1198/tast.
2011.11011.

Hoshyarmanesh:2016:CIS

[HKM16] Hossein Hoshyarmanesh, Amirhossein Karami, and Adel Mohammadpour. Confidence intervals for the scale parameter of exponential family of distributions. The American Statistician, 70(2): 134–137, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Harwell:2018:SRP

[HKPT18] Michael Harwell, Nidhi Kohli, and Yadira Peralta-Torres. A survey of reporting practices of computer simulation studies in statistical research. The American Statistician, 72(4):321–327, 2018. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2017.1342692.

Hirschberg:2010:GCD

[HL10] Joe Hirschberg and Jenny Lye. A geometric comparison of the delta and Fieller confidence intervals. The American Statistician, 64(3):234–241, August 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hughes:2015:NTS

[HL15] Timothy A. C. Hughes and Jaechoul Lee. A new test for short memory in long memory time series. The American Statistician, 69(3):182–190, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00031305.2015.1056829.

Hurlbert:2019:CGT

[HLU19] Stuart H. Hurlbert, Richard A. Levine, and Jessica Utts. Coup de grâce for a tough old bull: “Statistically significant” expires. The American Statistician, 73(S1):352–357, 2019. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2018.1543616.

Hu:2013:IPD

[HMR13] Jingchen Hu, Robin Mitra, and Jerome Reiter. Are independent parameter draws necessary for multiple imputation? The American Statistician, 67(3):143–149, 2013. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hogbin:2010:LE

[HN10] Martin Hogbin and W. Nijdam. Letter to the Editor. The American Statistician, 64(2):193–194, May 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Held:2016:HME

[HO16] Leonhard Held and Manuela Ott. How the maximal evidence of P - values against point null hypotheses depends on sample size. The American Statistician, 70(4):335–341, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hughes-Oliver:2017:MAD

[HO17] Jacqueline M. Hughes-Oliver. Mentoring to achieve diversity in graduate programs. The American Statistician, 71(1):55–60, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hodge:2017:LEA

[Hod17] Susan E. Hodge. Letter to the Editor: Average entropy does not measure uncertainty. The American Statistician, 71(1):89– 90, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See response [Kit17].

Hoef:2012:WID

[Hoe12] Jay M. Ver Hoef. Who invented the Delta Method? The American Statistician, 66(2):124–127, 2012. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Holland:2014:STJ

[Hol14] Bart Holland. Segal, N. L., and Torres, J. (2013), “A Repeated Grammatical Error Does Not Make it Right,” The American Statistician, 67, 266: Comment by Holland and Reply. The American Statistician, 68(2):127, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See [ST13].

Hong:2012:RAE

[Hon12] Liang Hong. A remark on the alternative expectation formula. The American Statistician, 66(4):232–233, 2012. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hong:2014:LE

[Hon14a] Liang Hong. Letter to the Editor. The American Statistician, 68 (3):220, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hong:2014:TNE

[Hon14b] Liang Hong. Two new elementary derivations of geometric expectation. The American Statistician, 68(3):188–190, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hong:2015:ADL

[Hon15a] Liang Hong. The absolute difference law for expectations. The American Statistician, 69(1):8–10, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hong:2015:ARA

[Hon15b] Liang Hong. Another remark on the alternative expectation formula. The American Statistician, 69(3):157–159, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic).  URL http://www.tandfonline.com/doi/abs/10.
1080/00031305.2015.1049710.

Horton:2013:HFD

[Hor13] Nicholas J. Horton. I hear, I forget. I do, I understand: A modified Moore-method mathematical statistics course. The American Statistician, 67(4):219–228, 2013. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Horton:2015:COS

[Hor15] Nicholas J. Horton. Challenges and opportunities for statistics and statistical education: Looking back, looking forward. The American Statistician, 69(2):138–145, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hodges:2010:ASC

[HR10] James S. Hodges and Brian J. Reich. Adding spatially-correlated errors can mess up the fixed effect you love. The American Statistician, 64(4):325–334, November 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Heller:2010:UCM

[HRS10] Ruth Heller, Paul R. Rosenbaum, and Dylan S. Small. Using the cross-match test to appraise covariate balance in matched pairs. The American Statistician, 64(4):299–309, November 2010. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hoerl:2010:MSP

[HS10] Roger W. Hoerl and Ronald D. Snee. Moving the statistics profession forward to the next level. The American Statistician, 64 (1):10–14, February 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hedayat:2012:RSE

[HS12] A. S. Hedayat and Guoqin Su. Robustness of the simultaneous estimators of location and scale from approximating a histogram by a normal density curve. The American Statistician, 66(1):25– 33, 2012. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hoerl:2017:SEI

[HS17] Roger W. Hoerl and Ronald D. Snee. Statistical engineering: An idea whose time has come? The American Statistician, 71 (3):209–219, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/00031305.2016.1247015.

Hubbard:2019:WAE

[Hub19] Raymond Hubbard. Will the ASA’s efforts to improve statistical practice be successful? Some evidence to the contrary. The American Statistician, 73(S1):31–35, 2019. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2018.1497540.

Hutson:2018:CWD

[Hut18] Alan Hutson. Comment on “What Do Interpolated Nonparametric Confidence Intervals for Population Quantiles Guarantee?”, Frey and Zhang (2017). The American Statistician, 72(3):302, 2018. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2018.1448893. See [FZ17].

Hutson:2018:CNB

[HV18] Alan D. Hutson and Albert Vexler. A cautionary note on beta families of distributions and the aliases within. The American Statistician, 72(2):121–129, 2018. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://amstat.tandfonline.com/doi/full/10.1080/
00031305.2016.1213661.

Hao:2017:NHD

[HZ17] Ning Hao and Hao Helen Zhang. A note on high-dimensional linear regression with interactions. The American Statistician, 71 (4):291–297, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/00031305.2016.1264311.

Ignatova:2012:CSM

[IDE12] Iliana Ignatova, Roland C. Deutsch, and Don Edwards. Closed sequential and multistage inference on binary responses with or without replacement. The American Statistician, 66(3):163–172, 2012. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See comment [MS14] and reply [IDE14].

Ignatova:2014:R

[IDE14] Lina Ignatova, Roland C. Deutsch, and Don Edwards. Reply. The American Statistician, 68(2):129, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See [IDE12].

Ignatova:2015:KJF

[IDE15] Iliana Ignatova, Roland Deutsch, and Don Edwards. Kirk, J. L., and Fay, M. P. “An Introduction to Practical Sequential Inferences Via Single-Arm Binary Response Studies Using the Binseqtest R Package,” The American Statistician, 68, 230–242. The American Statistician, 69(3):256–257, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00031305.2015.1053523. See [KF14].

Ionides:2017:RAS

[IGRP17] Edward L. Ionides, Alexander Giessing, Yaacov Ritov, and Scott E. Page. Response to the ASA’s statement on p-values: Context, process, and purpose. The American Statistician, 71(1): 88–89, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See [WL16b].

Inan:2014:KCK

[IIDdL14] Gul Inan, Ozlem Ilk-Dag, and Alexander de Leon. Kim, Y., Choi, Y.-K., and Emery, S. (2013), “Logistic Regression With Multiple Random Effects: A Simulation Study of Estimation Methods and Statistical Packages,” The American Statistician, 67, 171– 182. The American Statistician, 68(2):129–130, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See [KCE13].

Inlow:2010:MGF

[Inl10] Mark Inlow. A moment generating function proof of the Lindeberg–Lévy Central Limit Theorem. The American Statistician, 64(3):228–230, August 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Ioannidis:2019:WWL

[Ioa19] John P. A. Ioannidis. What have we (not) learnt from millions of scientific papers with P values? The American Statistician, 73(S1):20–25, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/00031305.2018.1447512.

Jaeger:2016:CTT

[Jae16] Adam Jaeger. Computation of two- and three-dimensional confidence regions with the likelihood ratio. The American Statistician, 70(4):395–398, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Jones:2014:PEI

[JJ14] Geoffrey Jones and Wesley O. Johnson. Prior elicitation: Interactive spreadsheet graphics with sliders can be fun, and informative. The American Statistician, 68(1):42–51, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Jeske:2018:RUP

[JM18] Daniel R. Jeske and Janet M. Myhre. Regression using pairs vs. regression on differences: A real-life case study for a Master’s level methods class. The American Statistician, 72(2):163– 168, 2018. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://amstat.tandfonline.com/doi/full/
10.1080/00031305.2017.1292956.

Jones:2019:IDI

[JMS19] M. C. Jones, Éric Marchand, and William E. Strawderman. On an intriguing distributional identity. The American Statistician, 73(1):16–21, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/00031305.2017.1375984.

Johnson:2013:CBS

[Joh13] Wesley O. Johnson. Comment: Bayesian statistics in the twenty first century. The American Statistician, 67(1):9–11, 2013. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Johnson:2019:EMS

[Joh19] Valen E. Johnson. Evidence from marginally significant t statistics. The American Statistician, 73(S1):129–134, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2018.1518788.

Jones:2019:LE

[Jon19] M. C. Jones. Letter to the Editor. The American Statistician, 73 (1):105, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://amstat.tandfonline.com/doi/
full/10.1080/00031305.2018.1556736.

Jones:2011:SIM

[JRP11] M. C. Jones, J. F. Rosco, and Arthur Pewsey. Skewness-invariant measures of kurtosis. The American Statistician, 65(2):89–95, May 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Jones:2011:IUH

[JS11] Hayley E. Jones and David J. Spiegelhalter. The identification of “Unusual” health-care providers from a hierarchical model. The American Statistician, 65(3):154–163, August 2011. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Jin:2017:ABF

[JTL17] Shaobo Jin, Måns Thulin, and Rolf Larsson. Approximate Bayesianity of frequentist confidence intervals for a binomial proportion. The American Statistician, 71(2):106–111, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2016.1208630.

Kadane:2013:C

[Kad13] Joseph B. Kadane. Comment. The American Statistician, 67 (4):202–203, 2013. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kaplan:2018:TSD

[Kap18] Daniel Kaplan. Teaching stats for data science. The American Statistician, 72(1):89–96, 2018. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://amstat.

tandfonline.com/doi/full/10.1080/00031305.2017.1398107.

Kim:2013:LRM

[KCE13] Yoonsang Kim, Young-Ku Choi, and Sherry Emery. Logistic regression with multiple random effects: A simulation study of estimation methods and statistical packages. The American Statistician, 67(3):171–182, 2013. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See comment [IIDdL14] and reply [KE14].

Knaeble:2017:RLS

[KD17] Brian Knaeble and Seth Dutter. Reversals of least-square estimates and model-invariant estimation for directions of unique effects. The American Statistician, 71(2):97–105, 2017. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2016.1226951. See comment [Chr18b].

Kim:2014:R

[KE14] Yoonsang Kim and Sherry Emery. Reply. The American Statistician, 68(2):130–131, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See [KCE13].

Kirk:2014:IPS

[KF14] Jennifer L. Kirk and Michael P. Fay. An introduction to practical sequential inferences via single-arm binary response studies using the binseqtest R package. The American Statistician, 68(4): 230–242, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See comments [IDE15].

Krueger:2019:PVP

[KH19] Joachim I. Krueger and Patrick R. Heck. Putting the P -value in its place. The American Statistician, 73(S1):122–128, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2018.1470033.

Khachatryan:2015:ISC

[Kha15] Davit Khachatryan. Incorporating statistical consulting case studies in introductory time series courses. The American Statistician, 69(4):387–396, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kittaneh:2017:RAE

[Kit17] Omar A. Kittaneh. Response to “Average entropy does not measure uncertainty”. The American Statistician, 71(1):91, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See [Hod17].

Kittaneh:2016:AEN

[KKAB16] Omar A. Kittaneh, Mohammad A. U. Khan, Muhammed Akbar, and Husam A. Bayoud. Average entropy: A new uncertainty measure with application to image segmentation. The American Statistician, 70(1):18–24, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kettenring:2015:COS

[KKM15] Jon R. Kettenring, Kenneth J. Koehler, and John D. McKenzie Jr. Challenges and opportunities for statistics in the next 25 years. The American Statistician, 69(2):86–90, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kessy:2018:OWD

[KLS18] Agnan Kessy, Alex Lewin, and Korbinian Strimmer. Optimal whitening and decorrelation. The American Statistician, 72(4): 309–314, 2018. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://amstat.tandfonline.com/doi/
full/10.1080/00031305.2016.1277159.

Keele:2019:RIO

[KM19] Luke Keele and Luke Miratrix. Randomization inference for outcomes with clumping at zero. The American Statistician, 73(2): 141–150, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://amstat.tandfonline.com/doi/
full/10.1080/00031305.2017.1385535.

Kmetz:2019:CCR

[Kme19] John L. Kmetz. Correcting corrupt research: Recommendations for the profession to stop misuse of p-values. The American Statistician, 73(S1):36–45, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://amstat.

tandfonline.com/doi/full/10.1080/00031305.2018.1518271.

Kulinskaya:2010:VSD

[KMS10] Elena Kulinskaya, Stephan Morgenthaler, and Robert G. Staudte. Variance stabilizing the difference of two binomial proportions. The American Statistician, 64(4):350–356, November 2010. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Keogh:2012:SAF

[KO12] Sarah Keogh and Donal O’Neill. A statistical analysis of the fairness of alternative handicapping systems in ten-pin bowling. The American Statistician, 66(4):209–213, 2012. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kong:2011:ESC

[Kon11] Yong Kong. The expectation of sample central moments. The American Statistician, 65(3):198–199, August 2011. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kotz:2010:TIU

[Kot10] Brian C. Kotz. Thoughts on the importance of the undergraduate statistics experience to the discipline’s (and society’s) future. The American Statistician, 64(1):15–18, February 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kocak:2012:SBE

[KOT12] Mehmet Kocak and Arzu Onar-Thomas. A simulation-based evaluation of the asymptotic power formulas for Cox models in small sample cases. The American Statistician, 66(3):173–179, 2012. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kourouklis:2012:NEV

[Kou12] Stavros Kourouklis. A new estimator of the variance based on minimizing mean squared error. The American Statistician, 66 (4):234–236, 2012. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Keeling:2011:SAS

[KP11] Kellie B. Keeling and Robert J. Pavur. Statistical accuracy of spreadsheet software. The American Statistician, 65(4):265–273, November 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Karrison:2012:EPF

[KRSR12] Theodore G. Karrison, Mark J. Ratain, Walter M. Stadler, and Gary L. Rosner. Estimation of progression-free survival for all treated patients in the randomized discontinuation trial design. The American Statistician, 66(3):155–162, 2012. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kuiper:2015:UOG

[KS15] Shonda Kuiper and Rodney X. Sturdivant. Using online gamebased simulations to strengthen students’ understanding of practical statistical issues in real-world data analysis. The American Statistician, 69(4):354–361, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kennedy-Shaffer:2019:BBU

[KS19] Lee Kennedy-Shaffer. Before p < 0.05 to beyond p < 0.05: Using history to contextualize p-values and significance testing. The American Statistician, 73(S1):82–90, 2019. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2018.1537891.

Katzfuss:2016:UEK

[KSW16] Matthias Katzfuss, Jonathan R. Stroud, and Christopher K. Wikle. Understanding the ensemble Kalman filter. The American Statistician, 70(4):350–357, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kroonenberg:2018:TCR

[KV18] P. M. Kroonenberg and Albert Verbeek. The tale of Cochran’s rule: My contingency table has so many expected values smaller than 5, what am I to do? The American Statistician, 72(2):175– 183, 2018. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://amstat.tandfonline.com/doi/full/
10.1080/00031305.2017.1286260.

Kuffner:2019:WVC

[KW19] Todd A. Kuffner and Stephen G. Walker. Why are p-values controversial? The American Statistician, 73(1):1–3, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2016.1277161.

Kwasny:2017:MAR

[Kwa17] Mary Kwasny. Mentoring in the ASA: A rejoinder. The American Statistician, 71(1):5, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kleiber:2016:VCD

[KZ16] Christian Kleiber and Achim Zeileis. Visualizing count data regressions using rootograms. The American Statistician, 70(3): 296–??, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lesser:2016:AHC

[LAK16] Martin L. Lesser, Meredith B. Akerman, and Nina Kohn. Analogies for helping clinicians and investigators better understand the principles and practice of biostatistics. The American Statistician, 70(2):166–170, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

LaMotte:2014:GSC

[LaM14] Lynn Roy LaMotte. The Gram–Schmidt construction as a basis for linear models. The American Statistician, 68(1):52–55, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lang:2017:MME

[Lan17] Joseph B. Lang. Mean–minimum exact confidence intervals. The American Statistician, 71(4):354–368, 2017. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2016.1256838.

Lavine:2019:FBO

[Lav19] Michael Lavine. Frequentist, Bayes, or other? The American Statistician, 73(S1):312–318, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://amstat.

tandfonline.com/doi/full/10.1080/00031305.2018.1459317.

Liao:2019:SJI

[LB19] J. G. Liao and Arthur Berg. Sharpening Jensen’s inequality. The American Statistician, 73(3):278–281, 2019. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2017.1419145.

Lee:2017:SRR

[LBR17] Jarod Y. L. Lee, James J. Brown, and Louise M. Ryan. Sufficiency revisited: Rethinking statistical algorithms in the big data era. The American Statistician, 71(3):202–208, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2016.1255659.

Lauzon:2009:EMC

[LC09] Carolyn Lauzon and Brian Caffo. Easy multiplicity control in equivalence testing using two one-sided tests. The American Statistician, 63(2):147–154, May 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See correction [CLR13].

Lee:2015:TGC

[LC15] Hyunju Lee and Ji Hwan Cha. On two general classes of discrete bivariate distributions. The American Statistician, 69(3):221– 230, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00031305.2015.1044564.

Lozada-Can:2010:TEA

[LCD10] C. Lozada-Can and A. C. Davison. Three examples of accurate likelihood inference. The American Statistician, 64(2):131–139, May 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lyon:2016:AUG

[LCG16] Merritt Lyon, Li C. Cheung, and Joseph L. Gastwirth. The advantages of using group means in estimating the Lorenz curve and Gini index from grouped data. The American Statistician, 70(1):25–32, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lynn:2010:CSA

[LDM10] Henry S. Lynn, Zhanjian Dong, and Zhe Mu. Comparison of software algorithms for calculating REML Wald type confidence limits for the between-group variance component in a small sample one-way random effects model example. The American Statistician, 64(1):83–87, February 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

LeBlanc:2019:BAN

[Le 19] Richard Le Blanc. Bayesian analysis on a noncentral Fisher– Student’s hypersphere. The American Statistician, 73(2):126–140, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2017.1377111.

Lesser:2016:LE

[Les16] Lawrence M. Lesser. Letter to the Editor. The American Statistician, 70(4):434, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Loy:2016:VPP

[LFH16] Adam Loy, Lendie Follett, and Heike Hofmann. Variations of Q– Q plots: The power of our eyes! The American Statistician, 70 (2):202–214, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lipsitz:2017:ECR

[LFS^{+}17] Stuart R. Lipsitz, Garrett M. Fitzmaurice, Debajyoti Sinha, Nathanael Hevelone, Edward Giovannucci, Quoc-Dien Trinh, and Jim C. Hu. Efficient computation of reduced regression models. The American Statistician, 71(2):171–176, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2017.1296375.

Liu:2019:SMI

[LG19] Dongmeng Liu and Jinko Graham. Simple measures of individual cluster-membership certainty for hard partitional clustering. The American Statistician, 73(1):70–79, 2019. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2018.1459315.

Lu:2011:ONM

[LGXB11] Bo Lu, Robert Greevy, Xinyi Xu, and Cole Beck. Optimal nonbipartite matching and its statistical applications. The American Statistician, 65(1):21–30, February 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lavine:2012:RSI

[LH12] Michael L. Lavine and James S. Hodges. On rigorous specification of ICAR models. The American Statistician, 66(1):42–49, 2012. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Leman:2015:DNI

[LHH15] Scotland Leman, Leanna House, and Andrew Hoegh. Developing a new interdisciplinary computational analytics undergraduate program: A qualitative-quantitative-qualitative approach. The American Statistician, 69(4):397–408, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Litterer:2019:BRS

[Lit19] Christian Litterer. Book review: Stochastic Processes: From Applications to Theory. The American Statistician, 73(4):418–419, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2019.1676116.

Levin:2013:NIB

[LL13] Bruce Levin and Cheng-Shiun Leu. Note on an identity between two unbiased variance estimators for the grand mean in a simple random effects model. The American Statistician, 67(1):42– 43, 2013. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Liao:2016:CBF

[LLB16] J. G. Liao, Duanping Liao, and Arthur Berg. Calibrated Bayes factors in assessing genetic association models. The American Statistician, 70(3):250–??, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lund:2016:NAA

[LLS16] Robert Lund, Gang Liu, and Qin Shao. A new approach to ANOVA methods for autocorrelated data. The American Statistician, 70(1):55–62, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Liu:2014:CFR

[LM14] Keli Liu and Xiao-Li Meng. Comment: A fruitful resolution to Simpson’s Paradox via multiresolution inference. The American Statistician, 68(1):17–29, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See [Arm14b, Arm14a].

Lee:2015:ENE

[LMZ15] Lisa M. Lee, Frances A. McCarty, and Tenny R. Zhang. Ethical numbers: Ethics training in U.S. graduate statistics programs, 2013–2014. The American Statistician, 69(1):11–16, 2015. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Liu:2016:NMD

[LNM16] Jia Liu, Daniel J. Nordman, and William Q. Meeker. The number of MCMC draws needed to compute Bayesian credible bounds. The American Statistician, 70(3):275–??, 2016. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lo:2019:DIT

[Lo19] Ambrose Lo. Demystifying the integrated tail probability expectation formula. The American Statistician, 73(4):367–374, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2018.1497541.

Locascio:2019:IRB

[Loc19] Joseph J. Locascio. The impact of results blind science publishing on statistical consultation and collaboration. The American Statistician, 73(S1):346–351, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://amstat.

tandfonline.com/doi/full/10.1080/00031305.2018.1505658.

Lomax:2013:SAI

[Lom13] Richard G. Lomax. Statistical accuracy of iPad applications: An initial examination. The American Statistician, 67(2):105–108, 2013. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Longford:2010:BDM

[Lon10] Nicholas T. Longford. Bayesian decision making about small binomial rates with uncertainty about the prior. The American Statistician, 64(2):164–169, May 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lee:2014:DCV

[LP14] Woojoo Lee and Yudi Pawitan. Direct calculation of the variance of maximum penalized likelihood estimates via EM algorithm. The American Statistician, 68(2):93–97, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Liu:2016:SHP

[LP16] Piaomu Liu and Edsel A. Peña. Sojourning with the homogeneous Poisson process. The American Statistician, 70(4):413– 423, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lum:2013:AMS

[LPB13a] Kristian Lum, Megan Emily Price, and David Banks. Applications of multiple systems estimation in human rights research. The American Statistician, 67(4):191–200, 2013. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lum:2013:R

[LPB13b] Kristian Lum, Megan Emily Price, and David Banks. Rejoinder. The American Statistician, 67(4):205–206, 2013. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lee:2016:EVP

[LR16] Joseph J. Lee and Donald B. Rubin. Evaluating the validity of post-hoc subgroup inferences: A case study. The American Statistician, 70(1):39–46, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lazar:2011:CCU

[LRF11] Nicole A. Lazar, Jaxk Reeves, and Christine Franklin. A capstone course for undergraduate statistics majors. The American Statistician, 65(3):183–189, August 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Louzada:2019:NBC

[LRR19] Francisco Louzada, Pedro L. Ramos, and Eduardo Ramos. A note on bias of closed-form estimators for the gamma distribution derived from likelihood equations. The American Statistician, 73 (2):195–199, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/00031305.2018.1513376.

Legler:2010:MIU

[LRZG^{+}10] Julie Legler, Paul Roback, Kathryn Ziegler-Graham, James Scott, Sharon Lane-Getaz, and Matthew Richey. A model for an interdisciplinary undergraduate research program. The American Statistician, 64(1):59–69, February 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Langford:2001:PBP

[LSO01] Eric Langford, Neil Schwertman, and Margaret Owens. Is the property of being positively correlated transitive? The American Statistician, 55(4):322–325, November 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2685695. See comment [Fri15].

Li:2010:CCI

[LTN10] Zhiguo Li, Jeremy M. G. Taylor, and Bin Nan. Construction of confidence intervals and regions for ordered binomial probabilities. The American Statistician, 64(4):291–298, November 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lu:2019:BRS

[Lu19] Minggen Lu. Book review: Survival Analysis with Interval-Censored Data: A Practical Approach with Examples in R, SAS, and BUGS. The American Statistician, 73(2):211–212, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2019.1603477.

Love:2017:DCP

[LVH^{+}17] Kim Love, Eric A. Vance, Frank E. Harrell, Jr., Dallas E. Johnson, Michael H. Kutner, Ronald D. Snee, and Doug Zahn. Developing a career in the practice of statistics: The mentor’s perspective. The American Statistician, 71(1):38–46, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Liu:2015:UAU

[LWM15] Shiyao Liu, Huaiqing Wu, and William Q. Meeker. Understanding and addressing the unbounded “Likelihood” problem. The American Statistician, 69(3):191–200, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/00031305.2014.1003968.

Lynn:2016:TNG

[Lyn16] Henry S. Lynn. Training the next generation of statisticians: From head to heart. The American Statistician, 70(2):149–151, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Malinovsky:2015:NMS

[MA15] Yaakov Malinovsky and Paul S. Albert. A note on the minimax solution for the two-stage group testing problem. The American Statistician, 69(1):45–52, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Malinovsky:2019:RNG

[MA19] Yaakov Malinovsky and Paul S. Albert. Revisiting nested group testing procedures: New results, comparisons, and robustness. The American Statistician, 73(2):117–125, 2019.

CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2017.1366367.

MacDonald:2017:MCD

[Mac17] Iain L. MacDonald. Models for count data. The American Statistician, 71(2):187–190, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://amstat.

tandfonline.com/doi/full/10.1080/00031305.2017.1291449.

See [ZSMS17] and reply [SMSZ17].

Mandel:2013:SBC

[Man13] Micha Mandel. Simulation-based confidence intervals for functions with complicated derivatives. The American Statistician, 67(2):76–81, 2013. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Manski:2019:TCT

[Man19] Charles F. Manski. Treatment choice with trial data: Statistical decision theory should supplant hypothesis testing. The American Statistician, 73(S1):296–304, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://amstat.tandfonline.com/doi/full/10.1080/
00031305.2018.1513377.

Martin:2017:SIC

[Mar17] Ryan Martin. A statistical inference course based on pvalues. The American Statistician, 71(2):128–136, 2017. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2016.1208629.

Matthews:2019:MTP

[Mat19] Robert A. J. Matthews. Moving towards the post p < 0.05 era via the analysis of credibility. The American Statistician, 73(S1): 202–212, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://amstat.tandfonline.com/doi/
full/10.1080/00031305.2018.1543136.

Mayo:2013:DBM

[May13] Deborah G. Mayo. Discussion: Bayesian methods: Applied? Yes. Philosophical defense? In flux. The American Statistician, 67(1): 11–15, 2013. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Marwick:2018:PDA

[MBM18] Ben Marwick, Carl Boettiger, and Lincoln Mullen. Packaging data analytical work reproducibly using R (and friends). The American Statistician, 72(1):80–88, 2018. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://amstat.tandfonline.com/doi/full/10.1080/
00031305.2017.1375986.

McCulloch:2014:C

[McC14] Robert McCulloch. Comment. The American Statistician, 68(2): 87–88, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See [Har14a].

Martinez:2011:EPC

[MCM^{+}11] Josue G. Martinez, Raymond J. Carroll, Samuel Müller, Joshua N. Sampson, and Nilanjan Chatterjee. Empirical performance of cross-validation with oracle methods in a genomics context. The American Statistician, 65(4):223–228, November 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

McNamara:2019:KAM

[McN19] Amelia McNamara. Key attributes of a modern statistical computing tool. The American Statistician, 73(4):375–384, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2018.1482784.

Morris:2011:NCP

[ME11] Nathan Morris and Robert Elston. A note on comparing the power of test statistics at low significance levels. The American Statistician, 65(3):164–166, August 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Meng:2010:RBT

[Men10] Xiao-Li Meng. Rejoinder: Better training, deeper thinking, and more policing. The American Statistician, 64(1):26–29, February 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Miao:2014:NST

[MG14] Weiwen Miao and Joseph L. Gastwirth. New statistical tests for detecting disparate impact arising from two-stage selection processes. The American Statistician, 68(3):146–157, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Malone:2010:RTI

[MGCR10] Christopher J. Malone, John Gabrosek, Phyllis Curtiss, and Matt Race. Resequencing topics in an introductory applied statistics course. The American Statistician, 64(1):52–58, February 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

McShane:2019:ASS

[MGG^{+}19] Blakeley B. McShane, David Gal, Andrew Gelman, Christian Robert, and Jennifer L. Tackett. Abandon statistical significance. The American Statistician, 73(S1):235–245, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2018.1527253.

Marks:2014:ABM

[MGRL14] Christopher E. Marks, Andrew G. Glen, Matthew W. Robinson, and Lawrence M. Leemis. Applying bootstrap methods to system reliability. The American Statistician, 68(3):174–182, 2014. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

McNamara:2018:WCD

[MH18] Amelia McNamara and Nicholas J. Horton. Wrangling categorical data in R. The American Statistician, 72(1):97–104, 2018. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2017.1356375.

Maurer:2019:CAV

[MHWB19] Karsten Maurer, Lynette Hudiburgh, Lisa Werwinski, and John Bailer. Content audit for p-value principles in introductory statistics. The American Statistician, 73(S1):385–391, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2018.1537890.

Michalopoulou:2014:UCP

[Mic14] Catherine Michalopoulou. A unique collaboration: Prominent statisticians’ survey work in Greece in 1946. The American Statistician, 68(3):196–203, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Missiakoulis:2015:LE

[Mis15] Spyros Missiakoulis. Letter to the Editor. The American Statistician, 69(1):62, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Missiakoulis:2019:PSL

[Mis19] Spyros Missiakoulis. Phlegon’s stem-and-leaf display. The American Statistician, 73(1):89–93, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://amstat.

tandfonline.com/doi/full/10.1080/00031305.2017.1328376.

Moore:2015:PAU

[MK15] Allison Amanda Moore and Jennifer J. Kaplan. Program assessment for an undergraduate statistics major. The American Statistician, 69(4):417–424, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

McDaid:2019:FDT

[MKR19] Aaron McDaid, Zoltán Kutalik, and Valentin Rousson. A fivedecision testing procedure to infer the value of a unidimensional parameter. The American Statistician, 73(4):321–326, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2018.1437075.

Muggeo:2014:T

[ML14] Vito M. R. Muggeo and Gianfranco Lovison. The “Three plus one” likelihood-based test statistics: Unified geometrical and graphical interpretations. The American Statistician, 68(4):302– 306, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

MacDonald:2016:EMD

[ML16] Iain L. MacDonald and Brendon M. Lapham. Even more direct calculation of the variance of a maximum penalized-likelihood estimator. The American Statistician, 70(1):114–118, 2016. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Mehrotra:2010:RBA

[MLL10] Devan V. Mehrotra, Xiaomin Lu, and Xiaoming Li. Rank-based analyses of stratified experiments: Alternatives to the van Elteren test. The American Statistician, 64(2):121–130, May 2010. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Mason:2015:BHA

[MM15] Robert L. Mason and John D. McKenzie Jr. A brief history of the American Statistical Association, 1990–2014. The American Statistician, 69(2):68–78, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Morganstein:2017:MAC

[Mor17] David Morganstein. Mentoring in the ASA: A commentary. The American Statistician, 71(1):3–4, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Maity:2019:BRL

[MPD19] Arnab Kumar Maity, Vivek Pradhan, and Ujjwal Das. Bias reduction in logistic regression with missing responses when the missing data mechanism is nonignorable. The American Statistician, 73(4):340–349, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00031305.2017.1407359.

Mutz:2019:PBT

[MPP19] Diana C. Mutz, Robin Pemantle, and Philip Pham. The perils of balance testing in experimental design: Messy analyses of clean data. The American Statistician, 73(1):32–42, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2017.1322143.

Michlin:2014:IDR

[MS14] Yefim Haim Michlin and Ofer Shaham. Ignatova, I., Deutsch, R. C., and Edwards, D. (2012), “Closed Sequential and Multistage Inference on Binary Responses With or Without Replacement,” The American Statistician, 66, 163–172: Comment by Michlin and Shaham and Reply. The American Statistician, 68 (2):128, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See [IDE12].

McDonald:2011:DCJ

[MT11] James B. McDonald and Patrick Turley. Distributional characteristics: Just a few more moments. The American Statistician, 65(2):96–103, May 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Manski:2019:TSN

[MT19] Charles F. Manski and Aleksey Tetenov. Trial size for nearoptimal choice between surveillance and aggressive treatment: Reconsidering MSLT-II. The American Statistician, 73(S1):305– 311, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://amstat.tandfonline.com/doi/full/
10.1080/00031305.2018.1543617.

McShane:2019:LSR

[MTBG19] Blakeley B. McShane, Jennifer L. Tackett, Ulf Böckenholt, and Andrew Gelman. Large-scale replication projects in contemporary psychological research. The American Statistician, 73(S1): 99–105, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://amstat.tandfonline.com/doi/
full/10.1080/00031305.2018.1505655.

Mukhopadhyay:2010:WFM

[Muk10] Nitis Mukhopadhyay. When finiteness matters: Counterexamples to notions of covariance, correlation, and independence. The American Statistician, 64(3):231–233, August 2010. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Mukhopadhyay:2014:CRW

[Muk14a] Nitis Mukhopadhyay. Comment and reply: Warr, R. L. and Erich, R. A. (2013), “Should the Interquartile Range Divided by the Standard Deviation be Used to Assess Normality?,” The American Statistician 67, 242–244. The American Statistician, 68(4): 316–317, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See [WE13].

Mukhopadhyay:2014:GDS

[Muk14b] Nitis Mukhopadhyay. Griffith, D. A. (2013), “Better Articulating Normal Curve Theory for Introductory Mathematical Statistics Students: Power Transformations and Their Back– Transformations,” The American Statistician, 67, 157-169: Comment by Mukhopadhyay and Reply. The American Statistician, 68(1):67, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See [Gri13, Gri14].

Nadarajah:2015:CGH

[Nad15] Saralees Nadarajah. On the computation of Gauss hypergeometric functions. The American Statistician, 69(2):146–148, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Nadarajah:2016:LE

[Nad16] Saralees Nadarajah. Letter to the Editor. The American Statistician, 70(2):224, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Narayanan:2012:RES

[Nar12] A. Narayanan. A review of eight software packages for structural equation modeling. The American Statistician, 66(2):129– 138, 2012. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://amstat.tandfonline.com/doi/full/
10.1080/00031305.2012.708641.

Nievergelt:2011:IDN

[Nie11] Yves Nievergelt. Increasing data with a negative slope. The American Statistician, 65(4):262, November 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Nolan:2010:CSC

[NL10] Deborah Nolan and Duncan Temple Lang. Computing in the statistics curricula. The American Statistician, 64(2):97–107, May 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Nolan:2015:ESR

[NL15] Deborah Nolan and Duncan Temple Lang. Explorations in statistics research: An approach to expose undergraduates to authentic data analysis. The American Statistician, 69(4):292–299, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Nadarajah:2018:EFC

[NL18] Saralees Nadarajah and Rui Li. An expression for fast computation of sample central moments. The American Statistician, 72 (2):169–171, 2018. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/00031305.2017.1286259.

Noguchi:2016:AEM

[NMR16] Kimihiro Noguchi and Fernando Marmolejo-Ramos. Assessing equality of means using the overlap of range-preserving confidence intervals. The American Statistician, 70(4):325–334, 2016. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Nolan:2016:TLD

[NP16] Deborah Nolan and Jamis Perrett. Teaching and learning data visualization: Ideas and assignments. The American Statistician, 70(3):260–??, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Niu:2019:NBI

[NR19] Xiaoyue Niu and James L. Rosenberger. Near-balanced incomplete block designs, with an application to poster competitions. The American Statistician, 73(2):159–164, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2017.1385534.

Nosedal-Sanchez:2012:RKH

[NSSLC12] Alvaro Nosedal-Sanchez, Curtis B. Storlie, Thomas C. M. Lee, and Ronald Christensen. Reproducing kernel Hilbert spaces for penalized regression: a tutorial. The American Statistician, 66(1): 50–60, 2012. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

OHagan:2019:EKE

[O’H19] Anthony O’Hagan. Expert knowledge elicitation: Subjective but scientific. The American Statistician, 73(S1):69–81, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2018.1518265.

Oldford:2016:SCQ

[Old16] R. Wayne Oldford. Self-calibrating quantile-quantile plots. The American Statistician, 70(1):74–90, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

ONeill:2014:SUM

[O’N14] B. O’Neill. Some useful moment results in sampling problems. The American Statistician, 68(4):282–296, 2014. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

ONeill:2016:C

[O’N16] Ben O’Neill. Corrigendum. The American Statistician, 70(3):323– ??, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Owen:2014:SLR

[OR14] A. B. Owen and P. A. Roediger. The sign of the logistic regression coefficient. The American Statistician, 68(4):297–301, 2014. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Oster:2011:SENa

[Ost11a] Robert A. Oster. Section editor’s notes. The American Statistician, 65(1):43, February 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Oster:2011:SENb

[Ost11b] Robert A. Oster. Section editor’s notes. The American Statistician, 65(2):113–114, May 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Oster:2011:SENc

[Ost11c] Robert A. Oster. Section editor’s notes. The American Statistician, 65(3):200, August 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Oster:2011:SEN

[Ost11d] Robert A. Oster. Section editor’s notes. The American Statistician, 65(4):263–264, November 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Oster:2012:SENa

[Ost12a] Robert A. Oster. Section Editor’s notes. The American Statistician, 66(2):128, 2012. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/00031305.2012.707548.

Oster:2012:SENb

[Ost12b] Robert A. Oster. Section Editor’s notes. The American Statistician, 66(4):238, 2012. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Oster:2013:SENa

[Ost13a] Robert A. Oster. Section Editor’s notes. The American Statistician, 67(2):104, 2013. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Oster:2013:SENb

[Ost13b] Robert A. Oster. Section Editor’s notes. The American Statistician, 67(3):170, 2013. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Oster:2014:SEN

[Ost14] Robert A. Oster. Section Editor’s notes. The American Statistician, 68(3):204, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Ormerod:2010:EVA

[OW10] J. T. Ormerod and M. P. Wand. Explaining variational approximations. The American Statistician, 64(2):140–153, May 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Park:2019:PTS

[Par19] Ryoungsun Park. Practical teaching strategies for hypothesis testing. The American Statistician, 73(3):282–287, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2018.1424034.

Pathak:2018:SPP

[Pat18] Ashok Kumar Pathak. A simple probabilistic proof for the alternating convolution of the central binomial coefficients. The American Statistician, 72(3):287–288, 2018. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2017.1358216.

Patriota:2019:MVT

[Pat19] Alexandre Galvão Patriota. On the mean value theorem for estimating functions. The American Statistician, 73(4):408–410, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2018.1558110.

Pentoney:2016:CIW

[PB16] Christopher S. Pentoney and Dale E. Berger. Confidence intervals and the within-the-bar bias. The American Statistician, 70(2): 215–220, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Perkins:2016:BPB

[PBD^{+}16] Susan M. Perkins, Peter Bacchetti, Cynthia S. Davey, Christopher J. Lindsell, Madhu Mazumdar, Robert A. Oster, Peter N. Peduzzi, David M. Rocke, Kyle D. Rudser, Mimi Kim, and the Biostatistics and Epidemiology and Research Design (BERD) Key Function Committee of the Clinical and Translational Science (CTSA) Consortium. Best practices for biostatistical consultation and collaboration in academic health centers. The American Statistician, 70(2):187–194, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Pearl:2014:CUS

[Pea14] Judea Pearl. Comment: Understanding Simpson’s Paradox. The American Statistician, 68(1):8–13, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Puig:2016:UAA

[PFG16] Xavier Puig, Martí Font, and Josep Ginebra. A unified approach to authorship attribution and verification. The American Statistician, 70(3):232–??, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Poulson:2012:THI

[PGA12] Robert S. Poulson, Gary L. Gadbury, and David B. Allison. Treatment heterogeneity and individual qualitative interaction. The American Statistician, 66(1):16–24, 2012. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Pfeifer:2014:PPC

[PGCL14] Phillip E. Pfeifer, Yael Grushka-Cockayne, and Kenneth C. Lichtendahl Jr. The promise of prediction contests. The American Statistician, 68(4):264–270, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Pawitan:2017:WGP

[PL17] Yudi Pawitan and Youngjo Lee. Wallet game: Probability, likelihood, and extended likelihood. The American Statistician, 71 (2):120–122, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/00031305.2016.1202140.

Pogrow:2019:HES

[Pog19] Stanley Pogrow. How effect size (practical significance) misleads clinical practice: The case for switching to practical benefit to assess applied research findings. The American Statistician, 73 (S1):223–234, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/00031305.2018.1549101.

Portnoy:2015:MPB

[Por15] Stephen Portnoy. Maximizing probability bounds under momentmatching restrictions. The American Statistician, 69(1):41–44, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Porter:2016:SAC

[Por16] Michael D. Porter. A statistical approach to crime linkage. The American Statistician, 70(2):152–165, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Portnoy:2019:IOO

[Por19] Stephen Portnoy. Invariance, optimality, and a 1-observation confidence interval for a normal mean. The American Statistician, 73(1):10–15, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/00031305.2017.1360796.

Pavlides:2010:EFP

[PP10] Marios G. Pavlides and Michael D. Perlman. On estimating the face probabilities of shaved dice with partial data. The American Statistician, 64(1):37–45, February 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Proschan:2010:BQQ

[PR10] Michael A. Proschan and Jeffrey S. Rosenthal. Beyond the quintessential quincunx. The American Statistician, 64(1):78– 82, February 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Plante:2011:SN

[PR11] Jean-François Plante and Nancy Reid. Statistics in the news. The American Statistician, 65(2):80–88, May 2011. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Priebe:2011:FCP

[Pri11] Carey E. Priebe. Fisher’s conditionality principle in statistical pattern recognition. The American Statistician, 65(3):167–169, August 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Phelps:2017:CLT

[PS17] Amy L. Phelps and Kathryn A. Szabat. The current landscape of teaching analytics to business students at institutions of higher education: Who is teaching what? The American Statistician, 71(2):155–161, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/00031305.2016.1277160.

Prendergast:2018:SEI

[PS18] Luke A. Prendergast and Robert G. Staudte. A simple and effective inequality measure. The American Statistician, 72(4):328– 343, 2018. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://amstat.tandfonline.com/doi/full/
10.1080/00031305.2017.1366366.

Picard:2013:REE

[PW13] Rick Picard and Brian Williams. Rare event estimation for computer models. The American Statistician, 67(1):22–32, 2013. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Quarfoot:2016:HRM

[QL16] David Quarfoot and Richard A. Levine. How robust are multirater interrater reliability indices to changes in frequency distribution? The American Statistician, 70(4):373–384, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Ramezan:2017:RBT

[Ram17] Reza Ramezan. Reviews of books and teaching materials. The American Statistician, 71(1):92–96, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Rosenblatt:2018:MAW

[RB18] Jonathan D. Rosenblatt and Yoav Benjamini. On mixture alternatives and Wilcoxon’s signed-rank test. The American Statistician, 72(4):344–347, 2018. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/00031305.2017.1360795.

Ruberg:2019:IDM

[RHGS^{+}19] Stephen J. Ruberg, Frank E. Harrell Jr., Margaret Gamalo-Siebers, Lisa LaVange, J. Jack Lee, Karen Price, and Carl Peck. Inference and decision making for 21st-Century drug development and approval. The American Statistician, 73(S1):319–327, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2019.1566091.

Rice:2010:DTF

[Ric10] Kenneth Rice. A decision-theoretic formulation of Fisher’s approach to testing. The American Statistician, 64(4):345–349, November 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Richardson:2019:CRN

[Ric19] Alice Richardson. A comparative review of nonparametric statistics textbooks. The American Statistician, 73(4):360–366, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2018.1437076.

Reich:2016:CSE

[RLS^{+}16] Nicholas G. Reich, Justin Lessler, Krzysztof Sakrejda, Stephen A. Lauer, Sopon Iamsirithaworn, and Derek A. T. Cummings. Case study in evaluating time series prediction models using the relative mean absolute error. The American Statistician, 70(3):285– ??, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Rose:2019:LVS

[RM19a] Sherri Rose and Thomas G. McGuire. Limitations of P -values and R-squared for stepwise regression building: A fairness demonstration in health policy risk adjustment. The American Statistician, 73(S1):152–156, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/00031305.2018.1518269.

Rouder:2019:TBT

[RM19b] Jeffrey N. Rouder and Richard D. Morey. Teaching Bayes’ Theorem: Strength of evidence as predictive accuracy. The American Statistician, 73(2):186–190, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://amstat.

tandfonline.com/doi/full/10.1080/00031305.2017.1341334.

Robinson:2019:WPM

[Rob19] Geoffrey K. Robinson. What properties might statistical inferences reasonably be expected to have? — Crisis and resolution in statistical inference. The American Statistician, 73(3):243– 252, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://amstat.tandfonline.com/doi/full/
10.1080/00031305.2017.1415971.

Rocks:2016:IEN

[Roc16] Brendan Rocks. Interval estimation for the “Net promoter score”. The American Statistician, 70(4):365–372, 2016. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Rodriguez:2015:WWC

[Rod15] Robert N. Rodriguez. Who will celebrate our 200th anniversary? Growing the next generation of ASA members. The American Statistician, 69(2):91–95, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Rougier:2019:VBF

[Rou19] Jonathan Rougier. p-values, Bayes factors, and sufficiency. The American Statistician, 73(S1):148–151, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://amstat.tandfonline.com/doi/full/10.1080/
00031305.2018.1502684.

Rosenbaum:2013:UEM

[RS13] Paul R. Rosenbaum and Jeffrey H. Silber. Using the exterior match to compare two entwined matched control groups. The American Statistician, 67(2):67–75, 2013. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Rossman:2015:APS

[RST15] Allan J. Rossman, Roy St. Laurent, and Josh Tabor. Advanced placement statistics: Expanding the scope of statistics education. The American Statistician, 69(2):121–126, 2015. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Sabbag:2019:BRH

[Sab19] Anelise G. Sabbag. Book review: Handbook of Educational Measurement and Psychometrics Using R, by Christopher D. Desjardins and Okan Bulut. Boca Raton, FL: Chapman Hall/ CRC Press, 2018, xxiii + 302 pp., $79.95 (H), ISBN: 978-1-49-877013-2. The American Statistician, 73(4):415–416, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2019.1676110.

Sambucini:2013:NSP

[Sam13] Valeria Sambucini. On the nature of the stationary point of a quadratic response surface: a Bayesian simulation-based approach. The American Statistician, 67(1):33–41, 2013. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Samsa:2015:IRB

[Sam15] Gregory P. Samsa. Has it really been demonstrated that most genomic research findings are false? The American Statistician, 69 (1):1–4, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Seier:2011:PVL

[SB11] Edith Seier and Douglas G. Bonett. A polyplot for visualizing location, spread, skewness, and kurtosis. The American Statistician, 65(4):258–261, November 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Schofield:2014:RBT

[SB14] Matthew Schofield and Patrick Breheny. Reviews of books and teaching materials. The American Statistician, 68(1):63–66, 2014.

CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Smucker:2015:BNP

[SB15] Byran J. Smucker and A. John Bailer. Beyond normal: Preparing undergraduates for the work force in a statistical consulting capstone. The American Statistician, 69(4):300–306, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Schell:2010:IKS

[Sch10a] Michael J. Schell. Identifying key statistical papers from 1985 to 2002 using citation data for applied biostatisticians. The American Statistician, 64(4):310–317, November 2010. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Schwarz:2010:CCD

[Sch10b] Wolf Schwarz. Comparing continuous and discrete birthday coincidences: “Same-Day” versus “Within 24 hours”. The American Statistician, 64(1):34–36, February 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Schofield:2012:CBL

[Sch12] Lynne Steuerle Schofield. Community-based learning versus traditional courses in statistics: Who takes them and why. The American Statistician, 66(2):118–123, 2012. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://amstat.

tandfonline.com/doi/full/10.1080/00031305.2012.701541.

Scheuren:2013:C

[Sch13] Fritz Scheuren. Comment. The American Statistician, 67(4):203– 205, 2013. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Spencer:2017:PET

[SCLP17] Richard G. Spencer, Benjamin D. Cortese, Vanessa A. Lukas, and Nancy Pleshko. Point estimates of test sensitivity and specificity from sample means and variances. The American Statistician, 71(1):81–87, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Schilling:2014:CPA

[SD14] Mark F. Schilling and Jimmy A. Doi. A coverage probability approach to finding an optimal binomial confidence procedure. The American Statistician, 68(3):133–145, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Schilling:2015:R

[SD15] Mark F. Schilling and Jimmy A. Doi. Reply. The American Statistician, 69(2):155–156, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Santi:2019:GTI

[SDE19] Flavio Santi, Maria Michela Dickson, and Giuseppe Espa. A graphical tool for interpreting regression coefficients of trinomial logit models. The American Statistician, 73(2):200–207, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2018.1442368.

South:2019:SPN

[SEC^{+}19] Charles South, Ryan Elmore, Andrew Clarage, Rob Sickorez, and Jing Cao. A starting point for navigating the world of daily fantasy basketball. The American Statistician, 73(2):179–185, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2017.1401559.

Sanchez-Espigares:2018:VTI

[SEGMA18] José A. Sánchez-Espigares, Pere Grima, and Lluís Marco-Almagro. Visualizing Type II error in normality tests. The American Statistician, 72(2):158–162, 2018. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://amstat.tandfonline.com/doi/full/10.1080/
00031305.2016.1278035.

Sen:2012:IBS

[Sen12] Ananda Sen. On the interrelation between the sample mean and the sample variance. The American Statistician, 66(2):112–117, 2012. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Schenfisch:2019:BRS

[SF19] Anna Schenfisch and Brittany Fasy. Book review: Statistical Analysis of Contingency Tables. The American Statistician, 73 (2):208, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://amstat.tandfonline.com/doi/
full/10.1080/00031305.2019.1571848.

Spratt:2017:SSE

[SFSM17] Heidi Spratt, Erin E. Fox, Nawar Shara, and Madhu Mazumdar. Strategies for success: Early-stage collaborating biostatistics faculty in an academic health center. The American Statistician, 71 (3):220–230, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/00031305.2016.1277157.

Swartz:2017:QPM

[SGBS17] Philippa Swartz, Mike Grosskopf, Derek Bingham, and Tim B. Swartz. The quality of pitches in major league baseball. The American Statistician, 71(2):148–154, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://amstat.tandfonline.com/doi/full/10.1080/
00031305.2016.1264313.

Shah:2017:WM

[Sha17] Aarti Shah. What is mentoring? The American Statistician, 71 (1):1–2, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Stanley:2010:CIB

[SJD10] T. D. Stanley, Stephen B. Jarrell, and Hristos Doucouliagos. Could it be better to discard 90% of the data? A statistical paradox. The American Statistician, 64(1):70–77, February 2010. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Steel:2019:BCC

[SLG19] E. Ashley Steel, Martin Liermann, and Peter Guttorp. Beyond calculations: A course in statistical thinking. The American Statistician, 73(S1):392–401, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://amstat.

tandfonline.com/doi/full/10.1080/00031305.2018.1505657.

Sibley:2018:FCE

[SLJ^{+}18] Alexander B. Sibley, Zhiguo Li, Yu Jiang, Yi-Ju Li, Cliburn Chan, Andrew Allen, and Kouros Owzar. Facilitating the calculation of the efficient score using symbolic computing. The American Statistician, 72(2):199–205, 2018. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2017.1392361.

Smith:2011:MMR

[Smi11] Marlene A. Smith. Missteps in multiple regression student projects: Beyond association-not-causation. The American Statistician, 65(3):190–197, August 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Smithline:2012:LE

[Smi12] Lawren Smithline. Letter to the Editor. The American Statistician, 66(3):207, 2012. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Sellers:2017:R

[SMSZ17] Kimberly F. Sellers, Darcy S. Morris, Galit Shmueli, and Li Zhu. Reply [to “Models for count data”]. The American Statistician, 71 (2):190, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://amstat.tandfonline.com/doi/
full/10.1080/00031305.2017.1296738.

Snee:2019:WSS

[Sne19] Ronald D. Snee. We stand on the shoulders of giants — pioneers of statistics in industry. The American Statistician, 73(4):400– 407, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2018.1543140.

Soler:2010:WTI

[Sol10] Frank P. Soler. Who is teaching introductory statistics? The American Statistician, 64(1):19–20, February 2010. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Sarkar:2016:VMM

[SR16] Jyotirmoy Sarkar and Mamunur Rashid. Visualizing mean, median, mean deviation, and standard deviation of a set of numbers. The American Statistician, 70(3):304–??, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Seaman:2012:HDS

[SSS12] John W. Seaman III, John W. Seaman Jr., and James D. Stamey. Hidden dangers of specifying noninformative priors. The American Statistician, 66(2):77–84, 2012. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Segal:2013:RGE

[ST13] Nancy L. Segal and Jorge Torres. A repeated grammatical error does not make it right. The American Statistician, 67(4): 266, 2013. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See comments [Hol14, VM14] and reply [ST14].

Segal:2014:R

[ST14] Nancy L. Segal and Jorge Luis Torres. Reply. The American Statistician, 68(2):127–128, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See [ST13].

Stark:2016:BRP

[Sta16] Philip B. Stark. Book review: Privacy, Big Data, and the Public Good: Frameworks for Engagement, Julia Lane, Victoria Stodden, Stefan Bender, and Helen Nissenbaum (eds.). New York, NY: Cambridge University Press, 2014, xix + 320 pp., $29.99 (P), ISBN: 978-1-107-63768-9. The American Statistician, 70(1): 119, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Stern:2014:C

[Ste14] Hal Stern. Comment. The American Statistician, 68(2):83–84, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See [Har14a].

Stigler:2010:CHR

[Sti10] Stephen M. Stigler. The changing history of robustness. The American Statistician, 64(4):277–281, November 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Stigler:2013:CBI

[Sti13a] Stephen Stigler. Comment: Bayesian inference: The Rodney Dangerfield of statistics? The American Statistician, 67(1):6– 7, 2013. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Stigler:2013:DAB

[Sti13b] Stephen M. Stigler. The digital approximation of the binomial by the Poisson. The American Statistician, 67(1):57–59, 2013. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Stigler:2016:LE

[Sti16] Stephen M. Stigler. Letter to the Editor. The American Statistician, 70(1):127, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Stine:2017:ENQ

[Sti17] Robert A. Stine. Explaining normal quantile-quantile plots through animation: The water-filling analogy. The American Statistician, 71(2):145–147, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://amstat.

tandfonline.com/doi/full/10.1080/00031305.2016.1200488.

Stone:2012:MEH

[Sto12] Daniel F. Stone. Measurement error and the hot hand. The American Statistician, 66(1):61–66, 2012. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Sundberg:2018:NSD

[Sun18] Rolf Sundberg. A note on “Shaved dice” inference. The American Statistician, 72(2):155–157, 2018. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://amstat.

tandfonline.com/doi/full/10.1080/00031305.2016.1277162.

Stander:2018:BSA

[SVCB18] Julian Stander, Luciana Dalla Valle, and Mario Cortina-Borja. A Bayesian survival analysis of a historical dataset: How long do popes live? The American Statistician, 72(4):368–375, 2018. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2017.1328374.

Stellato:2017:MCI

[SVG17] Bartolomeo Stellato, Bart P. G. Van Parys, and Paul J. Goulart. Multivariate Chebyshev inequality with estimated mean and variance. The American Statistician, 71(2):123–127, 2017. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2016.1186559.

Small:2011:STF

[SVR11] Dylan S. Small, Kevin G. Volpp, and Paul R. Rosenbaum. Structured testing of 2 × 2 factorial effects: An analytic plan requiring fewer observations. The American Statistician, 65(1):11–15, February 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Szabo:2019:TTM

[Sza19] Aniko Szabo. Test for trend with a multinomial outcome. The American Statistician, 73(4):313–320, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00031305.2017.

1407823. See comment [Chr20].

Tintle:2015:CAS

[TCC^{+}15] Nathan Tintle, Beth Chance, George Cobb, Soma Roy, Todd Swanson, and Jill VanderStoep. Combating anti-statistical thinking using simulation-based methods throughout the undergraduate curriculum. The American Statistician, 69(4):362–370, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Tsagbey:2017:ADW

[TdCP17] Sitsofe Tsagbey, Miguel de Carvalho, and Garritt L. Page. All data are wrong, but some are useful? Advocating the need for data auditing. The American Statistician, 71(3):231–235, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2017.1311282.

Tadinada:2018:SCV

[TG18] Sashi Kanth Tadinada and Abhinav Gupta. Simulation of constrained variables in engineering risk analyses. The American Statistician, 72(2):130–139, 2018. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://amstat.

tandfonline.com/doi/full/10.1080/00031305.2016.1255660.

Thaden:2018:SEM

[TK18] Hauke Thaden and Thomas Kneib. Structural equation models for dealing with spatial confounding. The American Statistician, 72(3):239–252, 2018. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/00031305.2017.1305290.

Taraldsen:2010:IPI

[TL10] Gunnar Taraldsen and Bo Henry Lindqvist. Improper priors are not improper. The American Statistician, 64(2):154–158, May 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Tillxe:2012:HBI

[TL12] Yves Tillé and Matti Langel. Histogram-based interpolation of the Lorenz curve and Gini index for grouped data. The American Statistician, 66(4):225–231, 2012. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Taylor:2018:FS

[TL18] Sean J. Taylor and Benjamin Letham. Forecasting at scale. The American Statistician, 72(1):37–45, 2018. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://amstat.tandfonline.com/doi/full/10.1080/
00031305.2017.1380080.

Tarpey:2016:SMI

[TO16] Thaddeus Tarpey and R. Todd Ogden. Statistical modeling to inform optimal game strategy: Markov plays H-O-R-S-E. The American Statistician, 70(2):181–186, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Tong:2019:SIE

[Ton19] Christopher Tong. Statistical inference enables bad science; statistical thinking enables good science. The American Statistician, 73(S1):246–261, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/00031305.2018.1518264.

Tarpey:2014:PRE

[TOPC14] Thaddeus Tarpey, R. Todd Ogden, Eva Petkova, and Ronald Christensen. A paradoxical result in estimating regression coefficients. The American Statistician, 68(4):271–276, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See comment [Din14a].

Tarpey:2015:R

[TOPC15] Thaddeus Tarpey, R. Todd Ogden, Eva Petkova, and Ronald Christensen. Reply. The American Statistician, 69(3):254–255, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00031305.2015.1056613.

Toulis:2017:UPQ

[Tou17] Panagiotis (Panos) Toulis. A useful pivotal quantity. The American Statistician, 71(3):272–274, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://amstat.

tandfonline.com/doi/full/10.1080/00031305.2016.1237894.

Tarpey:2019:LE

[TP19] Thaddeus Tarpey and Eva Petkova. Letter to the Editor. The American Statistician, 73(3):312, 2019. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2018.1537894.

Trafimow:2019:FNC

[Tra19] David Trafimow. Five nonobvious changes in editorial practice for Editors and reviewers to consider when evaluating submissions in a post p < 0.05 universe. The American Statistician, 73(S1):340– 345, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://amstat.tandfonline.com/doi/full/
10.1080/00031305.2018.1537888.

Tuyl:2017:NPM

[Tuy17] Frank Tuyl. A note on priors for the multinomial model. The American Statistician, 71(4):298–301, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://amstat.tandfonline.com/doi/full/10.1080/
00031305.2016.1222309.

Tuyl:2019:MHZ

[Tuy19] Frank Tuyl. A method to handle zero counts in the multinomial model. The American Statistician, 73(2):151–158, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2018.1444673.

Utts:2015:EN

[UL15] Jessica Utts and Nicole Lazar. Editors’ note. The American Statistician, 69(2):63, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Utts:2015:MFS

[Utt15] Jessica Utts. The many facets of statistics education: 175 years of common themes. The American Statistician, 69(2):100–107, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Vance:2015:RDT

[Van15] Eric A. Vance. Recent developments and their implications for the future of academic statistical consulting centers. The American Statistician, 69(2):127–137, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

vonCollani:2010:RDF

[vC10] Elart von Collani. Response to “Desired and Feared — What Do We Do Now and Over the Next 50 Years” by Xiao-Li Meng. The American Statistician, 64(1):23–25, February 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

vanDoorn:2018:BIK

[vDLMW18] Johnny van Doorn, Alexander Ly, Maarten Marsman, and Eric-Jan Wagenmakers. Bayesian inference for Kendall’s rank correlation coefficient. The American Statistician, 72(4):303–308, 2018. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2016.1264998.

vanDongen:2019:MPI

[vDvDG^{+}19] Noah N. N. van Dongen, Johnny B. van Doorn, Quentin F. Gronau, Don van Ravenzwaaij, Rink Hoekstra, Matthias N. Haucke, Daniel Lakens, Christian Hennig, Richard D. Morey, Saskia Homer, Andrew Gelman, Jan Sprenger, and Eric-Jan Wagenmakers. Multiple perspectives on inference for two simple statistical scenarios. The American Statistician, 73(S1):328–339, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2019.1565553.

Vellaisamy:2015:PPC

[Vel15] P. Vellaisamy. On probabilistic proofs of certain binomial identities. The American Statistician, 69(3):241–243, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic).  URL http://www.tandfonline.com/doi/abs/10.
1080/00031305.2015.1056381.

Velilla:2018:NCD

[Vel18a] Santiago Velilla. A note on collinearity diagnostics and centering. The American Statistician, 72(2):140–146, 2018. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2016.1264312. See comment [Chr18a].

Velilla:2018:R

[Vel18b] Santiago Velilla. Reply. The American Statistician, 72(1):117– 119, 2018. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://amstat.tandfonline.com/doi/full/
10.1080/00031305.2017.1398985. See [Chr18a, Vel18a].

Vollmer:2017:RDW

[VKP^{+}17] Lauren Vollmer, Aparna Keshaviah, Dmitriy Poznyak, Sharon Zhao, Fei Xing, and Nicholas Beyler. Re-defining the Who, When, and Where of mentoring for professional statisticians. The American Statistician, 71(1):34–37, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Vance:2017:BTS

[VLZ17] Eric A. Vance, Donna E. LaLonde, and Lin Zhang. The big tent for statistics: Mentoring required. The American Statistician, 71(1):15–22, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Vardeman:2013:MVI

[VM13] Stephen B. Vardeman and Max D. Morris. Majority voting by independent classifiers can increase error rates. The American Statistician, 67(2):94–96, 2013. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See comments [BXHH14].

Vardeman:2014:R

[VM14] Stephen B. Vardeman and Max D. Morris. Reply. The American Statistician, 68(2):127, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See [ST13].

Veresoglou:2015:EBD

[VR15] Stavros D. Veresoglou and Matthias C. Rillig. Evidence-based data analysis: Protecting the world from bad code? Comment by Veresoglou and Rillig. The American Statistician, 69(3):257, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00031305.2015.1056831.

Vexler:2014:SDB

[VTH14] Albert Vexler, Wan-Min Tsai, and Alan D. Hutson. A simple density-based empirical likelihood ratio test for independence. The American Statistician, 68(3):158–169, 2014. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Vance:2017:ESG

[VTK^{+}17] Eric A. Vance, Erin Tanenbaum, Amarjot Kaur, Mark C. Otto, and Richard Morris. An eight-step guide to creating and sustaining a mentoring program. The American Statistician, 71(1): 23–29, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Vardeman:2010:ESM

[VWB^{+}10] Stephen B. Vardeman, Joanne R. Wendelberger, Tom Burr, Michael S. Hamada, Leslie M. Moore, J. Marcus Jobe, Max D. Morris, and Huaiqing Wu. Elementary statistical methods and measurement error. The American Statistician, 64(1):46–51, February 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Vexler:2016:DDC

[VZH16] Albert Vexler, Li Zou, and Alan D. Hutson. Data-driven confidence interval estimation incorporating prior information with an adjustment for skewed data. The American Statistician, 70 (3):243–??, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wagaman:2016:MSN

[Wag16] Amy Wagaman. Meeting student needs for multivariate data analysis: A case study in teaching an undergraduate multivariate data analysis course. The American Statistician, 70(4):405– 412, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Waller:2018:DED

[Wal18] Lance A. Waller. Documenting and evaluating data science contributions in academic promotion in departments of statistics and biostatistics. The American Statistician, 72(1):11–19, 2018.

CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2017.1375988.

Wang:2010:CFP

[Wan10] Hsiuying Wang. Closed form prediction intervals applied for disease counts. The American Statistician, 64(3):250–256, August 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wang:2018:PCI

[Wan18] Weizhen Wang. A “Paradox” in confidence interval construction using sufficient statistics. The American Statistician, 72(4):315– 320, 2018. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://amstat.tandfonline.com/doi/full/
10.1080/00031305.2017.1305292.

Wang:2019:BRB

[Wan19] Xin Wang. Book review: Business Survival Analysis Using SAS: An Introduction to Lifetime Probabilities. The American Statistician, 73(2):208–209, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://amstat.

tandfonline.com/doi/full/10.1080/00031305.2018.1538851.

Ward:2017:BBR

[War17] Mark Daniel Ward. Building bridges: The role of an undergraduate mentor. The American Statistician, 71(1):30–33, 2017. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wasserstein:2015:CPI

[Was15] Ron Wasserstein. Communicating the power and impact of our profession: A heads up for the next Executive Directors of the ASA. The American Statistician, 69(2):96–99, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wolkewitz:2010:TPS

[WASB10] Martin Wolkewitz, Arthur Allignol, Martin Schumacher, and Jan Beyersmann. Two pitfalls in survival analyses of time-dependent exposure: a case study in a cohort of Oscar nominees. The American Statistician, 64(3):205–211, August 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Westgate:2017:CCS

[WB17] Philip M. Westgate and Woodrow W. Burchett. A comparison of correlation structure selection penalties for generalized estimating equations. The American Statistician, 71(4):344–353, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2016.1200490.

Wu:2018:MOP

[WB18] Steven Wu and Luke Bornn. Modeling offensive player movement in professional basketball. The American Statistician, 72(1):72– 79, 2018. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://amstat.tandfonline.com/doi/full/
10.1080/00031305.2017.1395365.

Wickham:2018:ISI

[WBL18] Hadley Wickham, Jennifer Bryan, and Nicole Lazar. Introduction: Special issue on data science. The American Statistician, 72(1):1, 2018. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://amstat.tandfonline.com/doi/
full/10.1080/00031305.2018.1438699.

White:2019:MEF

[WBNTF19] Philip A. White, Candace Berrett, E. Shannon Neeley-Tass, and Michael G. Findley. Modeling efficiency of foreign aid allocation in Malawi. The American Statistician, 73(4):385–399, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/00031305.2018.1470032.

Walker:2018:IBU

[WDCH18] M. L. Walker, Y. H. Dovoedo, S. Chakraborti, and C. W. Hilton. An improved boxplot for univariate data. The American Statistician, 72(4):348–353, 2018. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://amstat.

tandfonline.com/doi/full/10.1080/00031305.2018.1448891.

Warr:2013:SIR

[WE13] Richard L. Warr and Roger A. Erich. Should the interquartile range divided by the standard deviation be used to assess normality? The American Statistician, 67(4):242–244, 2013. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See comment [Muk14a] and reply [WE14].

Warr:2014:R

[WE14] Richard L. Warr and Roger A. Erich. Reply. The American Statistician, 68(4):317, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See [WE13].

Westfall:2014:KPR

[Wes14] Peter H. Westfall. Kurtosis as peakedness, 1905–2014. r.i.p. The American Statistician, 68(3):191–195, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

West:2011:OCS

[WG11] Brady T. West and Andrzej T. Galecki. An overview of current software procedures for fitting linear mixed models. The American Statistician, 65(4):274–282, November 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wetzels:2012:DBH

[WGW12] Ruud Wetzels, Raoul P. P. P. Grasman, and Eric-Jan Wagenmakers. A default Bayesian hypothesis test for ANOVA designs. The American Statistician, 66(2):104–111, 2012. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wechsler:2013:BLN

[WIE13] Sergio Wechsler, Rafael Izbicki, and Luís Gustavo Esteves. A Bayesian look at nonidentifiability: A simple example. The American Statistician, 67(2):90–93, 2013. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wierman:2016:CJC

[Wie16] John C. Wierman. The class joke contest: Encouraging creativity and improving attendance. The American Statistician, 70(3):257– ??, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Witmer:2017:BMU

[Wit17] Jeff Witmer. Bayes and MCMC for undergraduates. The American Statistician, 71(3):259–264, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://amstat.

tandfonline.com/doi/full/10.1080/00031305.2017.1305289.

Wang:2019:EDM

[WJ19] Jianjun Wang and Dallas E. Johnson. An examination of discrepancies in multiple imputation procedures between SAS^{T M} and SPSS^{T M}. The American Statistician, 73(1):80–88, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2018.1437078.

Webster:2013:EOS

[WK13] Anthony J. Webster and Richard Kemp. Estimating omissions from searches. The American Statistician, 67(2):82–89, 2013. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wright:2019:PVC

[WKW19] Tommy Wright, Martin Klein, and Jerzy Wieczorek. A primer on visualizations for comparing populations, including the issue of overlapping confidence intervals. The American Statistician, 73(2):165–178, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/00031305.2017.1392359.

Wang:2016:STS

[WL16a] Min Wang and Guangying Liu. A simple two-sample Bayesian t-test for hypothesis testing. The American Statistician, 70(2): 195–201, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wasserstein:2016:ASV

[WL16b] Ronald L. Wasserstein and Nicole A. Lazar. The ASA’s statement on p-values: Context, process, and purpose. The American Statistician, 70(2):129–133, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See response [IGRP17].

Wang:2019:ESC

[WRH19] Xiaofei Wang, Nicholas G. Reich, and Nicholas J. Horton. Enriching students’ conceptual understanding of confidence intervals: An interactive trivia-based classroom activity. The American Statistician, 73(1):50–55, 2019. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2017.1305294.

Wright:2012:ENO

[Wri12] Tommy Wright. The equivalence of Neyman optimum allocation for sampling and equal proportions for apportioning the U.S. House of Representatives. The American Statistician, 66(4):217– 224, 2012. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wright:2013:RIB

[Wri13] Kevin Wright. Revisiting Immer’s barley data. The American Statistician, 67(3):129–133, 2013. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wang:2015:MBS

[WS15] Wei Wang and Dylan S. Small. Monotone B-spline smoothing for a generalized linear model response. The American Statistician, 69(1):28–33, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Webb:2010:PIA

[WSF10] Russell Y. Webb, Peter J. Smith, and Abdulla Firag. On the probability of improved accuracy with increased sample size. The American Statistician, 64(3):257–262, August 2010. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wasserstein:2019:MWB

[WSL19] Ronald L. Wasserstein, Allen L. Schirm, and Nicole A. Lazar. Moving to a world beyond “p < 0.05”. The American Statistician, 73(S1):1–19, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/00031305.2019.1583913.

Wang:2011:UIN

[WW11] S. S. J. Wang and M. P. Wand. Using Infer.NET for statistical analyses. The American Statistician, 65(2):115–126, May 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wang:2019:DMG

[WWGL19] Feifei Wang, Jian Wang, Alan E. Gelfand, and Fan Li. Disease mapping with generative models. The American Statistician, 73 (3):213–223, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://amstat.tandfonline.com/
doi/full/10.1080/00031305.2017.1392358.

Wick:2017:BAS

[WYG17] Jo A. Wick, Hung-Wen Yeh, and Byron J. Gajewski. A Bayesian analysis of synchronous distance learning versus matched traditional control in graduate biostatistics courses. The American Statistician, 71(2):137–144, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://amstat.tandfonline.com/doi/full/10.1080/
00031305.2016.1247014.

Xu:2016:SSB

[XQL16] Xu Xu, Peter Z. G. Qian, and Qing Liu. Samurai Sudoku-based space-filling designs for data pooling. The American Statistician, 70(1):1–8, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Xu:2014:PGM

[Xu14] Shaoji Xu. A property of geometric mean regression. The American Statistician, 68(4):277–281, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Xia:2016:BIR

[XY16] Yan Xia and Yanyun Yang. Bias introduced by rounding in multiple imputation for ordered categorical variables. The American Statistician, 70(4):358–364, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Yuan:2010:CND

[YB10] Ke-Hai Yuan and Peter M. Bentler. Consistency of normaldistribution-based pseudo maximum likelihood estimates when data are missing at random. The American Statistician, 64 (3):263–267, August 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Ye:2017:CFE

[YC17] Zhi-Sheng Ye and Nan Chen. Closed-form estimators for the gamma distribution derived from likelihood equations. The American Statistician, 71(2):177–181, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://amstat.tandfonline.com/doi/full/10.1080/
00031305.2016.1209129.

Yan:2014:NAO

[YGP14] Jun Yan, Chao Guo, and Laurie E. Paarlberg. Are nonprofit antipoverty organizations located where they are needed? A spatial analysis of the Greater Hartford region. The American Statistician, 68(4):243–252, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Yee:2015:DCB

[YH15] Darrick Yee and Andrew Ho. Discreteness causes bias in percentage-based comparisons: A case study from educational testing. The American Statistician, 69(3):174–181, 2015. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic).  URL http://www.tandfonline.com/doi/abs/10.
1080/00031305.2015.1031828.

Young:2015:CDO

[YJCR15] Derek S. Young, Glenn F. Johnson, Mosuk Chow, and James L. Rosenberger. The challenges in developing an online applied statistics program: Lessons learned at Penn State University. The American Statistician, 69(3):213–220, 2015. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic).  URL http://www.tandfonline.com/doi/abs/10.
1080/00031305.2015.1038583.

Yang:2014:DCW

[YZS^{+}14] Fan Yang, José R. Zubizarreta, Dylan S. Small, Scott Lorch, and Paul R. Rosenbaum. Dissonant conclusions when testing the validity of an instrumental variable. The American Statistician, 68 (4):253–263, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Zabell:2013:PML

[Zab13] Sandy Zabell. Paul Meier on legal consulting. The American Statistician, 67(1):18–21, 2013. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Zhang:2014:UCM

[ZBGZ14] Kai Zhang, Lawrence D. Brown, Edward George, and Linda Zhao. Uniform correlation mixture of bivariate normal distributions and hypercubically contoured densities that are marginally normal. The American Statistician, 68(3):183–187, 2014. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Zhang:2015:BFN

[ZC15] Guangxiang Zhang and John J. Chen. Biostatistics faculty and NIH awards at U.S. medical schools. The American Statistician, 69(1):34–40, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Zhang:2017:CDG

[Zha17] Dabao Zhang. A coefficient of determination for generalized linear models. The American Statistician, 71(4):310–316, 2017. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2016.1256839.

Zhang:2018:MVC

[Zha18] Jin Zhang. Minimum volume confidence sets for two-parameter exponential distributions. The American Statistician, 72(3):213– 218, 2018. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://amstat.tandfonline.com/doi/full/
10.1080/00031305.2016.1264315.

Zhu:2013:DAP

[ZHM^{+}13] Yeyi Zhu, Ladia M. Hernandez, Peter Mueller, Yongquan Dong, and Michele R. Forman. Data acquisition and preprocessing in studies on humans: What is not taught in statistics classes? The American Statistician, 67(4):235–241, 2013. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Zigler:2016:CRB

[Zig16] Corwin Matthew Zigler. The central role of Bayes’ theorem for joint estimation of causal effects and propensity scores. The American Statistician, 70(1):47–54, 2016. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Ziliak:2019:HLY

[Zil19] Stephen T. Ziliak. How large are your G-values? Try Gosset’s Guinnessometrics when a little “p” is not enough. The American Statistician, 73(S1):281–290, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://amstat.tandfonline.com/doi/full/10.1080/
00031305.2018.1514325.

Zimmerman:2014:C

[Zim14] Dale L. Zimmerman. Comment. The American Statistician, 68 (2):85–86, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See [Har14a].

Zhang:2013:ERS

[ZLW13] Chong Zhang, Yufeng Liu, and Zhengxiao Wu. On the effect and remedies of shrinkage on classification probability estimation. The American Statistician, 67(3):134–142, 2013. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Zhu:2011:GSM

[ZNY11] Li Zhu, Liyun Ni, and Bin Yao. Group sequential methods and software applications. The American Statistician, 65(2):127–135, May 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Ziemer:2018:NLH

[ZPL^{+}18] Kathryn Schaefer Ziemer, Bianica Pires, Vicki Lancaster, Sallie Keller, Mark Orr, and Stephanie Shipp. A new lens on high school dropout: Use of correspondence analysis and the statewide longitudinal data system. The American Statistician, 72(2):191– 198, 2018. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://amstat.tandfonline.com/doi/full/
10.1080/00031305.2017.1322002.

Zeng:2011:PFD

[ZPM^{+}11] Chan Zeng, Zhaoxing Pan, Samantha MaWhinney, Anna E. Barón, and Gary O. Zerbe. Permutation and F distribution of tests in the multivariate general linear model. The American Statistician, 65(1):31–36, February 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Zhou:2010:NBI

[ZR10] Xiang Zhou and Jerome P. Reiter. A note on Bayesian inference after multiple imputation. The American Statistician, 64(2):159– 163, May 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Zubizarreta:2011:MSS

[ZRK^{+}11] José R. Zubizarreta, Caroline E. Reinke, Rachel R. Kelz, Jeffrey H. Silber, and Paul R. Rosenbaum. Matching for several sparse nominal variables in a case-control study of readmission following surgery. The American Statistician, 65(4):229–238, November 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Zhu:2017:BGG

[ZSMS17] Li Zhu, Kimberly F. Sellers, Darcy Steeg Morris, and Galit Shmueli. Bridging the gap: A generalized stochastic process for count data. The American Statistician, 71(1):71–80, 2017. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See letter [Mac17] and reply [SMSZ17].

Zuo:2010:CIO

[Zuo10] Yijun Zuo. Is the t confidence interval^{−}X±t_{α}(n−1)(s/\sqrt{\{n\}}) optimal? The American Statistician, 64(2):170–173, May 2010. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Zhang:2012:ELR

[ZXC12] Lingyun Zhang, Xinzhong Xu, and Gemai Chen. The exact likelihood ratio test for equality of two normal populations. The American Statistician, 66(3):180–184, 2012. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Zhang:2019:JCC

[ZZT^{+}19] Hongmei Zhang, Yubo Zou, Will Terry, Wilfried Karmaus, and Hasan Arshad. Joint clustering with correlated variables. The American Statistician, 73(3):296–306, 2019. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.tandfonline.com/doi/full/10.
1080/00031305.2018.1424033.